### v1

# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

Abstract
In this paper, we propose Text-based Open Molecule Generation Benchmark (TOMG-Bench), the first benchmark to evaluate the open-domain molecule generation capability of LLMs. TOMG-Bench encompasses a dataset of three major tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each task further contains three subtasks, with each subtask comprising 5,000 test samples. Given the inherent complexity of open molecule generation, we have also developed an automated evaluation system that helps measure both the quality and the accuracy of the generated molecules. Our comprehensive benchmarking of 25 LLMs reveals the current limitations and potential areas for improvement in text-guided molecule discovery. Furthermore, with the assistance of OpenMolIns, a specialized instruction tuning dataset proposed for solving challenges raised by TOMG-Bench, Llama3.1-8B could outperform all the open-source general LLMs, even surpassing GPT-3.5-turbo by 46.5% on TOMG-Bench. Our codes and datasets are available through https://github.com/phenixace/TOMG-Bench.
1 Introduction
Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals Keiser et al. (2010) to materials science Higuchi et al. (2023). Normally, molecule discovery is a trial and error process Ekins (2024), which requires repetitive experimentation and data analysis Mattern and Grosser (2023). Due to the inefficiency of the traditional techniques, it usually takes more than 10 years to bring a new drug candidate into the market Lee et al. (2018).
With the development of machine learning techniques and the advent of Graph Neural Networks (GNNs) Wang et al. (2023), there has been a significant step forward. As molecules could be represented as graphs, GNN-based methods can capture the structural patterns of the molecule and make accurate predictions. With the assistance of GNNs, researchers could analyse the properties of molecules Cai et al. (2022) and generate new molecule candidates Jin et al. (2018). However, challenges still exist. GNN-based methods struggle to generalize to different tasks Chen et al. (2024), necessitating costly data collection and preparation for different downstream tasks. Moreover, these methods are constrained in their capacity to generate molecules with specific, customized properties, limiting their flexibility in molecular design Li et al. (2024c).
Figure 1: Comparison of Text-Based Targeted Molecule Generation (a) v.s. Text-Based Open Molecule Generation (b).
In Contrast, Large Language Models (LLMs) have shown their great generalization capability Achiam et al. (2023) and could be easily adapted to different research fields. For instance, Cascella et al. (2023) utilizes ChatGPT for supporting clinical practice and Zhen et al. (2024) adopts LLMs as assistants for task planning in the field of Civil Engineering, showing the great potential of LLMs in scientific discovery.
As molecules can be represented as texts by Simplified Molecular Input Line Entry System (SMILES), a linear notation that encapsulates the structure of a chemical compound Weininger (1988), they can be processed and understood by LLMs, bridging the gap between molecules and natural languages. With advanced reasoning and in-context learning capabilities, LLMs are particularly adept at generalizing to the molecule domain Li et al. (2024b). The generalization capability makes LLMs a viable option in molecule discovery. Furthermore, by aligning molecules with textual data, LLMs can serve as powerful assistants to chemists Zhang et al. (2024); Li et al. (2024d). They could help interpret and generate chemical knowledge, suggest modifications to molecular structures, and even predict the properties and behaviours of compounds, which would potentially streamline the molecule discovery process, leading to breakthroughs in diverse research areas.
While the integration of LLMs into molecule discovery holds immense promise, the process of aligning molecules with textual data is challenging Li et al. (2024a). A significant challenge lies in the availability and diversity of datasets and benchmarks necessary for training and evaluation. Although the task of molecule-caption translation Edwards et al. (2022) is crucial for bridging the gap between the molecular and textual domains, it still has several limitations that need to be addressed:
On the one hand, there is a concern about the generalization of the molecule-caption translation task Li et al. (2024b). In real-world scenarios, molecule captions that describe molecular structures can be highly ambiguous, with multiple correct interpretations, while the current molecule-caption translation is actually a targeted generation task. In this case, these models often struggle to generalize to customized molecules, even for seemingly simple examples Li et al. (2024b). This suggests a fundamental mismatch between the molecular and textual spaces, raising questions about whether this task could truly guide LLMs well. On the other hand, a critical issue is the inability to propose new molecule structures. The ultimate goal of molecule discovery is not just to understand and describe existing chemical compounds but to innovate and discover new ones, particularly in the context of drug discovery, indicating that the current molecule-caption translation task and the corresponding evaluation metrics fall short in this regard. Therefore, addressing these challenges is essential for harnessing the full potential of LLMs in molecule discovery.
In our efforts to bridge the gap between the natural language and the molecular spaces and to further facilitate LLMs as chemist assistants in molecule discovery, we propose a novel benchmark, Text-based Open Molecule Generation Benchmark (TOMG-Bench). TOMG-Bench is designed to evaluate the open-domain generative capabilities of LLMs in the molecular domain through a series of structured instructions for molecule design and operations. As shown in Figure 1, different from the previous targeted molecule generation tasks, Text-based Open Molecule Generation does not set a specific target or enables LLMs to generate an exact matched molecule. Instead, we adopt chemical toolboxes like RDKit Landrum (2013) to test whether the generation meets the requirements. In other words, there can be multiple correct answers for a single question, and LLMs are only required to generate one of them. Notably, TOMG-Bench is meticulously categorized into three primary tasks, i.e., molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each category contains three subtasks, and each subtask is composed of 5,000 test samples, providing a comprehensive and robust assessment of whether LLMs truly grasp the molecular space. Meanwhile, we also propose a different set of evaluation metrics to evaluate and rank the performance of LLMs, which considers both the accuracy and quality of the generated molecules. Moreover, we propose an instruction-tuning dataset, OpenMolIns, by extracting and reformatting molecules from an existing molecule database. OpenMolIns is structured across five distinct data levels (i.e., light, small, medium, large, and extra-large) to tailor different training purposes.
To encapsulate our contributions, they are primarily threefold:
  • Firstly, we are the first to propose the task of Text-based Open Molecule Generation, along with a new set of evaluation metrics to benchmark the performance of LLMs in this task.
  • Secondly, we have benchmarked a suite of 25 LLMs, shedding light on the limitations of the targeted molecule generation.
  • Thirdly, we introduce an instruction tuning set, OpenMolIns, which helps Llama-3.1-8B-Instruct to outperform all the open-source general LLMs in TOMG-Bench.
2 Related Work
In this section, we briefly review related work about developing Artificial Intelligences (AIs) in Molecule Discovery and, more specifically, in text-based molecule generation tasks.
2.1 Development of AIs in Molecule Discovery
Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design Du et al. (2022). Thus, integrating artificial intelligence into molecule discovery has marked a transformative shift in the pharmaceutical landscape, significantly enhancing the efficiency and effectiveness of identifying and developing new therapeutic molecules. Recent advancements in machine learning (ML), deep learning (DL), and natural language processing (NLP) have enabled AI systems to analyze complex biological and chemical data more effectively than traditional methods Wigh et al. (2022); Wu et al. (2018); Zhou et al. (2023). For instance, MolReGPT  Li et al. (2024b) leverages large language models (LLMs) like ChatGPT to learn molecule SMILES strings representation for molecule-caption translation tasks. Moreover, existing studies have explored advanced methods that utilize various AI techniques to further enhance molecule discovery processes, including Convolutional Neural Networks (CNNs) Peng and Zhao (2019); Le et al. (2019), Recurrent Neural Networks (RNNs) Grisoni et al. (2020); Popova et al. (2019), Graph Neural Networks (GNNs) Wang et al. (2023); Sun et al. (2022), and Transformer-based networks Xia et al. (2023); Balaji et al. (2023); Edwards et al. (2022).
2.2 Text-based Molecule Generation
Text-based Molecule Generation (Text2Mol) Edwards et al. (2021) has recently emerged as a transformative approach to molecule discovery. This task centres on retrieving molecules using natural language descriptions as search queries, requiring the creation of paired datasets of molecules and their corresponding textual representations. This enables the learning of a shared semantic embedding space for efficient retrieval. Early approaches leveraged transformer-based models like MolT5 Edwards et al. (2022), employing self-supervised learning on large datasets to generate high-quality Simplified Molecular Input Line Entry System (SMILES) strings from textual inputs. Subsequent advancements, such as KV-PLM Zeng et al. (2022), MoMu Su et al. (2022), and BioT5 Pei et al. (2023), integrated molecular graphs and biochemical text to improve both understanding and generation capabilities. 3D-MoLM Li et al. (2024e) further enhanced this by incorporating spatial configurations, leading to more accurate and geometrically valid molecular representations. The application of large language models (LLMs) like MolReGPT Li et al. (2024b) and ICMA Li et al. (2024a) as in-context learners has also shown significant promise. These models demonstrate the ability of LLMs to adaptively generate molecules by retrieving and leveraging relevant examples from the provided context. Most recently, MolReFlect Li et al. (2024c) underscored the importance of fine-grained alignment between molecular structures and their textual descriptions, utilizing a teacher-student training paradigm to capture these nuanced relationships effectively. Unlike this targeted generation task, in this paper, we propose a Text-based Open Molecule Generation task to enable LLMs to generate an exactly matched molecule rather than set a specific target.
3 TOMG-Bench
In this section, we propose TOMG-Bench to comprehensively assess the performance of LLMs in molecular space. Specifically, the benchmark is composed of three basic tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). To ensure the integrity and effectiveness of TOMG-Bench, we have developed a robust set of evaluation metrics for different tasks. Additionally, we have created OpenMolIns, an instruction-tuning dataset aimed at enhancing the performance and adaptability of LLMs to the challenges presented by this benchmark.
Figure 2: Data construction workflow and evaluation process of TOMG-Bench.
3.1 Dataset Categorization
The categorization of the TOMG-Bench dataset initially considers the inherent characteristics of molecule SMILES representation and the role of LLMs serving as the chemist’s assistant, namely helping the chemists to edit, optimize and customize molecules as they want. Following the difficulty of the tasks, we demonstrate the content of the three basic tasks as well as their corresponding subtasks:
MolEdit emerges as the most straightforward task among the three domains, as an existing molecule is already provided, and LLMs are only required to make modifications to it, which tests the molecular structure knowledge of LLMs. In this case, we have crafted three subtasks for MolEdit: AddComponent, DelComponent, and SubComponent. In AddComponent, LLMs are instructed to add a specific functional group to the given molecule, and DelComponent challenges LLMs to remove a specified functional group from the provided molecule, while SubComponent is a hybrid of the previous two subtasks, requiring LLMs to first remove a designated functional group and then introduce a new one as specified to the molecule.
MolOpt challenges LLMs to not only edit molecules but also to discern whether the modification will steer the molecule towards a desired optimization target. To assess this capability, we concentrate on three pivotal properties that are vital for molecule discovery: LogP (Octanol-water partition coefficient, a metric of lipophilicity), MR (molecular refractivity, a proxy for the molar refractive index), and QED (Quantitative Estimate of Druglikeness, an assessment of drug-like characteristics). These metrics offer critical information about the potential pharmacological attributes of the molecule, which could help chemists filter molecules as viable drug candidates.
MolCustom is the most challenging task, where we have established three subtasks: AtomNum, BondNum, and FunctionalGroup. For AtomNum, LLMs are tasked with generating molecules that adhere to a specified count and type of atoms. BondNum involves the creation of molecules with a defined number and type of bonds. In FunctionalGroup, LLMs must generate a molecule that includes functional groups as specified. These subtasks may appear straightforward, yet they are deceptively challenging. They demand that LLMs have a sophisticated understanding of molecular syntax to precisely generate molecules that meet the complex criteria set forth.
3.2 Dataset Construction
Previously, molecule-related datasets and tasks have been hindered by the scarcity of human annotations. For instance, ChEBI-20  Edwards et al. (2022), a dataset for the molecule-caption translation task, contains only 33,000 samples for training, whereas image captioning datasets like MS COCO  Chen et al. (2015) have over 1,500,000 annotated captions on more than 330,000 images. Meanwhile, the annotation of molecules demands expertise from chemists and can sometimes require wet lab experiments, which are both expensive and time-consuming. In contrast, TOMG-Bench, as an open-domain generation task, does not rely on human annotations for construction. Instead, we focus on basic molecule structural properties and basic molecule operations that could be validated by chemical toolboxes to construct our dataset.
For MolCustom, we randomly generate 5,000 prompts as requests for each subtask, requiring different numbers and collections of atoms, bonds, and functional groups. For MolEdit and MolOpt, we sample molecules from a specific molecule database. Specifically, we select two molecule databases for this work: Zinc-250K Sterling and Irwin (2015), and PubChem  Kim et al. (2019). Zinc-250K has 250,000 molecules, which is smaller than PubChem, which has 10 million molecules. To facilitate the fast calculation of metrics mentioned in Section 3.5, we choose Zinc-250K for sampling the test molecules in TOMG-Bench. Each subtask is allocated 5,000 test samples. After sampling, we utilize RDKit Landrum (2013), a molecular informatics toolbox, to collect basic molecule statistics. There are functions available to calculate the required characteristics, especially the structural patterns and chemical properties like LogP, MR, and QED values, which can then be integrated into our pre-defined task prompts. Further details will be provided in the Appendix A.
3.3 Metric Design
The evaluation of the TOMG-Bench is facilitated through a set of carefully designed metrics tailored to the specific tasks within the benchmark.
For the MolCustom task, which includes subtasks such as AtomNum, BondNum, and FunctionalGroup, the following metrics are employed:
  • Accuracy: This metric measures the proportion of generated molecules that successfully meet the criteria. A high accuracy indicates that the LLM is effectively generating molecules according to the request.
  • Novelty: Novelty is a critical metric for MolCustom tasks, as it assesses the ability of LLMs to discover new molecular structures. This is particularly important for potential drug or material candidates, where innovation can lead to significant advancements.
  • Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax.
For the MolEdit and MolOpt tasks, the following metrics are adopted:
  • Accuracy: Similar to the MolCustom category, accuracy here measures the proportion of edited or optimized molecules that could pass the test process.
  • Similarity: Similarity between the original molecule and the edited molecule is an important point to concern. Since there are multiple valid solutions, only considering the accuracy is not sufficient for verification. For implementation, a high similarity indicates that the generated molecule can be modified from the original molecule via as few as possible steps. For instance, when removing a specific functional group, the edited molecule should still be similar to the original molecule. This process should not transform the molecule into an entirely new one nor inadvertently remove additional components.
  • Validity: Similarly, validity is still a crucial metric to confirm that the edited or optimized molecules adhere to the principles of the molecule textual representation.
Notably, the calculation of novelty and similarity metrics only considers valid molecules. Specifically, for the novelty metric, we assess the similarity between the generated molecule and those within the Zinc-250K database.
To comprehensively evaluate the average performance of LLMs on TOMG-Bench, we introduce a weighted average accuracy to rank the performance of LLMs. Considering that the novelty scores for MolCustom and similarity scores for MolEdit and MolOpt are also crucial metrics for evaluating the performance, especially the similarity scores, which help identify correct molecule editing operations. In this case, we adopt the novelty and similarity scores as the weights, and the weighted average accuracy can be computed as follows:
$\bar{w\!Acc}\!=\!\frac{1}{9}\!(\sum_{t\in\!\!M\!o\!l\!C\!u\!s\!t\!o\!m\!}\!(\!n_{t}*Acc_{t}\!)\!+\sum_{t\in\!\{\!M\!o\!l\!E\!d\!i\!t,\!M\!o\!l\!O\!p\!t\!\}}\!(\!s_{t}*Acc_{t}\!)),$ (1)
where $n_{t}$ represents the novelty score for the MolCustom tasks and $s_{t}$ is the similarity score for the MolEdit and MolOpt tasks. $Acc_{t}$ is the accuracy for each subtask $t$. This weighted average accuracy, $\bar{wAcc}$, provides a balanced measure of performance that considers both the accuracy and quality of the generation.
3.4 OpenMolIns: Instruction Tuning Dataset
In this section, we introduce OpenMolIns, a specialized dataset derived from the PubChem database to help LLMs get familiar with text-based open molecule generation via instruction tuning. This instruction tuning dataset is meticulously designed to ensure that the molecules it contains do not overlap with those in the Zinc-250K dataset to promote the generation of more novel molecular structures and avoid any potential data leakage that could compromise the integrity of the model performance.
We collect the instruction tuning dataset by introducing the samples of the nine subtasks in equal amounts. We still apply the RDKit toolbox to construct the instruction tuning dataset.
In the MolCustom domain, constructing the instruction tuning samples is rather straightforward. We calculate the molecular statistics and encapsulate them within prompts. For example, for AtomNum, we count all the atoms we would consider in the molecule, including their types and numbers. Then, we wrap these statistics with a random pre-defined prompt to construct the training sample. This method allows the generated molecules to better fit the distribution of molecular space.
For the MolEdit and MolOpt domains, we randomly select a functional group from the original molecule for addition or removal, utilizing the RDKit toolbox to execute these operations. Then, similarly, we wrap the original molecule and the edited molecule with a random pre-defined prompt to construct the training sample for MolEdit. In the case of MolOpt, we also evaluate the direction of the desired property changes by the functions in RDKit to determine whether the operation improves or decreases the property value, which is then altogether wrapped by the pre-defined prompts.
To investigate the impact of data scales on the performance of LLMs, we have established five distinct data levels tailored for different training purposes: light, small, medium, large, and xlarge, as illustrated in Section 3.5. Each level represents a different quantity of data, shown in Table 1, allowing us to analyze how the amount of training data influences the model’s ability to learn and generate or edit molecules effectively.
Item Data Size
TOMG-Bench
subtask 5,000
OpenMolIns
light 4,500
small 18,000
medium 45,000
large 90,000
xlarge 1,200,000
Table 1: Statisics of TOMG-Bench and OpenMolIns.
3.5 Statistics
In this section, we outline the basic statistics of TOMG-Bench along with the OpenMolIns dataset. Table 1 shows the details of the data size in TOMG-Bench as well as the OpenMolIns datset. For TOMG-Bench, we have three main tasks with nine subtasks in total, where each subtask contains 5,000 test samples. More details can be found in Appendix A.
For OpenMolIns, we have five distinct data scales: light, small, medium, large, and xlarge, ranging from 4,500 to 1,200,000 examples, which helps us investigate the data scaling law of applying LLMs to the Text-based Open Molecule Generation task. Notably, the nine subtasks within the TOMG-Bench are uniformly distributed in the OpenMolIns dataset.
4 Experiments
In this section, we present the experiment setup and results. Then, we illustrate our findings based on the observations.
4.1 Setup
4.1.1 Models
The models benchmarked are categorized into four groups: proprietary models, open-source general LLMs, open-source ChEBI-20 fine-tuned LLMs, and OpenMolIns fine-tuned LLMs.
Proprietary Models. This category includes LLMs that are only accessible via commercial API services. In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo Achiam et al. (2023), Claude-3.5 Anthropic (2024b), Claude-3 Anthropic (2024a), and Gemini-1.5-pro Deepmind (2024).
Open-source General LLMs. This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications. Specifically, we benchmark Llama-3-70B-Instruct, Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-1B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), Qwen2-7B-Instruct Yang et al. (2024), yi-1.5-9B Young et al. (2024), and chatglm-9B GLM et al. (2024).
Open-source ChEBI-20 Fine-tuned LLMs. LLMs fine-tuned on the ChEBI-20 dataset can grasp some extent of text-based molecule generation capability. In this case, our experiments also cover LLMs like MolT5-small, MolT5-base, MolT5-large Edwards et al. (2022), and BioT5-base Pei et al. (2023).
OpenMolIns Fine-tuned LLMs We further adopt LLMs like Galactica-125M Taylor et al. (2022), Llama3.2-1B-Instruct, and Llama-3.1-8B-Instruct on OpenMolIns dataset for instruction tuning. We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M.
4.1.2 Implementation Details
We implement various scripts to facilitate the testing of the aforementioned models. For proprietary models, we adopt the OpenAI API  [^1] https://platform.openai.com/docs/ framework. For open-source general LLMs, we utilize both the VLLM  [^2] https://docs.vllm.ai framework and the OpenAI framework. For the remaining LLMs, we adopt the Hugging Face transformers library  [^3] https://huggingface.co/docs/transformers/ for inference. Detailed hyper-parameters are demonstrated in Appendix B.
Furthermore, it is important to note that BioT5 is designed to use SELFIES as input instead of SMILES. Consequently, we convert the molecule SMILES strings into SELFIES format on BioT5.
4.2 Results
Figure 3: The performance of models benchmarked in TOMG-Bench. In TOMG-Bench, LLMs are divided into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs. Models whose parameters are known are plotted as dots, while models of unknown parameters are denoted as horizontal lines.
Figure 3 presents the performances and model sizes of different models benchmarked on TOMG-Bench, as well as the instruction-tuning performance of Galactica-125M on the five distinct data levels of OpenMolIns. More precise details are further illustrated in Appendix C.
4.3 Findings
Based on the above results, we observe the following key findings:
Text-based open molecule generation is challenging for LLMs. As illustrated in Figure 3, we have calculated the weighted average accuracy across all the nine subtasks. Among the LLMs benchmarked, Claude-3.5 stands out as the top performer, achieving a weighted average accuracy of 35.92%. Gemini-1.5-pro follows closely with a weighted average accuracy of 34.80%. These results underscore the considerable scope for improvement, even among the most advanced proprietary LLMs.
It is also worth noting that while the most advanced LLMs like Claude-3.5 and GPT-4o exhibit relatively strong performance in the MolEdit and MolOpt tasks, the more challenging MolCustom task still remains a challenge. In MolCustom tasks, no LLM has managed to achieve an accuracy exceeding 25% for a single subtask. This observation indicates that the generation of molecules from scratch demands a deep understanding of the molecular structural space, an area where current models are still striving to make significant strides.
Most powerful open-source general LLMs can already outperform GPT-3.5-turbo. In the TOMG-Bench, Llama-3-70B-Instruct has achieved an impressive weighted average accuracy of 23.93%, notably outperforming GPT-3.5-turbo, which scored 18.58%. Despite previously lagging behind proprietary models, open-source general LLMs have rapidly bridged the gap. The evolution of the Llama series, in particular, has been remarkable, finally surpassing GPT-3.5-turbo and demonstrating the fast development of open-source general LLMs.
More powerful LLMs inherit a better performance in TOMG-Bench. Across all the LLMs we benchmarked, a clear trend emerged: the more powerful the LLM is, the higher performance it can achieve on the TOMG-Bench. For instance, the GPT series has consistently demonstrated improved performance with each new iteration from GPT-3.5-turbo to GPT-4o.
Similarly, within the Llama-3 series, we could also observe that larger models tend to achieve superior results on the TOMG-Bench. These findings underscore a strong correlation between an LLM’s capabilities and its performance in our benchmark.
However, we encountered an unexpected anomaly with certain open-source LLMs. Notably, Qwen2-7B-Instruct, despite its impressive ability to solve mathematical problems and its size of 7 billion parameters, underperformed models with as few as 1 billion parameters. This result is particularly striking and suggests that the TOMG-Bench offers a unique and comprehensive evaluation that current races for LLMs may not have adequately addressed. This discovery also highlights the significance of the TOMG-Bench as a new benchmark for LLMs. It provides a broader and more diverse assessment that exposes potential blind spots in the development of LLMs.
ChEBI-20 dataset is insufficient for LLMs to master molecular structures and editing operations. The ChEBI-20 dataset and the associated molecule-caption translation task are designed to bridge the gap between molecular structures and textual descriptions. Despite this intention, LLMs trained on ChEBI-20 have demonstrated limited effectiveness in our TOMG-Bench benchmark. For instance, BioT5-base, which is claimed as the state-of-the-art (SOTA) model for text-based molecule generation on the ChEBI-20 dataset, only achieves a weighted average accuracy of 4.21% on the TOMG-Bench. In the MolEdit and MolOpt tasks, these models are unable to execute correct operations on provided molecules, resulting in disappointing similarity scores. Similarly, in the MolCustom task, which closely mirrors the text-based molecule generation task, the performance remains unsatisfactory, with no model achieving a score above 5% in a single subtask. This performance shortfall highlights a critical limitation of the ChEBI-20 dataset, as it lacks the data quantity and diversity necessary to effectively align molecules with textual descriptions.
In contrast, TOMG-Bench offers a more comprehensive and intricate evaluation framework for text-to-molecule generation. With a larger and more varied set of test examples, TOMG-Bench could robustly assess the capabilities of language models in translating textual descriptions into molecular structures. As such, it represents a significant advancement in the evaluation of text-based molecule generation.
OpenMolIns can enable LLMs to achieve better performance than the most powerful open source general LLMs. We have also developed OpenMolIns, an instruction-tuning dataset, to enhance LLMs’ proficiency in the tasks outlined in the TOMG-Bench. Across five distinct data scales, we observed a pronounced data scaling law: as the size of the corpus increases, the performance of LLMs also improves. In particular, for Galactica-125M, we assessed its capabilities comprehensively on both five distinct data scales. As shown in Figure 3, the outcomes were remarkable: Galactica-125M achieved a weighted average score of 25.73% on OpenMolIns-xlarge, surpassing even the 70B Llama-3-70B-Instruct and GPT-3.5-turbo, with only 125 million parameters. Meanwhile, the results of Galactica-125M show a clear data scaling law, denoting that LLMs are hungry for more molecule corpora to achieve better performance. Notably, OpenMolIns-large has also enabled Llama3.1-8B-Instruct to outperform all the existing open-source general LLMs in TOMG-Bench, showing the effectiveness of the dataset.
5 Conclusion
In this study, we introduce TOMG-Bench, the first benchmark designed to assess the capabilities of Large Language Models (LLMs) in the realm of open-domain molecule generation. Benchmarking 25 LLMs, TOMG-Bench highlights the limitations of existing targeted molecule generation tasks and demonstrates the potential of general LLMs in this domain. Additionally, through instruction tuning on our proposed OpenMolIns, LLMs exhibit significant potential on the TOMG-Bench, matching the performance of GPT-3.5-turbo. Our contributions not only lie in the development of a novel benchmark for molecular discovery but also provide a diverse indication of the capabilities of LLMs.
6 Limitations
Although TOMG-Bench is carefully designed and well-validated through our experiments, we still observe several limitations:
Prompt Diversity. Prompt diversity helps relieve the over-fitting of the instructions. While we adopt several different prompt templates and randomly choose from them, we still find the number of prompt templates is not enough to satisfy the prompt diversity.
Data Distribution. In our data construction process, we allocate distributions to atoms, bonds, and functional groups with the aim of making our benchmark more reflective of real-world distributions. Nevertheless, the distribution we use is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios accurately. This could potentially mask the true performance capabilities of LLMs in these specific tasks.
7 Acknowledgments
We thank all the reviewers for their insightful comments. We also thank GitHub Copilot for coding assistance and ChatGLM-4 for polishing the writing.
Appendix A Data Construction
In this section, we introduce the construction details of TOMG-Bench and OpenMolIns dataset, as well as the prompt templates.
A.1 MolEdit
For the molecule editing (MolEdit) task, we consider the common operations on modifying functional groups in a given molecule (i.e., add, drop, and substitute), which are simple tasks for human experts but challenging to LLMs. In this case, we further develop three corresponding subtasks: AddComponent, DelComponent, and SubComponent. Prompt templates for MolEdit are shown in Table 2. However, there are different kinds of functional groups, and some functional groups can play an important role in the molecule structure, such as connecting two separate parts of the molecule, which makes them unsuitable for these operations above as these operations will entirely change the structure of the molecule. In this case, we aim to make a slight change in the molecule structure and limit most of the functional groups we choose within the end groups.
Prompt Templates for MolEdit
AddComponent
Please add a {} to the molecule {}.
Modify the molecule {} by adding a {}.
Add a {} to the molecule {}.
DelComponent
Please remove a {} from the molecule {}.
Modify the molecule {} by removing a {}.
Remove a {} from the molecule {}.
SubComponent
Please substitute a {} in the molecule {} by {}.
Modify the molecule {} by replacing a {} by {}.
Replace a {} in the molecule {} by {}.
Please replace a {} in the molecule {} with {}.
Modify the molecule {} by substituting a {} with {}.
Substitute a {} in the molecule {} with {}.
Table 2: Prompt Templates for MolEdit
Table 3 presents the functional groups that are taken into account for AddComponent and DelComponent, along with their respective selection weights. To reflect the distribution of these functional groups in real-world scenarios, we have implemented a weighted random selection process for AddComponent, which ensures that less common functional groups are assigned a lower probability to be chosen, thereby refining the selection mechanism to better mirror practical occurrences.
Functional Group benzene ring hydroxyl aldehyde carboxyl amide
Weights 15 15 5 5 10
Functional Group amine nitro halo nitrile thiol
Weights 5 5 5 1 1
Table 3: Functional Groups that are considered in AddComponent and DelComponent, as well as their weights to be selected in AddComponent.
For SubComponent, our focus is exclusively on end groups, which include hydroxyl, aldehyde, carboxyl, nitro, halo, nitrile, and thiol, which ensures that the editing operations are confined to substituting the existing functional group with another from this list, thereby maintaining the integrity of the molecule’s overall structure without altering it fundamentally.
A.2 MolOpt
Molecule optimization (MolOpt), designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task. Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time. In contrast, TOMG-Bench requires one single LLM to optimize molecules with different metrics and directions. In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED. The prompt templates for MolOpt are illustrated in Table 4.
Prompt Templates for MolOpt
LogP
Please optimize the molecule {} to have a lower/higher LogP value.
Modify the molecule {} to decrease/increase its LogP value.
Optimize the molecule {} to have a lower/higher LogP value.
Please modify the molecule {} to decrease/increase its LogP value.
Modify the molecule {} to have a lower/higher LogP value.
MR
Please optimize the molecule {} to have a lower/higher MR value.
Modify the molecule {} to decrease/increase its MR value.
Optimize the molecule {} to have a lower/higher MR value.
Please modify the molecule {} to decrease/increase its MR value.
Modify the molecule {} to have a lower/higher MR value.
QED
Please optimize the molecule {} to have a lower/higher QED value.
Modify the molecule {} to decrease/increase its QED value.
Optimize the molecule {} to have a lower/higher QED value.
Please modify the molecule {} to decrease/increase its QED value.
Modify the molecule {} to have a lower/higher QED value.
Table 4: Prompt Templates for MolOpt
LogP refers to the logarithm of the partition coefficient, which is a measure of a molecule’s hydrophilicity or lipophilicity. It is an important factor in determining a compound’s bioavailability and membrane permeability.
Molecular Refractivity (MR) is a measure of the molar refractive index, which provides insight into the molecular size and the degree of molecular branching. It is used to assess the overall shape and bulk of a molecule.
Quantitative Estimation of Drug-Likeness (QED) is a computational metric that evaluates the drug-likeness of a molecule based on a set of predefined rules. A higher QED score suggests a greater likelihood that the molecule will have favourable pharmacological properties.
A.3 MolCustom
To enable customized design of molecules, we think of three fundamental features for describing the molecule, including atoms, bonds, and functional groups. Given the specified category and number of atoms, bonds, and functional groups, LLMs should generate the molecule as we request. The prompt templates for MolCustom are shown in Table 5. Below, we present the construction details of the three subtasks for MolCustom:
Prompt Templates for MolCustom
AtomNum
Please generate a molecule with {} atom(s).
Please generate a molecule composed of {} atom(s).
Please generate a molecule consisting {} atom(s).
The molecule has {} atom(s).
The molecule is composed of {} atom(s).
The molecule consists of {} atom(s).
There is a molecule with {} atom(s).
There is a molecule composed of {} atom(s).
There is a molecule consisting of {} atom(s).
The molecule contains {} atom(s).
BondNum
Please generate a molecule with {} bond(s).
Please generate a molecule composed of {} bond(s).
Please generate a molecule consisting {} bond(s).
The molecule has {} bond(s).
The molecule is composed of {} bond(s).
The molecule consists of {} bond(s).
There is a molecule with {} bond(s).
There is a molecule composed of {} bond(s).
There is a molecule consisting of {} bond(s).
The molecule contains {} bond(s).
FunctionalGroup
Please generate a molecule with {} group(s).
Please generate a molecule composed of {} group(s).
Please generate a molecule consisting {} group(s).
The molecule has {} group(s).
The molecule is composed of {} group(s).
The molecule consists of {} group(s).
There is a molecule with {} group(s).
There is a molecule composed of {} group(s).
There is a molecule consisting of {} group(s).
The molecule contains {} group(s).
Table 5: Prompt Templates for MolCustom
Atom carbon oxygen nitrogen sulfur fluorine chlorine bromine iodine phosphorus
Weights [Mandatory] 5 3 3 2 2 2 2 1
Atom boron silicon selenium tellurium arsenic antimony bismuth polonium
Weights 1 1 1 1 1 1 1 1
Table 6: Atoms that are considered in AtomNum, as well as their weights to be selected.
AtomNum. Table 6 shows the atoms we consider in AtomNum, as well as their weights to be selected. Notably, carbon, as the basic unit in organic chemicals, is a mandatory option. The number of carbon atoms ranges from 1 to 40, while the number of other selected atoms ranges from 1 to 5. This setting relieves the difficulty for generation, as LLMs could generate a carbon backbone first and attach the remaining atoms to the backbone one by one.
Bond single double triple rotatable aromatic
Weights 5 4 3 1 1
Table 7: Chemical bonds that are considered in BondNum, as well as their weights to be selected.
BondNum. Similarly, we select five different kinds of chemical bonds: single, double, triple, rotatable, and aromatic, as shown in Table 7. For the single bond, if selected, the number can vary from 1 to 50. For the aromatic bond, the number follows the rules of the formation of aromatic bonds, varying from 5 to 20. Moreover, the number of these remaining bonds, if selected, is specified from 1 to 5.
Functional Group benzene ring hydroxyl anhydride aldehyde ketone carboxyl ester amide amine nitro
Weights 15 15 2 5 5 10 5 5 5 2
Functional Group halo thioether nitrile thiol sulfide disulfide sulfoxide sulfone borane
Weights 2 1 1 1 1 1 1 1 1
Table 8: Functional Groups that are considered in FunctionalGroup, as well as their weights to be selected.
FunctionalGroup. Lastly, we also specify functional groups in the molecule structure. Table 8 shows the range of functional groups and their weights that are taken into consideration.
Notably, in MolCustom, if not specified, LLMs can generate any number of these atoms, bonds, and functional groups. However, for these specified items, LLMs should strictly follow the requirements.
Appendix B Hyper Parameters
In this section, we illustrate the detailed parameters adopted in this work, as shown in Table 9.
Item Value
Generation
temperature 0.75
top_p 0.85
num_beams 1
max_new_tokens 512
Instruction Tuning
batchsize 32
lr 3e-4
cutoff_len 1024
Lora Settings
r 64
$\alpha$ 128
dropout 0.1
Table 9: Hyper-parameters
Appendix C Detailed Results
In this section, we first show the leaderboard of TOMG-Bench in Table 10, where Claude-3.5 achieves first place with a weighted average accuracy of 35.92%. Notably, via instruction tuning on our OpenMolIns dataset, Llama-3.1-8B achieves 6th place, which outperforms all the existing open-source LLMs and is just behind Cladue-3.
Model #Parameters (B) $\bar{Acc}$ (%) $\bar{wAcc}(\%)$
Claude-3.5 Anthropic (2024b) - 51.10 35.92
Gemini-1.5-pro Deepmind (2024) - 52.25 34.80
GPT-4-turbo Achiam et al. (2023) - 50.74 34.23
GPT-4o Achiam et al. (2023) - 49.08 32.29
Claude-3 Anthropic (2024a) - 46.14 30.47
OpenMolIns-large (Llama-3.1-8B) 8 43.1 27.22
OpenMolIns-xlarge (Galactica-125M) 0.125 44.48 25.73
Llama3-70B-Instruct (Int4) Dubey et al. (2024) 70 38.54 23.93
OpenMolIns-large (Galactica-125M) 0.125 39.28 23.42
OpenMolIns-medium (Galactica-125M) 0.125 34.54 19.89
GPT-3.5-turbo Achiam et al. (2023) - 28.93 18.58
OpenMolIns-small (Galactica-125M) 0.125 24.17 15.18
Llama3.1-8B-Instruct Dubey et al. (2024) 8 26.26 14.09
Llama3-8B-Instruct Dubey et al. (2024) 8 26.40 13.75
chatglm-9B GLM et al. (2024) 9 18.50 13.13(7)
OpenMolIns-light (Galactica-125M) 0.125 20.95 13.13(6)
OpenMolIns-large (Llama3.2-1B) 1 14.11 8.10
yi-1.5-9B Young et al. (2024) 9 14.10 7.32
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 7 11.17 4.81
BioT5-base Pei et al. (2023) 0.25 24.19 4.21
MolT5-large Edwards et al. (2022) 0.78 23.11 2.89
Llama-3.1-1B-Instruct Dubey et al. (2024) 1 3.95 1.99
MolT5-base Edwards et al. (2022) 0.25 11.11 1.30(0)
MolT5-small Edwards et al. (2022) 0.08 11.55 1.29(9)
Qwen2-7B-Instruct Yang et al. (2024) 7 0.18 0.15
Table 10: Leaderboard of TOMG-Benchmark.
Then, we present the detailed results of all the subtasks in Table 11, 12, and 13.
Models AddComponent DelComponent SubComponent
Accuracy Similarity Validity Accuracy Similarity Validity Accuracy Similarity Validity
GPT-4o Achiam et al. (2023) 0.6188 0.6782 0.7412 0.7012 0.6038 0.8474 0.7992 0.7225 0.9368
GPT-4-turbo Achiam et al. (2023) 0.699 0.6936 0.7934 0.7244 0.5735 0.906 0.7778 0.7323 0.916
GPT-3.5-turbo Achiam et al. (2023) 0.5832 0.6545 0.798 0.3082 0.7797 0.8468 0.2918 0.6333 0.6822
Claude-3.5 Anthropic (2024b) 0.6832 0.7017 0.4414 0.5414 0.6678 0.796 0.8104 0.731 0.9588
Claude-3 Anthropic (2024a) 0.6766 0.684 0.818 0.5556 0.6408 0.8984 0.655 0.7159 0.9184
Gemini-1.5-pro Deepmind (2024) 0.7058 0.6792 0.8254 0.759 0.5949 0.9158 0.7148 0.7139 0.8684
Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.5198 0.6801 0.5922 0.6122 0.5637 0.7182 0.5094 0.717 0.6822
Llama3-8B-Instruct Dubey et al. (2024) 0.3914 0.6649 0.5374 0.4348 0.5058 0.57 0.2602 0.6841 0.4838
Llama3.1-8B-Instruct Dubey et al. (2024) 0.2992 0.6088 0.4954 0.4336 0.5257 0.591 0.3401 0.6424 0.5076
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.1868 0.6251 0.376 0.2018 0.3774 0.359 0.0602 0.6227 0.355
Qwen2-7B-Instruct Yang et al. (2024) 0.001 0.2527 0.0036 0.0006 0.4024 0.0012 0.0004 0.2895 0.0068
Yi-1.5-9B Young et al. (2024) 0.1742 0.417 0.4216 0.2858 0.5936 0.4909 0.137 0.4619 0.4368
Chatglm-9B GLM et al. (2024) 0.2932 0.7622 0.5686 0.2956 0.7494 0.6914 0.1498 0.715 0.5084
Llama-3.2-1B-Instruct Dubey et al. (2024) 0.0374 0.5343 0.1982 0.0768 0.575 0.3028 0.0102 0.3671 0.1468
MolT5-small Edwards et al. (2022) 0.122 0.1027 0.449 0.1598 0.1125 0.4504 0.0708 0.1029 0.4876
MolT5-base Edwards et al. (2022) 0.1354 0.1066 0.4686 0.1562 0.1144 0.4472 0.0584 0.1028 0.4426
MolT5-large Edwards et al. (2022) 0.2834 0.1084 0.9282 0.2228 0.1201 0.9198 0.1692 0.0932 0.941
BioT5-base Pei et al. (2023) 0.3462 0.1567 1 0.1668 0.1597 1 0.0684 0.1576 0.9998
OpenMolIns-large (Llama-3.2-1B) 0.1756 0.5676 0.3216 0.1816 0.4963 0.2466 0.0844 0.5415 0.2958
OpenMolIns-large (Llama-3.1-8B) 0.5822 0.6541 0.673 0.5104 0.5074 0.6896 0.544 0.6258 0.84
OpenMolIns-light (Galactica-125M) 0.3786 0.5958 0.3786 0.2062 0.6521 0.7048 0.3102 0.5879 0.6674
OpenMolIns-small (Galactica-125M) 0.3472 0.6172 0.5356 0.3258 0.6025 0.5758 0.2692 0.6181 0.5692
OpenMolIns-medium (Galactica-125M) 0.4736 0.5682 0.7442 0.4886 0.5184 0.7488 0.3282 0.5975 0.6958
OpenMolIns-large (Galactica-125M) 0.5866 0.5876 0.8228 0.6078 0.5577 0.7934 0.3438 0.6491 0.8438
OpenMolIns-xlarge (Galactica-125M) 0.5842 0.5859 0.8438 0.6526 0.5084 0.8286 0.1872 0.6024 0.8538
Table 11: Results on MolEdit. For each task, we highlight the best accuracy and underline the second best accuracy.
Models LogP MR QED
Accuracy Similarity Validity Accuracy Similarity Validity Accuracy Similarity Validity
GPT-4o Achiam et al. (2023) 0.719 0.6586 0.8796 0.6864 0.642 0.8352 0.3952 0.618 0.857
GPT-4-turbo Achiam et al. (2023) 0.7662 0.6984 0.9048 0.7388 0.6821 0.8848 0.3946 0.6587 0.905
GPT-3.5-turbo Achiam et al. (2023) 0.4048 0.6327 0.854 0.412 0.6263 0.8486 0.3316 0.5635 0.8354
Claude-3.5 Anthropic (2024b) 0.797 0.7124 0.9422 0.6962 0.7112 0.911 0.5361 0.7042 0.8604
Claude-3 Anthropic (2024a) 0.7984 0.6067 0.9096 0.6094 0.6398 0.9062 0.4678 0.5855 0.9044
Gemini-1.5-pro Deepmind (2024) 0.7712 0.7022 0.9274 0.7876 0.6744 0.8926 0.4704 0.6077 0.9484
Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.5984 0.6028 0.6482 0.5684 0.6032 0.6272 0.2774 0.4828 0.634
Llama3-8B-Instruct Dubey et al. (2024) 0.4642 0.3658 0.6086 0.4332 0.4793 0.5704 0.2568 0.4547 0.6112
Llama3.1-8B-Instruct Dubey et al. (2024) 0.399 0.4235 0.5122 0.4164 0.483 0.5238 0.2655 0.4499 0.6158
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.222 0.4501 0.2802 0.1908 0.2578 0.3795 0.121 0.3244 0.2532
Qwen2-7B-Instruct Yang et al. (2024) 0 0.2923 0.0004 0.0002 0.4123 0.0004 0 0 0
Yi-1.5-9B 0.2884 0.5461 0.4927 0.205 0.3724 0.4126 0.1064 0.6596 0.4526
Chatglm-9B 0.3666 0.6902 0.4736 0.3514 0.682 0.5 0.1832 0.6506 0.4342
Llama-3.2-1B-Instruct Dubey et al. (2024) 0.0644 0.5055 0.1664 0.0822 0.441 0.1604 0.0714 0.4757 0.1796
MolT5-small 0.2158 0.1052 0.4302 0.2316 0.1011 0.442 0.2214 0.1031 0.4326
MolT5-base 0.2074 0.1051 0.4168 0.1856 0.1073 0.3796 0.2358 0.1054 0.4536
MolT5-large 0.4244 0.1015 0.8156 0.4496 0.1072 0.8678 0.4654 0.119 0.9214
BioT5-base 0.5158 0.1526 1 0.506 0.1597 1 0.5068 0.158 1
OpenMolIns-large (Llama-3.2-1B) 0.2898 0.5951 0.385 0.2644 0.5956 0.3678 0.1996 0.5849 0.349
OpenMolIns-large (Llama-3.1-8B) 0.8054 0.6678 0.872 0.7122 0.6548 0.8514 0.5224 0.6398 0.8802
OpenMolIns-light (Galactica-125M) 0.3202 0.6547 0.6416 0.3508 0.6435 0.6358 0.269 0.6521 0.638
OpenMolIns-small (Galactica-125M) 0.4172 0.642 0.5568 0.3958 0.6452 0.5338 0.2956 0.6385 0.5376
OpenMolIns-medium (Galactica-125M) 0.5904 0.5812 0.789 0.5874 0.5873 0.7384 0.4608 0.5859 0.7768
OpenMolIns-large (Galactica-125M) 0.6454 0.5927 0.8198 0.6388 0.5973 0.8028 0.495 0.5962 0.81
OpenMolIns-xlarge (Galactica-125M) 0.7362 0.5744 0.8902 0.7124 0.5697 0.8612 0.5786 0.5677 0.8626
Table 12: Results on MolOpt. For each task, we highlight the best accuracy and underline the second best accuracy.
Models AtomNum BondNum FunctionalGroup
Accuracy Novelty Validity Accuracy Novelty Validity Accuracy Novelty Validity
GPT-4o Achiam et al. (2023) 0.1998 0.6703 0.5852 0.065 0.6336 0.8564 0.233 0.6513 0.859
GPT-4-turbo Achiam et al. (2023) 0.1702 0.6991 0.4904 0.0774 0.6301 0.9068 0.218 0.6605 0.8778
GPT-3.5-turbo Achiam et al. (2023) 0.107 0.5054 0.6947 0.0518 0.6871 0.5522 0.1136 0.6585 0.8686
Claude-3.5 Anthropic (2024b) 0.1928 0.6926 0.6548 0.1058 0.6584 0.886 0.2364 0.6582 0.8892
Claude-3 Anthropic (2024a) 0.1044 0.6833 0.591 0.1042 0.6598 0.8696 0.1816 0.9158 0.6644
Gemini-1.5-pro Deepmind (2024) 0.1742 0.6902 0.6774 0.0708 0.6522 0.8688 0.2486 0.6673 0.924
Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.1404 0.6675 0.5474 0.067 0.6478 0.7378 0.1752 0.6576 0.765
Llama3-8B-Instruct Dubey et al. (2024) 0.0242 0.6649 0.3812 0.026 0.6303 0.57 0.0848 0.6167 0.7216
Llama3.1-8B-Instruct Dubey et al. (2024) 0.0228 0.702 0.3862 0.0395 0.6541 0.6387 0.13 0.6274 0.6905
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.0078 0.6732 0.2986 0.0102 0.6309 0.4524 0.0048 0.6012 0.402
Qwen2-7B-Instruct Yang et al. (2024) 0.011 0.9061 0.2622 0.001 0.8645 0.0796 0.0022 0.8601 0.0622
Yi-1.5-9B Young et al. (2024) 0.0392 0.6848 0.617 0.0208 0.6407 0.7072 0.0126 0.6945 0.6521
Chatglm-9B GLM et al. (2024) 0.0002 0.7483 0.2131 0.0254 0.7189 0.4682 0 0.6908 0.5926
Llama-3.2-1B-Instruct Dubey et al. (2024) 0.004 0.6807 0.185 0.008 0.7465 0.2226 0.0008 0.7461 0.2818
MolT5-small Edwards et al. (2022) 0.0006 0.6586 0.661 0.0064 0.598 0.6202 0.0114 0.5287 0.8354
MolT5-base Edwards et al. (2022) 0.0008 0.6868 0.756 0.007 0.6509 0.8422 0.013 0.5464 0.8382
MolT5-large Edwards et al. (2022) 0.015 0.7103 0.8412 0.0118 0.5611 0.8916 0.0382 0.6088 0.9406
BioT5-base Pei et al. (2023) 0.0118 0.8353 0.995 0.0078 0.6667 0.9992 0.0476 0.6792 0.9998
OpenMolIns-large (LLama-3.2-1B) 0.0144 0.649 0.5616 0.035 0.615 0.6186 0.0252 0.6373 0.4412
OpenMolIns-large (LLama-3.1-8B) 0.0136 0.6634 0.7582 0.0544 0.6614 0.7456 0.1344 0.6396 0.6435
OpenMolIns-light (Galactica-125M) 0.0044 0.6054 0.793 0.0216 0.5724 0.7596 0.0244 0.5756 0.8442
OpenMolIns-small (Galactica-125M) 0.0146 0.6568 0.8424 0.053 0.6365 0.7926 0.057 0.5954 0.8874
OpenMolIns-medium (Galactica-125M) 0.0294 0.6553 0.8698 0.0622 0.6473 0.7474 0.0882 0.6091 0.8932
OpenMolIns-large (Galactica-125M) 0.0464 0.6729 0.9116 0.0716 0.6695 0.7374 0.0996 0.6276 0.8966
OpenMolIns-xlarge (Galactica-125M) 0.1862 0.6899 0.9308 0.1656 0.6887 0.7952 0.2006 0.6445 0.9162
Table 13: Results on MolCustom. For each task, we highlight the best accuracy and underline the second best accuracy.

### v2

@@
 TOMG-Bench encompasses a dataset of three major tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom).
-Each task further contains three subtasks, with each subtask comprising 5,000 test samples.
-Given the inherent complexity of open molecule generation, we have also developed an automated evaluation system that helps measure both the quality and the accuracy of the generated molecules.
-Our comprehensive benchmarking of 25 LLMs reveals the current limitations and potential areas for improvement in text-guided molecule discovery.
-Furthermore, with the assistance of OpenMolIns, a specialized instruction tuning dataset proposed for solving challenges raised by TOMG-Bench, Llama3.1-8B could outperform all the open-source general LLMs, even surpassing GPT-3.5-turbo by 46.5% on TOMG-Bench.
+Each major task further contains three subtasks, while each subtask comprises 5,000 test samples.
+Given the inherent complexity of open molecule generation evaluation, we also developed an automated evaluation system that helps measure both the quality and the accuracy of the generated molecules.
+Our comprehensive benchmarking of 25 LLMs reveals the current limitations as well as potential areas for improvement in text-guided molecule discovery.
+Furthermore, we propose OpenMolIns, a specialized instruction tuning dataset established for solving challenges raised by TOMG-Bench.
+Fine-tuned on OpenMolIns, Llama3.1-8B could outperform all the open-source general LLMs, even surpassing GPT-3.5-turbo by 46.5% on TOMG-Bench.
 Our codes and datasets are available through https://github.com/phenixace/TOMG-Bench.
 1 Introduction
-Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals Keiser et al.
-(2010) to materials science Higuchi et al.
-(2023).
-Normally, molecule discovery is a trial and error process Ekins (2024), which requires repetitive experimentation and data analysis Mattern and Grosser (2023).
-Due to the inefficiency of the traditional techniques, it usually takes more than 10 years to bring a new drug candidate into the market Lee et al.
-(2018).
-With the development of machine learning techniques and the advent of Graph Neural Networks (GNNs) Wang et al.
-(2023), there has been a significant step forward.
+Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023).
+Normally, molecule discovery is a trial-and-error process (Ekins, 2024), which requires repetitive experimentation and data analysis (Mattern & Grosser, 2023).
+Due to the inefficiency of the traditional techniques, it usually takes more than 10 years to bring a new drug candidate into the market (Lee et al., 2018).
+With the development of machine learning techniques and the advent of Graph Neural Networks (GNNs) (Wang et al., 2023), there has been a significant step forward.
 As molecules could be represented as graphs, GNN-based methods can capture the structural patterns of the molecule and make accurate predictions.
-With the assistance of GNNs, researchers could analyse the properties of molecules Cai et al.
-(2022) and generate new molecule candidates Jin et al.
-(2018).
+With the assistance of GNNs, researchers could analyse the properties of molecules (Cai et al., 2022) and generate new molecule candidates (Jin et al., 2018).
 However, challenges still exist.
-GNN-based methods struggle to generalize to different tasks Chen et al.
-(2024), necessitating costly data collection and preparation for different downstream tasks.
-Moreover, these methods are constrained in their capacity to generate molecules with specific, customized properties, limiting their flexibility in molecular design Li et al.
-(2024c).
+GNN-based methods struggle to generalize to different tasks (Chen et al., 2024), necessitating costly data collection and preparation for different downstream tasks.
+Moreover, these methods are constrained in their capacity to generate molecules with specific, customized properties, limiting their flexibility in molecular design (Li et al., 2024c).
+In contrast, Large Language Models (LLMs) have shown their great generalization capability (OpenAI, 2023) and could be easily adapted to different research fields.
+For instance, Cascella et al.
+(2023) utilize ChatGPT for supporting clinical practice and Zhen et al.
+(2024) adopts LLMs as assistants for task planning in the field of Civil Engineering, showing the great potential of LLMs in scientific discovery.
 Figure 1: Comparison of Text-based Targeted Molecule Generation (a) v.s. Text-based Open Molecule Generation (b).
-In Contrast, Large Language Models (LLMs) have shown their great generalization capability Achiam et al.
-(2023) and could be easily adapted to different research fields.
-For instance, Cascella et al.
-(2023) utilizes ChatGPT for supporting clinical practice and Zhen et al.
-(2024) adopts LLMs as assistants for task planning in the field of Civil Engineering, showing the great potential of LLMs in scientific discovery.
-As molecules can be represented as texts by Simplified Molecular Input Line Entry System (SMILES), a linear notation that encapsulates the structure of a chemical compound Weininger (1988), they can be processed and understood by LLMs, bridging the gap between molecules and natural languages.
-With advanced reasoning and in-context learning capabilities, LLMs are particularly adept at generalizing to the molecule domain Li et al.
-(2024b).
+As molecules can be represented as texts by Simplified Molecular Input Line Entry System (SMILES), a linear notation that encapsulates the structure of a chemical compound (Weininger, 1988), they can be processed and understood by LLMs, bridging the gap between molecules and natural languages.
+With advanced reasoning and in-context learning capabilities, LLMs are particularly adept at generalizing to the molecule domain (Li et al., 2024b).
 The generalization capability makes LLMs a viable option in molecule discovery.
-Furthermore, by aligning molecules with textual data, LLMs can serve as powerful assistants to chemists Zhang et al.
-(2024); Li et al.
-(2024d).
+Furthermore, by aligning molecules with textual data, LLMs can serve as powerful assistants to chemists (Zhang et al., 2024; Li et al., 2024d).
 They could help interpret and generate chemical knowledge, suggest modifications to molecular structures, and even predict the properties and behaviours of compounds, which would potentially streamline the molecule discovery process, leading to breakthroughs in diverse research areas.
-While the integration of LLMs into molecule discovery holds immense promise, the process of aligning molecules with textual data is challenging Li et al.
-(2024a).
+While the integration of LLMs into molecule discovery holds immense promise, the process of aligning molecules with textual data is challenging (Li et al., 2024a).
 A significant challenge lies in the availability and diversity of datasets and benchmarks necessary for training and evaluation.
-Although the task of molecule-caption translation Edwards et al.
-(2022) is crucial for bridging the gap between the molecular and textual domains, it still has several limitations that need to be addressed:
-On the one hand, there is a concern about the generalization of the molecule-caption translation task Li et al.
-(2024b).
+Although the molecule-caption translation datasets including ChEBI-20 (Edwards et al., 2022), PubChem324K (Liu et al., 2023) and L+M-24 (Edwards et al., 2024) are crucial for bridging the gap between the molecular and textual domains, they still face several limitations that need to be addressed:
+On the one hand, there is a concern about the generalization of models trained on the molecule-caption translation task (Li et al., 2024b).
 In real-world scenarios, molecule captions that describe molecular structures can be highly ambiguous, with multiple correct interpretations, while the current molecule-caption translation is actually a targeted generation task.
-In this case, these models often struggle to generalize to customized molecules, even for seemingly simple examples Li et al.
-(2024b).
+In this case, these models often struggle to generalize to customized molecules, even for seemingly simple examples (Li et al., 2024b).
 This suggests a fundamental mismatch between the molecular and textual spaces, raising questions about whether this task could truly guide LLMs well.
@@
 TOMG-Bench is designed to evaluate the open-domain generative capabilities of LLMs in the molecular domain through a series of structured instructions for molecule design and operations.
-As shown in Figure 1, different from the previous targeted molecule generation tasks, Text-based Open Molecule Generation does not set a specific target or enables LLMs to generate an exact matched molecule.
-Instead, we adopt chemical toolboxes like RDKit Landrum (2013) to test whether the generation meets the requirements.
+As shown in Figure 1, different from the previous targeted molecule generation tasks, Text-based Open Molecule Generation does not set a specific target or enables LLMs to generate an exactly matched molecule.
+Instead, we design automatic evaluation processes with chemical toolboxes like RDKit (Landrum, 2013) to test whether the generation meets the requirements.
 In other words, there can be multiple correct answers for a single question, and LLMs are only required to generate one of them.
 Notably, TOMG-Bench is meticulously categorized into three primary tasks, i.e., molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom).
-Each category contains three subtasks, and each subtask is composed of 5,000 test samples, providing a comprehensive and robust assessment of whether LLMs truly grasp the molecular space.
+Each category contains three subtasks, and each subtask is composed of 5,000 test samples, providing a comprehensive assessment of whether LLMs truly grasp the molecular space.
 Meanwhile, we also propose a different set of evaluation metrics to evaluate and rank the performance of LLMs, which considers both the accuracy and quality of the generated molecules.
-Moreover, we propose an instruction-tuning dataset, OpenMolIns, by extracting and reformatting molecules from an existing molecule database.
-OpenMolIns is structured across five distinct data levels (i.e., light, small, medium, large, and extra-large) to tailor different training purposes.
-To encapsulate our contributions, they are primarily threefold:
-  • Firstly, we are the first to propose the task of Text-based Open Molecule Generation, along with a new set of evaluation metrics to benchmark the performance of LLMs in this task.
-  • Secondly, we have benchmarked a suite of 25 LLMs, shedding light on the limitations of the targeted molecule generation.
-  • Thirdly, we introduce an instruction tuning set, OpenMolIns, which helps Llama-3.1-8B-Instruct to outperform all the open-source general LLMs in TOMG-Bench.
+To solve the challenges raised by TOMG-Bench, we propose an instruction-tuning dataset, OpenMolIns, by extracting and reformatting molecules with meticulously designed rules from an existing molecule database.
+OpenMolIns is structured across five distinct data levels (i.e., light, small, medium, large, and xlarge) to study the data scaling law of LLMs in the text-based open molecule generation task.
+Our contributions are primarily threefold:
+  • 1.
+    TOMG-Bench introduces the task of Text-based Open Molecule Generation, which consists of three major tasks and their nine subtasks in total to synthetically benchmark the performance of LLMs in this task, demonstrating the limitations of the targeted molecule generation datasets like ChEBI-20.
+  • 2.
+    We benchmarked a suite of 25 LLMs, providing a brand-new perspective for the evaluation of LLMs.
+  • 3.
+    We introduce OpenMolIns, an instruction tuning dataset, enabling Llama-3.1-8B-Instruct to outperform all the open-source general LLMs in TOMG-Bench.
 2 Related Work
-In this section, we briefly review related work about developing Artificial Intelligences (AIs) in Molecule Discovery and, more specifically, in text-based molecule generation tasks.
 2.1 Development of AIs in Molecule Discovery
-Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design Du et al.
-(2022).
+Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design (Du et al., 2022).
 Thus, integrating artificial intelligence into molecule discovery has marked a transformative shift in the pharmaceutical landscape, significantly enhancing the efficiency and effectiveness of identifying and developing new therapeutic molecules.
-Recent advancements in machine learning (ML), deep learning (DL), and natural language processing (NLP) have enabled AI systems to analyze complex biological and chemical data more effectively than traditional methods Wigh et al.
-(2022); Wu et al.
-(2018); Zhou et al.
-(2023).
-For instance, MolReGPT  Li et al.
-(2024b) leverages large language models (LLMs) like ChatGPT to learn molecule SMILES strings representation for molecule-caption translation tasks.
-Moreover, existing studies have explored advanced methods that utilize various AI techniques to further enhance molecule discovery processes, including Convolutional Neural Networks (CNNs) Peng and Zhao (2019); Le et al.
-(2019), Recurrent Neural Networks (RNNs) Grisoni et al.
-(2020); Popova et al.
-(2019), Graph Neural Networks (GNNs) Wang et al.
-(2023); Sun et al.
-(2022), and Transformer-based networks Xia et al.
-(2023); Balaji et al.
-(2023); Edwards et al.
-(2022).
+Recent advancements in machine learning (ML), deep learning (DL), and natural language processing (NLP) have enabled AI systems to analyze complex biological and chemical data more effectively than traditional methods (Wigh et al., 2022; Wu et al., 2018; Zhou et al., 2023).
+For instance, MolReGPT  (Li et al., 2024b) leverages LLMs like ChatGPT to learn molecule SMILES strings representation for molecule-caption translation tasks.
+Moreover, existing studies have explored advanced methods that utilize various AI techniques to further enhance molecule discovery processes, including Convolutional Neural Networks (CNNs) (Peng & Zhao, 2019; Le et al., 2019), Recurrent Neural Networks (RNNs) (Grisoni et al., 2020; Popova et al., 2019), Graph Neural Networks (GNNs) (Wang et al., 2023; Sun et al., 2022), and Transformer-based networks (Xia et al., 2023; Balaji et al., 2023; Edwards et al., 2022).
 2.2 Text-based Molecule Generation
-Text-based Molecule Generation (Text2Mol) Edwards et al.
-(2021) has recently emerged as a transformative approach to molecule discovery.
-This task centres on retrieving molecules using natural language descriptions as search queries, requiring the creation of paired datasets of molecules and their corresponding textual representations.
-This enables the learning of a shared semantic embedding space for efficient retrieval.
-Early approaches leveraged transformer-based models like MolT5 Edwards et al.
-(2022), employing self-supervised learning on large datasets to generate high-quality Simplified Molecular Input Line Entry System (SMILES) strings from textual inputs.
-Subsequent advancements, such as KV-PLM Zeng et al.
-(2022), MoMu Su et al.
-(2022), and BioT5 Pei et al.
-(2023), integrated molecular graphs and biochemical text to improve both understanding and generation capabilities.
-3D-MoLM Li et al.
-(2024e) further enhanced this by incorporating spatial configurations, leading to more accurate and geometrically valid molecular representations.
-The application of large language models (LLMs) like MolReGPT Li et al.
-(2024b) and ICMA Li et al.
-(2024a) as in-context learners has also shown significant promise.
+Text-based Molecule Generation (Edwards et al., 2021) has recently emerged as a transformative task to molecule discovery.
+Previously, this task centres on generating targeted molecules based on natural language descriptions, requiring the creation of paired datasets of molecules and their corresponding textual representations.
+Early approaches leveraged transformer-based models like MolT5 (Edwards et al., 2022), employing self-supervised learning on large datasets to generate high-quality SMILES strings from textual inputs.
+Subsequent advancements, such as KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023), integrated molecular graphs or biochemical text to improve both understanding and generation capabilities.
+3D-MoLM (Li et al., 2024e) further enhanced this by incorporating spatial configurations, leading to more accurate and geometrically valid molecular representations.
+Meanwhile, the application of LLMs like MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) as in-context learners has also shown significant promise.
 These models demonstrate the ability of LLMs to adaptively generate molecules by retrieving and leveraging relevant examples from the provided context.
-Most recently, MolReFlect Li et al.
-(2024c) underscored the importance of fine-grained alignment between molecular structures and their textual descriptions, utilizing a teacher-student training paradigm to capture these nuanced relationships effectively.
-Unlike this targeted generation task, in this paper, we propose a Text-based Open Molecule Generation task to enable LLMs to generate an exactly matched molecule rather than set a specific target.
+Most recently, MolReFlect (Li et al., 2024c) underscored the importance of fine-grained alignment between molecular structures and their textual descriptions, utilizing a teacher-student training paradigm to capture these nuanced relationships effectively.
 3 TOMG-Bench
@@
 To ensure the integrity and effectiveness of TOMG-Bench, we have developed a robust set of evaluation metrics for different tasks.
-Additionally, we have created OpenMolIns, an instruction-tuning dataset aimed at enhancing the performance and adaptability of LLMs to the challenges presented by this benchmark.
-Figure 2: Data construction workflow and evaluation process of TOMG-Bench.
+Additionally, we created OpenMolIns, an instruction-tuning dataset aimed at enhancing the performance and adaptability of LLMs to the challenges presented by this benchmark.
+Figure 2: Task illustrations of TOMG-Bench. Both molecules to the right of the arrow are acceptable solutions.
 3.1 Dataset Categorization
 The categorization of the TOMG-Bench dataset initially considers the inherent characteristics of molecule SMILES representation and the role of LLMs serving as the chemist’s assistant, namely helping the chemists to edit, optimize and customize molecules as they want.
-Following the difficulty of the tasks, we demonstrate the content of the three basic tasks as well as their corresponding subtasks:
+As shown in Figure 2, following the difficulty of the tasks, we demonstrate the three basic tasks as well as their corresponding subtasks:
 MolEdit emerges as the most straightforward task among the three domains, as an existing molecule is already provided, and LLMs are only required to make modifications to it, which tests the molecular structure knowledge of LLMs.
@@
 MolOpt challenges LLMs to not only edit molecules but also to discern whether the modification will steer the molecule towards a desired optimization target.
-To assess this capability, we concentrate on three pivotal properties that are vital for molecule discovery: LogP (Octanol-water partition coefficient, a metric of lipophilicity), MR (molecular refractivity, a proxy for the molar refractive index), and QED (Quantitative Estimate of Druglikeness, an assessment of drug-like characteristics).
+To assess this capability, we concentrate on three pivotal properties that are vital for molecule discovery and can be effectively validated: LogP (Octanol-water partition coefficient, a metric of lipophilicity), MR (molecular refractivity, a proxy for the molar refractive index), and QED (Quantitative Estimate of Druglikeness, an assessment of drug-like characteristics).
 These metrics offer critical information about the potential pharmacological attributes of the molecule, which could help chemists filter molecules as viable drug candidates.
@@
 They demand that LLMs have a sophisticated understanding of molecular syntax to precisely generate molecules that meet the complex criteria set forth.
+Compared to previous text-based target molecule generation datasets, the task design in TOMG-Bench is rather naive, but focuses more on the details of the molecular substructure, which puts even more rigorous requirements on both the molecular understanding and generation capabilities of LLMs.
+Furthermore, the accurate understanding on molecule substructures and editing operations also ensures the explainability and trustworthiness of generated molecules.
 3.2 Dataset Construction
 Previously, molecule-related datasets and tasks have been hindered by the scarcity of human annotations.
-For instance, ChEBI-20  Edwards et al.
-(2022), a dataset for the molecule-caption translation task, contains only 33,000 samples for training, whereas image captioning datasets like MS COCO  Chen et al.
-(2015) have over 1,500,000 annotated captions on more than 330,000 images.
+For instance, ChEBI-20  (Edwards et al., 2022), a dataset for the molecule-caption translation task, contains only 26,407 samples for training.
+In contrast, datasets for the text-to-image generation task in the computer vision field like MS COCO  (Chen et al., 2015) have over 1,500,000 annotated captions on more than 330,000 images.
 Meanwhile, the annotation of molecules demands expertise from chemists and can sometimes require wet lab experiments, which are both expensive and time-consuming.
 In contrast, TOMG-Bench, as an open-domain generation task, does not rely on human annotations for construction.
-Instead, we focus on basic molecule structural properties and basic molecule operations that could be validated by chemical toolboxes to construct our dataset.
-For MolCustom, we randomly generate 5,000 prompts as requests for each subtask, requiring different numbers and collections of atoms, bonds, and functional groups.
+Instead, we focus on basic molecule structural properties and molecule operations that could be validated by chemical toolboxes, which makes it easy to increase the volume of data, thereby resolving data hunger in chemical corpora.
+In TOMG-Bench, three major tasks consist of nine subtasks in total. Each subtask is allocated 5,000 test samples. Figure 4 illustrates the data construction process.
 For MolEdit and MolOpt, we sample molecules from a specific molecule database.
-Specifically, we select two molecule databases for this work: Zinc-250K Sterling and Irwin (2015), and PubChem  Kim et al.
-(2019).
-Zinc-250K has 250,000 molecules, which is smaller than PubChem, which has 10 million molecules.
-To facilitate the fast calculation of metrics mentioned in Section 3.5, we choose Zinc-250K for sampling the test molecules in TOMG-Bench.
-Each subtask is allocated 5,000 test samples.
-After sampling, we utilize RDKit Landrum (2013), a molecular informatics toolbox, to collect basic molecule statistics.
-There are functions available to calculate the required characteristics, especially the structural patterns and chemical properties like LogP, MR, and QED values, which can then be integrated into our pre-defined task prompts.
-Further details will be provided in the Appendix A.
+We mainly consider two molecule databases: Zinc-250K (Sterling & Irwin, 2015), and PubChem  (Kim et al., 2019).
+Zinc-250K has 250,000 molecules, while PubChem has about 10 million molecules.
+Therefore, the data size of Zinc-250K is more suitable for us to build the test sets, while PubChem could be further sampled to generate the instruction tuning dataset.
+After sampling, we utilize RDKit (Landrum, 2013), a molecular informatics toolbox, to collect basic molecule statistics, including the structural patterns and chemical properties like LogP, MR, and QED values, which can then be integrated into our pre-defined task prompts.
+For the MolCustom task, we randomly generate 5,000 instructions as requests for each subtask, requiring different numbers and collections of atoms, bonds, and functional groups.
+Further details are provided in the Appendix D.
 3.3 Metric Design
 The evaluation of the TOMG-Bench is facilitated through a set of carefully designed metrics tailored to the specific tasks within the benchmark.
-For the MolCustom task, which includes subtasks such as AtomNum, BondNum, and FunctionalGroup, the following metrics are employed:
-  • Accuracy: This metric measures the proportion of generated molecules that successfully meet the criteria.
-A high accuracy indicates that the LLM is effectively generating molecules according to the request.
-  • Novelty: Novelty is a critical metric for MolCustom tasks, as it assesses the ability of LLMs to discover new molecular structures.
-This is particularly important for potential drug or material candidates, where innovation can lead to significant advancements.
+Task (Weighted) Success Rate Similarity Novelty Validity
+MolEdit
+MolCustom
+MolOpt
+Table 1: All used metrics are listed here. ’✓’ means that a metric is calculated on a task.
+3.3.1 MolEdit & MolOpt Evaluation
+For the MolEdit and MolOpt tasks, as shown in Table 1, the following metrics are adopted:
+Success Rate: We have developed automated testing processes with RDKit, which compare the generated molecule with the original molecule and check if the number of the specified functional group is correct after modification (for MolEdit task) and if the specified chemical property is optimized as required (for MolOpt task).
+Details of the automated testing process are shown in Appendix E.
+The success rate reflects the percentage of molecules that have successfully passed the rigorous testing processes.
+However, as it is impossible to track every change between the generated molecules and the original molecule (See Appendix E for detailed explanation), successfully passing the test processes does not necessarily mean that the generated molecule is modified from the original molecule.
+Similarity: In this case, we need to evaluate the similarity between the original molecule and the generated molecule.
+When removing a specified functional group, the modification should follow the minimal change principle.
+Therefore, a high similarity indicates that the generated molecule can be modified from the original molecule via as few as possible steps, instead of transforming the molecule into an entirely new one or inadvertently removing additional components.
+For implementation, we calculate the Tanimoto Similarity between the generated molecule and the original molecule by transforming them into the Morgan Fingerprints (Butina, 1999), which can be described by the following equation:
+$\delta(m^{g},m^{o})=\frac{\left|fp_{m^{g}}\cap fp_{m^{o}}\right|}{\left|fp_{m^{g}}\cup fp_{m^{o}}\right|},$ (1)
+where $m^{g}$ denotes the generated molecule, $m^{o}$ is the original molecule, and $\delta(m^{g},m^{o})$ is the Tanimoto Similarity between the generated molecule and the original molecule, while $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints.
+$\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
+Then, the similarity score for a subtask $t$ in MolEdit and MolOpt, denoted as $\delta_{t}$, is determined by taking the average of the similarity scores from all individual test samples.
 Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax.
-For the MolEdit and MolOpt tasks, the following metrics are adopted:
-  • Accuracy: Similar to the MolCustom category, accuracy here measures the proportion of edited or optimized molecules that could pass the test process.
-  • Similarity: Similarity between the original molecule and the edited molecule is an important point to concern.
-Since there are multiple valid solutions, only considering the accuracy is not sufficient for verification.
-For implementation, a high similarity indicates that the generated molecule can be modified from the original molecule via as few as possible steps.
-For instance, when removing a specific functional group, the edited molecule should still be similar to the original molecule.
-This process should not transform the molecule into an entirely new one nor inadvertently remove additional components.
-  • Validity: Similarly, validity is still a crucial metric to confirm that the edited or optimized molecules adhere to the principles of the molecule textual representation.
-Notably, the calculation of novelty and similarity metrics only considers valid molecules.
-Specifically, for the novelty metric, we assess the similarity between the generated molecule and those within the Zinc-250K database.
-To comprehensively evaluate the average performance of LLMs on TOMG-Bench, we introduce a weighted average accuracy to rank the performance of LLMs.
-Considering that the novelty scores for MolCustom and similarity scores for MolEdit and MolOpt are also crucial metrics for evaluating the performance, especially the similarity scores, which help identify correct molecule editing operations.
-In this case, we adopt the novelty and similarity scores as the weights, and the weighted average accuracy can be computed as follows:
-$\bar{w\!Acc}\!=\!\frac{1}{9}\!(\sum_{t\in\!\!M\!o\!l\!C\!u\!s\!t\!o\!m\!}\!(\!n_{t}*Acc_{t}\!)\!+\sum_{t\in\!\{\!M\!o\!l\!E\!d\!i\!t,\!M\!o\!l\!O\!p\!t\!\}}\!(\!s_{t}*Acc_{t}\!)),$ (1)
-where $n_{t}$ represents the novelty score for the MolCustom tasks and $s_{t}$ is the similarity score for the MolEdit and MolOpt tasks.
-$Acc_{t}$ is the accuracy for each subtask $t$.
-This weighted average accuracy, $\bar{wAcc}$, provides a balanced measure of performance that considers both the accuracy and quality of the generation.
+We use RDKit to parse the SMILES strings.
+Only the molecules that successfully pass the parser are considered valid molecules.
+A higher validity means that the model is more familiar with the grammar of molecule textual representations.
+3.3.2 MolCustom Evaluation
+For the MolCustom task, the metrics adopted are also shown in Table 1.
+As LLMs are required to generate the molecule from scratch, we employ a different testing process for the evaluation of success rate and introduce the novelty score instead of similarity:
+Success Rate: For different subtasks in MolCustom, we have designed different testing processes by the chemical toolbox.
+We calculate the number of the specified atoms, functional groups, or bonds for the three subtasks to check if the generation meets the requirements.
+Detailed testing process of MolCustom could also be found in Appendix E.
+A high success rate indicates that the LLM is effectively generating molecules according to the request.
+Novelty: However, since there may be numerous valid solutions, fulfilling the basic requirements alone does not guarantee the quality of generation.
+To enable the exploration of new molecule structures, we employ novelty as an indication of the quality, which compares the similarity between the newly generated molecules and existing ones in the database.
+The novelty score is particularly important for discovering potential drug or material candidates, where innovation can lead to significant advancements.
+For implementation, we choose the Zinc250K as the database and calculate the average Tanimoto Similarity between the generated molecules and the existing molecules.
+The novelty $n$ for the generated molecule $m^{g}$ can be calculated as follow:
+$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc250K}\delta(m^{g},m^{k})}{\left|Zinc250K\right|},$ (2)
+Then, the novelty score forthe subtask $t$ in MolCustom, denoted as $n_{t}$, is determined by taking the average of the novelty scores from all individual test samples.
+Validity: Similarly, validity is also an important metric for the MolCustom task.
+3.3.3 Average Weighted Success Rate
+To comprehensively evaluate the average performance of LLMs on TOMG-Bench, we introduce an average weighted success rate to rank the performance of LLMs.
+Considering that the success rate alone can not ensure the accurate generation for MolEdit and MolOpt tasks and the novelty score for MolCustom is also a critical metric for the quality of molecule generation, we adopt similarity and novelty scores as weights of the success rate:
+$W\!S\!R_{t}=\begin{cases}n_{t}\times S\!R_{t},&t\in\{MolCustom\}\\ \delta_{t}\times S\!R_{t},&t\in\{MolEdit,MolOpt\}\end{cases},$ (3)
+where $W\!S\!R_{t}$ denotes the weighted success rate for a subtask $t$, while $\delta_{t}$ is the similarity score for the MolEdit and MolOpt tasks, $n_{t}$ represents the novelty score for the MolCustom tasks, and $S\!R_{t}$ is the corresponding success rate of the subtask.
+Then, the average weighted success rate $\overline{W\!S\!R}$ could evaluate the synthetic performance of LLMs among all the nine subtasks:
+$\overline{W\!S\!R}=\frac{1}{9}\sum_{t}W\!S\!R_{t},$ (4)
+where the average weighted success rate $\overline{W\!S\!R}$ provides a balanced measure of performance that considers both the success rate and quality of the generation.
 3.4 OpenMolIns: Instruction Tuning Dataset
@@
 This instruction tuning dataset is meticulously designed to ensure that the molecules it contains do not overlap with those in the Zinc-250K dataset to promote the generation of more novel molecular structures and avoid any potential data leakage that could compromise the integrity of the model performance.
-We collect the instruction tuning dataset by introducing the samples of the nine subtasks in equal amounts. We still apply the RDKit toolbox to construct the instruction tuning dataset.
+We collect the instruction tuning dataset by constructing the samples of the nine subtasks in equal amounts.
+The RDKit toolbox is applied to construct the instruction tuning dataset.
+Details of the construction process and the pre-defined prompt templates are shown in Appendix D.
+For the MolEdit and MolOpt domains, we randomly select a functional group from the original molecule for addition or removal.
+Then, similarly, we wrap the original molecule and the edited molecule with a random pre-defined prompt to construct the training sample for MolEdit.
+In the case of MolOpt, we also evaluate the direction of the desired property changes by calculating the properties to determine whether the operation improves or decreases the property value.
+Then, the original molecule, the optimized molecule, and the optimization direction are altogether wrapped by a pre-defined prompt.
 In the MolCustom domain, constructing the instruction tuning samples is rather straightforward.
@@
 Then, we wrap these statistics with a random pre-defined prompt to construct the training sample.
-This method allows the generated molecules to better fit the distribution of molecular space.
-For the MolEdit and MolOpt domains, we randomly select a functional group from the original molecule for addition or removal, utilizing the RDKit toolbox to execute these operations.
-Then, similarly, we wrap the original molecule and the edited molecule with a random pre-defined prompt to construct the training sample for MolEdit.
-In the case of MolOpt, we also evaluate the direction of the desired property changes by the functions in RDKit to determine whether the operation improves or decreases the property value, which is then altogether wrapped by the pre-defined prompts.
-To investigate the impact of data scales on the performance of LLMs, we have established five distinct data levels tailored for different training purposes: light, small, medium, large, and xlarge, as illustrated in Section 3.5.
-Each level represents a different quantity of data, shown in Table 1, allowing us to analyze how the amount of training data influences the model’s ability to learn and generate or edit molecules effectively.
-Item Data Size
-TOMG-Bench
-subtask 5,000
-OpenMolIns
-light 4,500
-small 18,000
-medium 45,000
-large 90,000
-xlarge 1,200,000
+This method allows the generated molecules to better fit the real distribution of molecular space.
+To investigate the impact of data scales on the performance of LLMs, we have established five distinct data levels: light, small, medium, large, and xlarge, as illustrated in Section 3.5.
+Each level represents a different quantity of data, shown in Table 2, allowing researchers to analyze how the amount of training data influences the model’s ability to generate and edit molecules effectively.
+Dataset TOMG-Bench OpenMolIns
+item each subtask total light small medium large xlarge
+Data Size 5,000 45, 000 4,500 18,000 45,000 90,000 1,200,000
 Table 2: Statisics of TOMG-Bench and OpenMolIns.
 3.5 Statistics
-In this section, we outline the basic statistics of TOMG-Bench along with the OpenMolIns dataset.
-Table 1 shows the details of the data size in TOMG-Bench as well as the OpenMolIns datset.
+In this section, we outline the data sizes of TOMG-Bench along with the OpenMolIns dataset, which are shown in Table 2.
 For TOMG-Bench, we have three main tasks with nine subtasks in total, where each subtask contains 5,000 test samples.
-More details can be found in Appendix A.
-For OpenMolIns, we have five distinct data scales: light, small, medium, large, and xlarge, ranging from 4,500 to 1,200,000 examples, which helps us investigate the data scaling law of applying LLMs to the Text-based Open Molecule Generation task.
-Notably, the nine subtasks within the TOMG-Bench are uniformly distributed in the OpenMolIns dataset.
+OpenMolIns has five distinct data scales: light, small, medium, large, and xlarge, ranging from 4,500 to 1,200,000 examples, which helps us investigate the data scaling law of applying LLMs to the text-based open molecule generation task.
+The nine subtasks within the TOMG-Bench are uniformly distributed in OpenMolIns.
 4 Experiments
-In this section, we present the experiment setup and results. Then, we illustrate our findings based on the observations.
-4.1 Setup
-4.1.1 Models
+4.1 Models
 The models benchmarked are categorized into four groups: proprietary models, open-source general LLMs, open-source ChEBI-20 fine-tuned LLMs, and OpenMolIns fine-tuned LLMs.
@@
 This category includes LLMs that are only accessible via commercial API services.
-In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo Achiam et al.
-(2023), Claude-3.5 Anthropic (2024b), Claude-3 Anthropic (2024a), and Gemini-1.5-pro Deepmind (2024).
+In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful language understanding and generation capabilities.
+These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
 Open-source General LLMs.
 This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications.
-Specifically, we benchmark Llama-3-70B-Instruct, Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-1B-Instruct Dubey et al.
-(2024), Mistral-7B-Instruct-v0.2 Jiang et al.
-(2023), Qwen2-7B-Instruct Yang et al.
-(2024), yi-1.5-9B Young et al.
-(2024), and chatglm-9B GLM et al.
-(2024).
+Specifically, we benchmark Llama-3-70B-Instruct, Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), and chatglm-9B (GLM et al., 2024).
+These LLMs are top-tier open source general LLMs within our computation budget.
 Open-source ChEBI-20 Fine-tuned LLMs.
 LLMs fine-tuned on the ChEBI-20 dataset can grasp some extent of text-based molecule generation capability.
-In this case, our experiments also cover LLMs like MolT5-small, MolT5-base, MolT5-large Edwards et al.
-(2022), and BioT5-base Pei et al.
-(2023).
-OpenMolIns Fine-tuned LLMs We further adopt LLMs like Galactica-125M Taylor et al.
-(2022), Llama3.2-1B-Instruct, and Llama-3.1-8B-Instruct on OpenMolIns dataset for instruction tuning.
+In this case, our experiments also cover LLMs like MolT5-small, MolT5-base, MolT5-large (Edwards et al., 2022), and BioT5-base (Pei et al., 2023).
+These models have been the state-of-art models in the molecule-caption translation task.
+OpenMolIns Fine-tuned LLMs We further adopt LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama-3.1-8B-Instruct on OpenMolIns dataset for instruction tuning.
 We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M.
-4.1.2 Implementation Details
-We implement various scripts to facilitate the testing of the aforementioned models.
-For proprietary models, we adopt the OpenAI API  [^1] https://platform.openai.com/docs/ framework.
-For open-source general LLMs, we utilize both the VLLM  [^2] https://docs.vllm.ai framework and the OpenAI framework.
-For the remaining LLMs, we adopt the Hugging Face transformers library  [^3] https://huggingface.co/docs/transformers/ for inference.
-Detailed hyper-parameters are demonstrated in Appendix B.
-Furthermore, it is important to note that BioT5 is designed to use SELFIES as input instead of SMILES. Consequently, we convert the molecule SMILES strings into SELFIES format on BioT5.
-4.2 Results
+We select Galatica-125M because it has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023).
+Meanwhile, a small size model can also help us study the data scaling law within reasonable budget.
+Besides, Llama3.2-1B-Instruct and Llama-3.1-8B-Instruct are selected due to their advanced capabilities and similar architectures.
 Figure 3: The performance of models benchmarked in TOMG-Bench.
 In TOMG-Bench, LLMs are divided into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs.
-Models whose parameters are known are plotted as dots, while models of unknown parameters are denoted as horizontal lines.
+Models of unknown parameters are denoted as horizontal lines.
+4.2 Findings
 Figure 3 presents the performances and model sizes of different models benchmarked on TOMG-Bench, as well as the instruction-tuning performance of Galactica-125M on the five distinct data levels of OpenMolIns.
-More precise details are further illustrated in Appendix C.
-4.3 Findings
+More precise details for each subtask are further illustrated in Appendix F.
 Based on the above results, we observe the following key findings:
-Text-based open molecule generation is challenging for LLMs.
-As illustrated in Figure 3, we have calculated the weighted average accuracy across all the nine subtasks.
-Among the LLMs benchmarked, Claude-3.5 stands out as the top performer, achieving a weighted average accuracy of 35.92%.
-Gemini-1.5-pro follows closely with a weighted average accuracy of 34.80%.
-These results underscore the considerable scope for improvement, even among the most advanced proprietary LLMs.
-It is also worth noting that while the most advanced LLMs like Claude-3.5 and GPT-4o exhibit relatively strong performance in the MolEdit and MolOpt tasks, the more challenging MolCustom task still remains a challenge.
-In MolCustom tasks, no LLM has managed to achieve an accuracy exceeding 25% for a single subtask.
-This observation indicates that the generation of molecules from scratch demands a deep understanding of the molecular structural space, an area where current models are still striving to make significant strides.
+Text-based open molecule generation is both promising and challenging for LLMs.
+As illustrated in Figure 3, we have calculated the average weighted success rate across all the nine subtasks.
+Among the LLMs benchmarked, Claude-3.5 stands out as the top performer, achieving an average weighted success rate of 35.92%.
+Gemini-1.5-pro follows closely with an average weighted success rate of 34.80%.
+Meanwhile, fine-tuned on OpenMolIns-large, Llama-3.1-8B-Instruct achieves an average weighted success rate of 27.22%, outperforming all the open-source general LLMs.
+However, these results still underscore the considerable scope for improvement, even among the most advanced proprietary LLMs.
+What’s more, despite LLMs like Claude-3.5 and GPT-4o exhibit relatively strong performance in the MolEdit and MolOpt tasks, the more challenging MolCustom task still remains a challenge.
+In MolCustom tasks, no LLM has managed to achieve a success rate exceeding 25% for a single subtask.
+Meanwhile, as shown in Figure 5, among the three subtasks of MolCustom, BondNum is the most difficult task, as it requires LLMs to understand the connections or edges in the molecule graph merely from the textual representation.
+These observations indicate that the generation of molecules from scratch demands a deep understanding of the molecular structures, an area where current models are still striving to make significant strides.
 Most powerful open-source general LLMs can already outperform GPT-3.5-turbo.
-In the TOMG-Bench, Llama-3-70B-Instruct has achieved an impressive weighted average accuracy of 23.93%, notably outperforming GPT-3.5-turbo, which scored 18.58%.
-Despite previously lagging behind proprietary models, open-source general LLMs have rapidly bridged the gap.
-The evolution of the Llama series, in particular, has been remarkable, finally surpassing GPT-3.5-turbo and demonstrating the fast development of open-source general LLMs.
-More powerful LLMs inherit a better performance in TOMG-Bench.
+In TOMG-Bench, Llama-3-70B-Instruct has achieved an impressive average weighted success rate of 23.93%, notably outperforming GPT-3.5-turbo, which scored 18.58%.
+Compared to the proprietary models, open-source general LLMs seem to have been lack of chemical corpora for the pre-training process.
+However, recent development of open-source general LLMs have rapidly bridged the gap.
+The evolution of the Llama series, in particular, has been remarkable, finally surpassing GPT-3.5-turbo and showing the increasing diversity of pre-training corpora.
+More powerful LLMs usually inherit a better performance in TOMG-Bench.
 Across all the LLMs we benchmarked, a clear trend emerged: the more powerful the LLM is, the higher performance it can achieve on the TOMG-Bench.
 For instance, the GPT series has consistently demonstrated improved performance with each new iteration from GPT-3.5-turbo to GPT-4o.
-Similarly, within the Llama-3 series, we could also observe that larger models tend to achieve superior results on the TOMG-Bench.
-These findings underscore a strong correlation between an LLM’s capabilities and its performance in our benchmark.
+Similarly, within the Llama-3 series, we could also observe that larger and more powerful models tend to achieve superior results on the TOMG-Bench.
+These findings underscore a strong correlation between an LLM’s capabilities and its performance in text-based open molecule generation.
 However, we encountered an unexpected anomaly with certain open-source LLMs.
@@
 It provides a broader and more diverse assessment that exposes potential blind spots in the development of LLMs.
-ChEBI-20 dataset is insufficient for LLMs to master molecular structures and editing operations.
-The ChEBI-20 dataset and the associated molecule-caption translation task are designed to bridge the gap between molecular structures and textual descriptions.
-Despite this intention, LLMs trained on ChEBI-20 have demonstrated limited effectiveness in our TOMG-Bench benchmark.
-For instance, BioT5-base, which is claimed as the state-of-the-art (SOTA) model for text-based molecule generation on the ChEBI-20 dataset, only achieves a weighted average accuracy of 4.21% on the TOMG-Bench.
-In the MolEdit and MolOpt tasks, these models are unable to execute correct operations on provided molecules, resulting in disappointing similarity scores.
-Similarly, in the MolCustom task, which closely mirrors the text-based molecule generation task, the performance remains unsatisfactory, with no model achieving a score above 5% in a single subtask.
+ChEBI-20 is insufficient for LLMs to master molecular structures and editing operations.
+The ChEBI-20 dataset is designed to bridge the gap between molecules and textual descriptions.
+Despite this intention, LLMs trained on ChEBI-20 have demonstrated limited effectiveness in TOMG-Bench.
+For instance, BioT5-base, which is claimed as the state-of-the-art model for text-based molecule generation on the ChEBI-20 dataset, only achieves an average weighted success rate of 4.21% on the TOMG-Bench.
+In the MolEdit and MolOpt tasks, although these models obtain a high success rate, the generations are actually not modified from the original molecules, leading to poor similarity scores.
+This also shows that the average weighted success rate provides a more accurate reflection of the actual performance.
+Similarly, in the MolCustom task, which closely mirrors the text-based molecule generation task, the performance remains unsatisfactory, with no model achieving a success rate above 5% in a single subtask.
 This performance shortfall highlights a critical limitation of the ChEBI-20 dataset, as it lacks the data quantity and diversity necessary to effectively align molecules with textual descriptions.
-In contrast, TOMG-Bench offers a more comprehensive and intricate evaluation framework for text-to-molecule generation.
-With a larger and more varied set of test examples, TOMG-Bench could robustly assess the capabilities of language models in translating textual descriptions into molecular structures.
-As such, it represents a significant advancement in the evaluation of text-based molecule generation.
-OpenMolIns can enable LLMs to achieve better performance than the most powerful open source general LLMs.
-We have also developed OpenMolIns, an instruction-tuning dataset, to enhance LLMs’ proficiency in the tasks outlined in the TOMG-Bench.
-Across five distinct data scales, we observed a pronounced data scaling law: as the size of the corpus increases, the performance of LLMs also improves.
-In particular, for Galactica-125M, we assessed its capabilities comprehensively on both five distinct data scales.
-As shown in Figure 3, the outcomes were remarkable: Galactica-125M achieved a weighted average score of 25.73% on OpenMolIns-xlarge, surpassing even the 70B Llama-3-70B-Instruct and GPT-3.5-turbo, with only 125 million parameters.
-Meanwhile, the results of Galactica-125M show a clear data scaling law, denoting that LLMs are hungry for more molecule corpora to achieve better performance.
-Notably, OpenMolIns-large has also enabled Llama3.1-8B-Instruct to outperform all the existing open-source general LLMs in TOMG-Bench, showing the effectiveness of the dataset.
+The data scaling law varies among different tasks.
+Across five distinct data scales of OpenMolIns and the experiments on Galactica-125M, we observe a general data scaling law: as the size of the corpus increases, the performance of LLMs also improves.
+As shown in Figure 3 and Table 12, the outcomes were remarkable: Galactica-125M achieved a weighted average score of 25.73% on OpenMolIns-xlarge, surpassing even the 70B Llama-3-70B-Instruct and GPT-3.5-turbo, with only 125 million parameters.
+However, an interesting observation shows that the data scaling law varies among different tasks.
+For instance, in MolCustom tasks, we witnessed a substantial average improvement of 286% when we scaling the data size from large to xlarge, which suggests that, for specific tasks, increasing data size has a beneficial effect on model performance.
+Conversely, in the MolCustom tasks, the performance gains from data scaling were notably modest.
+More concerning, in the MolEdit tasks, we even observed a decline in performance with increased data size, which indicates that the model’s capacity is not solely data-driven but is also constrained by the number of parameters.
+It suggests that for certain subtasks within TOMG-Bench, simply increasing the volume of data may not lead to proportional improvements in performance, and other factors such as model complexity or optimization strategies may need to be considered.
 5 Conclusion
-In this study, we introduce TOMG-Bench, the first benchmark designed to assess the capabilities of Large Language Models (LLMs) in the realm of open-domain molecule generation.
+In this study, we introduce TOMG-Bench, the first benchmark designed to assess the capabilities of LLMs in the realm of open-domain molecule generation.
 Benchmarking 25 LLMs, TOMG-Bench highlights the limitations of existing targeted molecule generation tasks and demonstrates the potential of general LLMs in this domain.
-Additionally, through instruction tuning on our proposed OpenMolIns, LLMs exhibit significant potential on the TOMG-Bench, matching the performance of GPT-3.5-turbo.
-Our contributions not only lie in the development of a novel benchmark for molecular discovery but also provide a diverse indication of the capabilities of LLMs.
-6 Limitations
-Although TOMG-Bench is carefully designed and well-validated through our experiments, we still observe several limitations:
-Prompt Diversity.
-Prompt diversity helps relieve the over-fitting of the instructions.
-While we adopt several different prompt templates and randomly choose from them, we still find the number of prompt templates is not enough to satisfy the prompt diversity.
-Data Distribution.
-In our data construction process, we allocate distributions to atoms, bonds, and functional groups with the aim of making our benchmark more reflective of real-world distributions.
-Nevertheless, the distribution we use is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios accurately.
-This could potentially mask the true performance capabilities of LLMs in these specific tasks.
-7 Acknowledgments
-We thank all the reviewers for their insightful comments. We also thank GitHub Copilot for coding assistance and ChatGLM-4 for polishing the writing.
-Appendix A Data Construction
+Additionally, we propose an instruction tuning dataset, OpenMolIns, to fully explore the potential of LLMs on the TOMG-Bench, enabling Llama-3.1-8B-Instruct to outperform GPT-3.5-turbo.
+Our contributions not only lie in the development of a novel benchmark for LLMs in molecular discovery but also provide a diverse indication of the LLM evaluation.
+Appendix A Comparison with Existing Text-based Molecule Benchmarks or Datasets
+Benchmarks or Datasets Task Type # Training Data # Test Data Public Access
+PCdes(Zeng et al., 2022) Targeted Generation 10,500 3000 $\times$
+ChEBI-20 (Edwards et al., 2021) Targeted Generation 26,407 3,000
+PubChem (Liu et al., 2023) Targeted Generation 12,000 2,000
+Mol-Instructions (description guided molecule design) (Fang et al., 2024) Targeted Generation 297,319 1000
+TOMG-Bench & OpenMolIns Open Generation 1,200,000 45,000
+Table 3: Comparison with existing text-based molecule generation benchmarks or datasets.
+Table 3 shows the comparison between TOMG-Bench & OpenMolIns and the existing text-based molecule generation benchmarks or datasets.
+TOMG-Bench introduces the brand-new open-ended molecule generation task for LLMs, which boosts the diversity of text-based molecule generation task.
+Compared to the existing benchmarks or datasets, OpenMolIns has the largest training set, which could better aligns the molecular space with the natural language space.
+Meanwhile, TOMG-Bench owns three major tasks and nine subtasks with 45,000 test samples in total, whose test data size is also the largest one among all the existing benchmarks and datasets, enabling more accurate evaluation and comprehensive assessment of LLMs’ performance in text-based molecule generation.
+Appendix B Implementation Details
+We implement various scripts to facilitate the testing of the aforementioned models.
+For proprietary models, we adopt the OpenAI API [^1] https://platform.openai.com/docs/ framework.
+For open-source general LLMs, we utilize both the VLLM [^2] https://docs.vllm.ai framework and the OpenAI framework.
+For the remaining LLMs, we adopt the Hugging Face transformers library [^3] https://huggingface.co/docs/ for inference.
+Furthermore, BioT5 is designed to use SELFIES, an alternative string representation of molecules by using special tokens to ensure valid generations (Krenn et al., 2019), as input instead of SMILES.
+Consequently, we convert the molecule SMILES strings into SELFIES format on BioT5.
+Appendix C Hyper Parameters
+In this section, we illustrate the detailed parameters adopted in this work, as shown in Table 4.
+Item Value
+Generation
+temperature 0.75
+top_p 0.85
+num_beams 1
+max_new_tokens 512
+Instruction Tuning
+batchsize 32
+lr 3e-4
+cutoff_len 1024
+Lora Settings
+r 64
+$\alpha$ 128
+dropout 0.1
+Table 4: Hyper-parameters.
+Appendix D Data Construction
 In this section, we introduce the construction details of TOMG-Bench and OpenMolIns dataset, as well as the prompt templates.
-A.1 MolEdit
-For the molecule editing (MolEdit) task, we consider the common operations on modifying functional groups in a given molecule (i.e., add, drop, and substitute), which are simple tasks for human experts but challenging to LLMs.
+Figure 4 demonstrates the overall data construction process for TOMG-Bench as well as the OpenMolIns.
+Generally, we utilize Zinc250K to construct the test samples in TOMG-Bench, while adopting PubChem for the training samples in OpenMolIns.
+Notably, when constructing the OpenMolIns, we exclude all the molecules in Zinc250K to avoid data leakage and ensure the novelty score of the generated molecules.
+Figure 4: Data construction workflow of TOMG-Bench & OpenMolIns.
+D.1 MolEdit
+For the MolEdit task, we consider the common operations on modifying functional groups in a given molecule (i.e., add, drop, and substitute), which are simple tasks for human experts but challenging to LLMs.
 In this case, we further develop three corresponding subtasks: AddComponent, DelComponent, and SubComponent.
@@
 Substitute a {} in the molecule {} with {}.
-Table 2: Prompt Templates for MolEdit
+Table 5: Prompt Templates for MolEdit.
 Table 6 presents the functional groups that are taken into account for AddComponent and DelComponent, along with their respective selection weights.
-To reflect the distribution of these functional groups in real-world scenarios, we have implemented a weighted random selection process for AddComponent, which ensures that less common functional groups are assigned a lower probability to be chosen, thereby refining the selection mechanism to better mirror practical occurrences.
+Given that certain functional groups are encountered less frequently in real-world scenarios, we have implemented a weighted random selection process for AddComponent, which ensures that less common functional groups are assigned a lower probability to be chosen.
 Functional Group benzene ring hydroxyl aldehyde carboxyl amide
@@
 Table 6: Functional Groups that are considered in AddComponent and DelComponent, as well as their weights to be selected in AddComponent.
-For SubComponent, our focus is exclusively on end groups, which include hydroxyl, aldehyde, carboxyl, nitro, halo, nitrile, and thiol, which ensures that the editing operations are confined to substituting the existing functional group with another from this list, thereby maintaining the integrity of the molecule’s overall structure without altering it fundamentally.
-A.2 MolOpt
-Molecule optimization (MolOpt), designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task.
+For SubComponent, our focus is exclusively on end groups for simplification, which include hydroxyl, aldehyde, carboxyl, nitro, halo, nitrile, and thiol, which ensures that the editing operations are confined to substituting the existing functional group with another from this list, thereby maintaining the integrity of the molecule’s overall structure without altering it fundamentally.
+D.2 MolOpt
+MolOpt is designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task.
 Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time.
-In contrast, TOMG-Bench requires one single LLM to optimize molecules with different metrics and directions.
+In contrast, TOMG-Bench requires one single LLM to optimize molecules with different metrics and optimization directions.
 In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED.
-The prompt templates for MolOpt are illustrated in Table 4.
+The prompt templates for MolOpt are listed in Table 7.
 Prompt Templates for MolOpt
@@
 Modify the molecule {} to have a lower/higher QED value.
-Table 4: Prompt Templates for MolOpt
+Table 7: Prompt Templates for MolOpt.
+Below are the detailed explanations of the chemical properties adopted in the three subtasks of MolOpt:
 LogP refers to the logarithm of the partition coefficient, which is a measure of a molecule’s hydrophilicity or lipophilicity.
@@
 A higher QED score suggests a greater likelihood that the molecule will have favourable pharmacological properties.
-A.3 MolCustom
-To enable customized design of molecules, we think of three fundamental features for describing the molecule, including atoms, bonds, and functional groups.
+D.3 MolCustom
+To enable customized design of molecules, we consider three fundamental features for describing the molecule, including atoms, bonds, and functional groups.
 Given the specified category and number of atoms, bonds, and functional groups, LLMs should generate the molecule as we request.
@@
 The molecule contains {} group(s).
-Table 5: Prompt Templates for MolCustom
+Table 8: Prompt Templates for MolCustom.
 Atom carbon oxygen nitrogen sulfur fluorine chlorine bromine iodine phosphorus
@@
 Table 9: Atoms that are considered in AtomNum, as well as their weights to be selected.
+Higher weights indicate that they are more likely to be selected.
+For simplification, we only consider normal atoms in the molecule structures.
 AtomNum.
@@
 Table 10: Chemical bonds that are considered in BondNum, as well as their weights to be selected.
+Higher weights indicate that they are more likely to be selected.
+For simplification, we only consider normal bonds in the molecule structures.
 BondNum.
@@
 Table 11: Functional Groups that are considered in FunctionalGroup, as well as their weights to be selected.
+Higher weights indicate that they are more likely to be selected.
+For simplification, we only consider normal and edge functional groups in the molecule structures.
 FunctionalGroup. Lastly, we also specify functional groups in the molecule structure. Table 11 shows the range of functional groups and their weights that are taken into consideration.
-Notably, in MolCustom, if not specified, LLMs can generate any number of these atoms, bonds, and functional groups. However, for these specified items, LLMs should strictly follow the requirements.
-Appendix B Hyper Parameters
-In this section, we illustrate the detailed parameters adopted in this work, as shown in Table 9.
-Item Value
-Generation
-temperature 0.75
-top_p 0.85
-num_beams 1
-max_new_tokens 512
-Instruction Tuning
-batchsize 32
-lr 3e-4
-cutoff_len 1024
-Lora Settings
-r 64
-$\alpha$ 128
-dropout 0.1
-Table 9: Hyper-parameters
-Appendix C Detailed Results
-In this section, we first show the leaderboard of TOMG-Bench in Table 10, where Claude-3.5 achieves first place with a weighted average accuracy of 35.92%.
-Notably, via instruction tuning on our OpenMolIns dataset, Llama-3.1-8B achieves 6th place, which outperforms all the existing open-source LLMs and is just behind Cladue-3.
-Model #Parameters (B) $\bar{Acc}$ (%) $\bar{wAcc}(\%)$
-Claude-3.5 Anthropic (2024b) - 51.10 35.92
-Gemini-1.5-pro Deepmind (2024) - 52.25 34.80
-GPT-4-turbo Achiam et al. (2023) - 50.74 34.23
-GPT-4o Achiam et al. (2023) - 49.08 32.29
-Claude-3 Anthropic (2024a) - 46.14 30.47
-OpenMolIns-large (Llama-3.1-8B) 8 43.1 27.22
-OpenMolIns-xlarge (Galactica-125M) 0.125 44.48 25.73
-Llama3-70B-Instruct (Int4) Dubey et al. (2024) 70 38.54 23.93
-OpenMolIns-large (Galactica-125M) 0.125 39.28 23.42
-OpenMolIns-medium (Galactica-125M) 0.125 34.54 19.89
-GPT-3.5-turbo Achiam et al. (2023) - 28.93 18.58
-OpenMolIns-small (Galactica-125M) 0.125 24.17 15.18
-Llama3.1-8B-Instruct Dubey et al. (2024) 8 26.26 14.09
-Llama3-8B-Instruct Dubey et al. (2024) 8 26.40 13.75
-chatglm-9B GLM et al. (2024) 9 18.50 13.13(7)
-OpenMolIns-light (Galactica-125M) 0.125 20.95 13.13(6)
-OpenMolIns-large (Llama3.2-1B) 1 14.11 8.10
-yi-1.5-9B Young et al. (2024) 9 14.10 7.32
-Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 7 11.17 4.81
-BioT5-base Pei et al. (2023) 0.25 24.19 4.21
-MolT5-large Edwards et al. (2022) 0.78 23.11 2.89
-Llama-3.1-1B-Instruct Dubey et al. (2024) 1 3.95 1.99
-MolT5-base Edwards et al. (2022) 0.25 11.11 1.30(0)
-MolT5-small Edwards et al. (2022) 0.08 11.55 1.29(9)
-Qwen2-7B-Instruct Yang et al. (2024) 7 0.18 0.15
+Notably, in MolCustom, if not required, LLMs can generate any number of these unspecified atoms, bonds, and functional groups.
+However, for these specified items, LLMs should strictly follow the requirements.
+Appendix E Testing Process
+In this section, we present the testing process used to calculate the success rate of each subtask.
+Algorithm 1 MolEdit Testing Process
+0:  Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Group-to-add $\alpha$, Group-to-delete $\Delta$
+0:  Pass or Not (Bool) if $t$ is AddComponent then    if $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false else if $t$ is DelComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 then        return true     else        return false else if $t$ is SubComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 and $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false
+The testing process for MolEdit task is demonstrated in Algorithm 1.
+However, in this process, we can not guarantee that the modified molecule differs the original molecule only in the specified functional group.
+The reasons are to follow:
+  • The complexity of molecular structures makes it difficult to use regular matching to monitor the exact number of functional groups present.
+This is because molecules can have a wide variety of bonding patterns and spatial arrangements that are not always amenable to simple, rule-based matching techniques.
+  • Modifying functional groups in specific positions can indeed lead to unintended changes in the molecule’s overall structure.
+This is due to the interconnected nature of atoms within a molecule, where altering one part can have a ripple effect on the stability and geometry of the entire structure.
+  • The tools and algorithms currently available may not have the necessary sophistication to predict and control all the possible outcomes of a functional group modification, particularly when dealing with complex molecules or reactions that are not fully understood.
+Therefore, the similarity score needs to be considered to ensure the accuracy and quality of the generation.
+Algorithm 2 MolOpt Testing Process
+0:  : Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Optimization Direction Requirement $R$
+0:  : Pass or Not (Bool) if $t$ is LogP then    if $LogP$($m^{g}$) $>$ $LogP$($m^{o})$ and R is $\uparrow$ then        return true     else if $LogP$($m^{g}$) $<$ $LogP$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is MR then    if $M\!R$($m^{g}$) $>$ $M\!R$($m^{o})$ and R is $\uparrow$ then        return true     else if $M\!R$($m^{g}$) $<$ $M\!R$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is QED then    if $Q\!E\!D$($m^{g}$) $>$ $Q\!E\!D$($m^{o})$ and R is $\uparrow$ then        return true     else if $Q\!E\!D$($m^{g}$) $<$ $Q\!E\!D$($m^{o})$ and R is $\downarrow$ then        return true     else        return false
+For MolOpt, we adopt Algorithm 2 for the testing process. Similarly, merely passing the test does not necessarily mean that the generated molecule follows the principle of minimal change.
+Algorithm 3 MolCustom Testing Process
+0:  : Generated Molecule $m^{g}$, Subtask $t$, Atom List $A$, Bond List $B$, Functional Group List $G$, Requirements $R$
+0:  : Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for $atom$ in $A$ do         if $|atom|$($m^{g}$) $\neq$ $R$[$atom$then            flag = false else if $t$ is BondNum then    for $bond$ in $B$ do         if $|bond|$($m^{g}$) $\neq$ $R$[$bond$then            flag = false else if $t$ is FunctionalGroup then    for $group$ in $G$ do         if $|group|$($m^{g}$) $\neq$ $R$[$group$then            flag = false return flag
+For MolCustom, the testing process is shown in Algorithm 3, which test whether the specified number of the atom, bond, or functional group satisfies the requirements.
+Appendix F Detailed Results
+In this section, we first show the leaderboard of TOMG-Bench in Table 12, where Claude-3.5 achieves first place with an average weighted success rate of 35.92%.
+Notably, via instruction tuning on our OpenMolIns dataset, Llama-3.1-8B achieves 6th place, which outperforms all the existing open-source general LLMs.
+Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
+Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
+Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
+GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
+GPT-4o (OpenAI, 2023) - 49.08 32.29
+Claude-3 (Anthropic, 2024a) - 46.14 30.47
+Llama-3.1-8B (OpenMolIns-large) 8 43.1 27.22
+Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
+Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
+Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
+Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
+GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
+Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
+Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
+Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
+chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
+Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
+Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
+yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
+Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
+BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
+MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
+Llama-3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
+MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
+MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
+Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
 Table 12: Leaderboard of TOMG-Benchmark.
-Then, we present the detailed results of all the subtasks in Table 11, 12, and 13.
+Then, we compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5.
+Obviously, it can be concluded that Claude-3.5 outperforms almost all the other LLMs in these subtasks, except for the DelComponent task and the BondNum task.
+In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5, while Llama-3.1-8B (OpenMolIns-large) achieves the best performance in the BondNum task.
+Meanwhile, these radar maps also imply the difficulties of the subtasks.
+For example, for majority of the LLMs, they are struggling with the BondNum task.
+Even the most advanced proprietary LLMs like Claude-3.5 failed to achieve an ideal performance.
+Moreover, in the MolOpt task, QED, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization.
+In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
+Figure 5: The radar maps compare the performance ($\overline{W\!S\!R}$) of some representative LLMs from different model categories in each subtask of TOMG-Bench.
+Finally, we present the detailed experimental results of all the subtasks in Table 13, 14, and 15.
 Models AddComponent DelComponent SubComponent
-Accuracy Similarity Validity Accuracy Similarity Validity Accuracy Similarity Validity
-GPT-4o Achiam et al. (2023) 0.6188 0.6782 0.7412 0.7012 0.6038 0.8474 0.7992 0.7225 0.9368
-GPT-4-turbo Achiam et al. (2023) 0.699 0.6936 0.7934 0.7244 0.5735 0.906 0.7778 0.7323 0.916
-GPT-3.5-turbo Achiam et al. (2023) 0.5832 0.6545 0.798 0.3082 0.7797 0.8468 0.2918 0.6333 0.6822
-Claude-3.5 Anthropic (2024b) 0.6832 0.7017 0.4414 0.5414 0.6678 0.796 0.8104 0.731 0.9588
-Claude-3 Anthropic (2024a) 0.6766 0.684 0.818 0.5556 0.6408 0.8984 0.655 0.7159 0.9184
-Gemini-1.5-pro Deepmind (2024) 0.7058 0.6792 0.8254 0.759 0.5949 0.9158 0.7148 0.7139 0.8684
-Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.5198 0.6801 0.5922 0.6122 0.5637 0.7182 0.5094 0.717 0.6822
-Llama3-8B-Instruct Dubey et al. (2024) 0.3914 0.6649 0.5374 0.4348 0.5058 0.57 0.2602 0.6841 0.4838
-Llama3.1-8B-Instruct Dubey et al. (2024) 0.2992 0.6088 0.4954 0.4336 0.5257 0.591 0.3401 0.6424 0.5076
-Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.1868 0.6251 0.376 0.2018 0.3774 0.359 0.0602 0.6227 0.355
-Qwen2-7B-Instruct Yang et al. (2024) 0.001 0.2527 0.0036 0.0006 0.4024 0.0012 0.0004 0.2895 0.0068
-Yi-1.5-9B Young et al. (2024) 0.1742 0.417 0.4216 0.2858 0.5936 0.4909 0.137 0.4619 0.4368
-Chatglm-9B GLM et al. (2024) 0.2932 0.7622 0.5686 0.2956 0.7494 0.6914 0.1498 0.715 0.5084
-Llama-3.2-1B-Instruct Dubey et al. (2024) 0.0374 0.5343 0.1982 0.0768 0.575 0.3028 0.0102 0.3671 0.1468
-MolT5-small Edwards et al. (2022) 0.122 0.1027 0.449 0.1598 0.1125 0.4504 0.0708 0.1029 0.4876
-MolT5-base Edwards et al. (2022) 0.1354 0.1066 0.4686 0.1562 0.1144 0.4472 0.0584 0.1028 0.4426
-MolT5-large Edwards et al. (2022) 0.2834 0.1084 0.9282 0.2228 0.1201 0.9198 0.1692 0.0932 0.941
-BioT5-base Pei et al. (2023) 0.3462 0.1567 1 0.1668 0.1597 1 0.0684 0.1576 0.9998
-OpenMolIns-large (Llama-3.2-1B) 0.1756 0.5676 0.3216 0.1816 0.4963 0.2466 0.0844 0.5415 0.2958
-OpenMolIns-large (Llama-3.1-8B) 0.5822 0.6541 0.673 0.5104 0.5074 0.6896 0.544 0.6258 0.84
-OpenMolIns-light (Galactica-125M) 0.3786 0.5958 0.3786 0.2062 0.6521 0.7048 0.3102 0.5879 0.6674
-OpenMolIns-small (Galactica-125M) 0.3472 0.6172 0.5356 0.3258 0.6025 0.5758 0.2692 0.6181 0.5692
-OpenMolIns-medium (Galactica-125M) 0.4736 0.5682 0.7442 0.4886 0.5184 0.7488 0.3282 0.5975 0.6958
-OpenMolIns-large (Galactica-125M) 0.5866 0.5876 0.8228 0.6078 0.5577 0.7934 0.3438 0.6491 0.8438
-OpenMolIns-xlarge (Galactica-125M) 0.5842 0.5859 0.8438 0.6526 0.5084 0.8286 0.1872 0.6024 0.8538
-Table 11: Results on MolEdit. For each task, we highlight the best accuracy and underline the second best accuracy.
+SR Similarity WSR Validity SR Similarity WSR Validity SR Similarity WSR Validity
+GPT-4o (OpenAI, 2023) 0.6188 0.6782 0.4197 0.7412 0.7012 0.6038 0.4234 0.8474 0.7992 0.7225 0.5774 0.9368
+GPT-4-turbo (OpenAI, 2023) 0.6990 0.6936 0.4848 0.7934 0.7244 0.5735 0.4154 0.9060 0.7778 0.7323 0.5696 0.9160
+GPT-3.5-turbo (OpenAI, 2023) 0.5832 0.6545 0.3817 0.7980 0.3082 0.7797 0.2403 0.8468 0.2918 0.6333 0.1848 0.6822
+Claude-3.5 (Anthropic, 2024b) 0.6832 0.7017 0.4794 0.4414 0.5414 0.6678 0.3615 0.7960 0.8104 0.7310 0.5924 0.9588
+Claude-3 (Anthropic, 2024a) 0.6766 0.6840 0.4628 0.8180 0.5556 0.6408 0.3560 0.8984 0.6550 0.7159 0.4689 0.9184
+Gemini-1.5-pro (Deepmind, 2024) 0.7058 0.6792 0.4794 0.8254 0.7590 0.5949 0.4515 0.9158 0.7148 0.7139 0.5103 0.8684
+Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.5198 0.6801 0.3535 0.5922 0.6122 0.5637 0.3451 0.7182 0.5094 0.7170 0.3652 0.6822
+Llama3-8B-Instruct (Dubey et al., 2024) 0.3914 0.6649 0.2602 0.5374 0.4348 0.5058 0.2199 0.5700 0.2602 0.6841 0.1780 0.4838
+Llama3.1-8B-Instruct (Dubey et al., 2024) 0.2992 0.6088 0.1822 0.4954 0.4336 0.5257 0.2279 0.5910 0.3401 0.6424 0.2185 0.5076
+Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.1868 0.6251 0.1168 0.3760 0.2018 0.3774 0.0762 0.3590 0.0602 0.6227 0.0375 0.3550
+Qwen2-7B-Instruct (Yang et al., 2024) 0.0010 0.2527 0.0003 0.0036 0.0006 0.4024 0.0002 0.0012 0.0004 0.2895 0.0001 0.0068
+Yi-1.5-9B (Young et al., 2024) 0.1742 0.4170 0.0726 0.4216 0.2858 0.5936 0.1697 0.4909 0.1370 0.4619 0.0633 0.4368
+Chatglm-9B (GLM et al., 2024) 0.2932 0.7622 0.2235 0.5686 0.2956 0.7494 0.2215 0.6914 0.1498 0.7150 0.1071 0.5084
+Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0374 0.5343 0.0200 0.1982 0.0768 0.5750 0.0442 0.3028 0.0102 0.3671 0.0037 0.1468
+MolT5-small (Edwards et al., 2022) 0.1220 0.1027 0.0125 0.4490 0.1598 0.1125 0.0180 0.4504 0.0708 0.1029 0.0073 0.4876
+MolT5-base (Edwards et al., 2022) 0.1354 0.1066 0.0144 0.4686 0.1562 0.1144 0.0179 0.4472 0.0584 0.1028 0.0060 0.4426
+MolT5-large (Edwards et al., 2022) 0.2834 0.1084 0.0307 0.9282 0.2228 0.1201 0.0268 0.9198 0.1692 0.0932 0.0158 0.9410
+BioT5-base (Pei et al., 2023) 0.3462 0.1567 0.0542 1.0000 0.1668 0.1597 0.0266 1.0000 0.0684 0.1576 0.0108 0.9998
+Llama-3.2-1B (OpenMolIns-large) 0.1756 0.5676 0.0997 0.3216 0.1816 0.4963 0.0901 0.2466 0.0844 0.5415 0.0457 0.2958
+Llama-3.1-8B (OpenMolIns-large) 0.5822 0.6541 0.3808 0.6730 0.5104 0.5074 0.2590 0.6896 0.5440 0.6258 0.3404 0.8400
+Galactica-125M (OpenMolIns-light) 0.3786 0.5958 0.2256 0.6842 0.2062 0.6521 0.1345 0.7048 0.3102 0.5879 0.1824 0.6674
+Galactica-125M (OpenMolIns-small) 0.3472 0.6172 0.2143 0.5356 0.3258 0.6025 0.1963 0.5758 0.2692 0.6181 0.1664 0.5692
+Galactica-125M (OpenMolIns-medium) 0.4736 0.5682 0.2691 0.7442 0.4886 0.5184 0.2533 0.7488 0.3282 0.5975 0.1961 0.6958
+Galactica-125M (OpenMolIns-large) 0.5866 0.5876 0.3447 0.8228 0.6078 0.5577 0.3390 0.7934 0.3438 0.6491 0.2232 0.8438
+Galactica-125M (OpenMolIns-xlarge) 0.5842 0.5859 0.3423 0.8438 0.6526 0.5084 0.3318 0.8286 0.1872 0.6024 0.1128 0.8538
+Table 13: Detailed results on MolEdit. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
 Models LogP MR QED
-Accuracy Similarity Validity Accuracy Similarity Validity Accuracy Similarity Validity
-GPT-4o Achiam et al. (2023) 0.719 0.6586 0.8796 0.6864 0.642 0.8352 0.3952 0.618 0.857
-GPT-4-turbo Achiam et al. (2023) 0.7662 0.6984 0.9048 0.7388 0.6821 0.8848 0.3946 0.6587 0.905
-GPT-3.5-turbo Achiam et al. (2023) 0.4048 0.6327 0.854 0.412 0.6263 0.8486 0.3316 0.5635 0.8354
-Claude-3.5 Anthropic (2024b) 0.797 0.7124 0.9422 0.6962 0.7112 0.911 0.5361 0.7042 0.8604
-Claude-3 Anthropic (2024a) 0.7984 0.6067 0.9096 0.6094 0.6398 0.9062 0.4678 0.5855 0.9044
-Gemini-1.5-pro Deepmind (2024) 0.7712 0.7022 0.9274 0.7876 0.6744 0.8926 0.4704 0.6077 0.9484
-Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.5984 0.6028 0.6482 0.5684 0.6032 0.6272 0.2774 0.4828 0.634
-Llama3-8B-Instruct Dubey et al. (2024) 0.4642 0.3658 0.6086 0.4332 0.4793 0.5704 0.2568 0.4547 0.6112
-Llama3.1-8B-Instruct Dubey et al. (2024) 0.399 0.4235 0.5122 0.4164 0.483 0.5238 0.2655 0.4499 0.6158
-Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.222 0.4501 0.2802 0.1908 0.2578 0.3795 0.121 0.3244 0.2532
-Qwen2-7B-Instruct Yang et al. (2024) 0 0.2923 0.0004 0.0002 0.4123 0.0004 0 0 0
-Yi-1.5-9B 0.2884 0.5461 0.4927 0.205 0.3724 0.4126 0.1064 0.6596 0.4526
-Chatglm-9B 0.3666 0.6902 0.4736 0.3514 0.682 0.5 0.1832 0.6506 0.4342
-Llama-3.2-1B-Instruct Dubey et al. (2024) 0.0644 0.5055 0.1664 0.0822 0.441 0.1604 0.0714 0.4757 0.1796
-MolT5-small 0.2158 0.1052 0.4302 0.2316 0.1011 0.442 0.2214 0.1031 0.4326
-MolT5-base 0.2074 0.1051 0.4168 0.1856 0.1073 0.3796 0.2358 0.1054 0.4536
-MolT5-large 0.4244 0.1015 0.8156 0.4496 0.1072 0.8678 0.4654 0.119 0.9214
-BioT5-base 0.5158 0.1526 1 0.506 0.1597 1 0.5068 0.158 1
-OpenMolIns-large (Llama-3.2-1B) 0.2898 0.5951 0.385 0.2644 0.5956 0.3678 0.1996 0.5849 0.349
-OpenMolIns-large (Llama-3.1-8B) 0.8054 0.6678 0.872 0.7122 0.6548 0.8514 0.5224 0.6398 0.8802
-OpenMolIns-light (Galactica-125M) 0.3202 0.6547 0.6416 0.3508 0.6435 0.6358 0.269 0.6521 0.638
-OpenMolIns-small (Galactica-125M) 0.4172 0.642 0.5568 0.3958 0.6452 0.5338 0.2956 0.6385 0.5376
-OpenMolIns-medium (Galactica-125M) 0.5904 0.5812 0.789 0.5874 0.5873 0.7384 0.4608 0.5859 0.7768
-OpenMolIns-large (Galactica-125M) 0.6454 0.5927 0.8198 0.6388 0.5973 0.8028 0.495 0.5962 0.81
-OpenMolIns-xlarge (Galactica-125M) 0.7362 0.5744 0.8902 0.7124 0.5697 0.8612 0.5786 0.5677 0.8626
-Table 12: Results on MolOpt. For each task, we highlight the best accuracy and underline the second best accuracy.
+SR Similarity WSR Validity SR Similarity WSR Validity SR Similarity WSR Validity
+GPT-4o (OpenAI, 2023) 0.7190 0.6586 0.4735 0.8796 0.6864 0.6420 0.4407 0.8352 0.3952 0.6180 0.2442 0.8570
+GPT-4-turbo (OpenAI, 2023) 0.7662 0.6984 0.5351 0.9048 0.7388 0.6821 0.5039 0.8848 0.3946 0.6587 0.2599 0.9050
+GPT-3.5-turbo (OpenAI, 2023) 0.4048 0.6327 0.2561 0.8540 0.4120 0.6263 0.2580 0.8486 0.3316 0.5635 0.1869 0.8354
+Claude-3.5 (Anthropic, 2024b) 0.7970 0.7124 0.5678 0.9422 0.6962 0.7112 0.4951 0.9110 0.5361 0.7042 0.3775 0.8604
+Claude-3 (Anthropic, 2024a) 0.7984 0.6067 0.4844 0.9096 0.6094 0.6398 0.3899 0.9062 0.4678 0.5855 0.2739 0.9044
+Gemini-1.5-pro (Deepmind, 2024) 0.7712 0.7022 0.5415 0.9274 0.7876 0.6744 0.5312 0.8926 0.4704 0.6077 0.2859 0.9484
+Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.5984 0.6028 0.3607 0.6482 0.5684 0.6032 0.3429 0.6272 0.2774 0.4828 0.1339 0.6340
+Llama3-8B-Instruct (Dubey et al., 2024) 0.4642 0.3658 0.1698 0.6086 0.4332 0.4793 0.2076 0.5704 0.2568 0.4547 0.1168 0.6112
+Llama3.1-8B-Instruct (Dubey et al., 2024) 0.3990 0.4235 0.1690 0.5122 0.4336 0.5257 0.2279 0.5910 0.2655 0.4499 0.1194 0.6158
+Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.2220 0.4501 0.0999 0.2802 0.1908 0.2578 0.0492 0.3795 0.1210 0.3244 0.0393 0.2532
+Qwen2-7B-Instruct (Yang et al., 2024) 0.0000 0.2923 0.0000 0.0004 0.0002 0.4123 0.0001 0.0004 0.0000 0.0000 0.0000 0.0000
+Yi-1.5-9B (Young et al., 2024) 0.2884 0.5461 0.1575 0.4927 0.2050 0.3724 0.0763 0.4126 0.1064 0.6596 0.0702 0.4526
+Chatglm-9B (GLM et al., 2024) 0.3666 0.6902 0.2530 0.4736 0.3514 0.6820 0.2397 0.5000 0.1832 0.6506 0.1192 0.4342
+Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0644 0.5055 0.0326 0.1664 0.0822 0.4410 0.0363 0.1604 0.0714 0.4757 0.0340 0.1796
+MolT5-small (Edwards et al., 2022) 0.2158 0.1052 0.0227 0.4302 0.2316 0.1011 0.0234 0.4420 0.2214 0.1031 0.0228 0.4326
+MolT5-base (Edwards et al., 2022) 0.2074 0.1051 0.0218 0.4168 0.1856 0.1073 0.0199 0.3796 0.2358 0.1054 0.0249 0.4536
+MolT5-large (Edwards et al., 2022) 0.4244 0.1015 0.0431 0.8156 0.4496 0.1072 0.0482 0.8678 0.4654 0.1190 0.0554 0.9214
+BioT5-base (Pei et al., 2023) 0.5158 0.1526 0.0787 1.0000 0.5060 0.1597 0.0808 1.0000 0.5068 0.1580 0.0801 1.0000
+Llama-3.2-1B (OpenMolIns-large) 0.2898 0.5951 0.1725 0.3850 0.2644 0.5956 0.1575 0.3678 0.1996 0.5849 0.1167 0.3490
+Llama-3.1-8B (OpenMolIns-large) 0.8054 0.6678 0.5378 0.8720 0.7122 0.6548 0.4663 0.8514 0.5224 0.6398 0.3342 0.8802
+Galactica-125M (OpenMolIns-light) 0.3202 0.6547 0.2096 0.6416 0.3508 0.6435 0.2257 0.6358 0.2690 0.6521 0.1754 0.6380
+Galactica-125M (OpenMolIns-small) 0.4172 0.6420 0.2678 0.5568 0.3958 0.6452 0.2554 0.5338 0.2956 0.6385 0.1887 0.5376
+Galactica-125M (OpenMolIns-medium) 0.5904 0.5812 0.3431 0.7890 0.5874 0.5873 0.3450 0.7384 0.4608 0.5859 0.2700 0.7768
+Galactica-125M (OpenMolIns-large) 0.6454 0.5927 0.3825 0.8198 0.6388 0.5973 0.3816 0.8028 0.4950 0.5962 0.2951 0.8100
+Galactica-125M (OpenMolIns-xlarge) 0.7362 0.5744 0.4229 0.8902 0.7124 0.5697 0.4059 0.8612 0.5786 0.5677 0.3285 0.8626
+Table 14: Detailed results on MolOpt. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
 Models AtomNum BondNum FunctionalGroup
-Accuracy Novelty Validity Accuracy Novelty Validity Accuracy Novelty Validity
-GPT-4o Achiam et al. (2023) 0.1998 0.6703 0.5852 0.065 0.6336 0.8564 0.233 0.6513 0.859
-GPT-4-turbo Achiam et al. (2023) 0.1702 0.6991 0.4904 0.0774 0.6301 0.9068 0.218 0.6605 0.8778
-GPT-3.5-turbo Achiam et al. (2023) 0.107 0.5054 0.6947 0.0518 0.6871 0.5522 0.1136 0.6585 0.8686
-Claude-3.5 Anthropic (2024b) 0.1928 0.6926 0.6548 0.1058 0.6584 0.886 0.2364 0.6582 0.8892
-Claude-3 Anthropic (2024a) 0.1044 0.6833 0.591 0.1042 0.6598 0.8696 0.1816 0.9158 0.6644
-Gemini-1.5-pro Deepmind (2024) 0.1742 0.6902 0.6774 0.0708 0.6522 0.8688 0.2486 0.6673 0.924
-Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.1404 0.6675 0.5474 0.067 0.6478 0.7378 0.1752 0.6576 0.765
-Llama3-8B-Instruct Dubey et al. (2024) 0.0242 0.6649 0.3812 0.026 0.6303 0.57 0.0848 0.6167 0.7216
-Llama3.1-8B-Instruct Dubey et al. (2024) 0.0228 0.702 0.3862 0.0395 0.6541 0.6387 0.13 0.6274 0.6905
-Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.0078 0.6732 0.2986 0.0102 0.6309 0.4524 0.0048 0.6012 0.402
-Qwen2-7B-Instruct Yang et al. (2024) 0.011 0.9061 0.2622 0.001 0.8645 0.0796 0.0022 0.8601 0.0622
-Yi-1.5-9B Young et al. (2024) 0.0392 0.6848 0.617 0.0208 0.6407 0.7072 0.0126 0.6945 0.6521
-Chatglm-9B GLM et al. (2024) 0.0002 0.7483 0.2131 0.0254 0.7189 0.4682 0 0.6908 0.5926
-Llama-3.2-1B-Instruct Dubey et al. (2024) 0.004 0.6807 0.185 0.008 0.7465 0.2226 0.0008 0.7461 0.2818
-MolT5-small Edwards et al. (2022) 0.0006 0.6586 0.661 0.0064 0.598 0.6202 0.0114 0.5287 0.8354
-MolT5-base Edwards et al. (2022) 0.0008 0.6868 0.756 0.007 0.6509 0.8422 0.013 0.5464 0.8382
-MolT5-large Edwards et al. (2022) 0.015 0.7103 0.8412 0.0118 0.5611 0.8916 0.0382 0.6088 0.9406
-BioT5-base Pei et al. (2023) 0.0118 0.8353 0.995 0.0078 0.6667 0.9992 0.0476 0.6792 0.9998
-OpenMolIns-large (LLama-3.2-1B) 0.0144 0.649 0.5616 0.035 0.615 0.6186 0.0252 0.6373 0.4412
-OpenMolIns-large (LLama-3.1-8B) 0.0136 0.6634 0.7582 0.0544 0.6614 0.7456 0.1344 0.6396 0.6435
-OpenMolIns-light (Galactica-125M) 0.0044 0.6054 0.793 0.0216 0.5724 0.7596 0.0244 0.5756 0.8442
-OpenMolIns-small (Galactica-125M) 0.0146 0.6568 0.8424 0.053 0.6365 0.7926 0.057 0.5954 0.8874
-OpenMolIns-medium (Galactica-125M) 0.0294 0.6553 0.8698 0.0622 0.6473 0.7474 0.0882 0.6091 0.8932
-OpenMolIns-large (Galactica-125M) 0.0464 0.6729 0.9116 0.0716 0.6695 0.7374 0.0996 0.6276 0.8966
-OpenMolIns-xlarge (Galactica-125M) 0.1862 0.6899 0.9308 0.1656 0.6887 0.7952 0.2006 0.6445 0.9162
-Table 13: Results on MolCustom. For each task, we highlight the best accuracy and underline the second best accuracy.
+SR Novelty WSR Validity SR Novelty WSR Validity SR Novelty WSR Validity
+GPT-4o (OpenAI, 2023) 0.1998 0.6703 0.1339 0.5852 0.0650 0.6336 0.0412 0.8564 0.2330 0.6513 0.1518 0.8590
+GPT-4-turbo (OpenAI, 2023) 0.1702 0.6991 0.1190 0.4904 0.0774 0.6301 0.0488 0.9068 0.2180 0.6605 0.1440 0.8778
+GPT-3.5-turbo (OpenAI, 2023) 0.1070 0.5054 0.0541 0.6947 0.0518 0.6871 0.0356 0.5522 0.1136 0.6585 0.0748 0.8686
+Claude-3.5 (Anthropic, 2024b) 0.1928 0.6926 0.1335 0.6548 0.1058 0.6584 0.0697 0.8860 0.2364 0.6582 0.1556 0.8892
+Claude-3 (Anthropic, 2024a) 0.1044 0.6833 0.0713 0.5910 0.1042 0.6598 0.0688 0.8696 0.1816 0.9158 0.1663 0.6644
+Gemini-1.5-pro (Deepmind, 2024) 0.1742 0.6902 0.1202 0.6774 0.0708 0.6522 0.0462 0.8688 0.2486 0.6673 0.1659 0.9240
+Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.1404 0.6675 0.0937 0.5474 0.0670 0.6478 0.0434 0.7378 0.1752 0.6576 0.1152 0.7650
+Llama3-8B-Instruct (Dubey et al., 2024) 0.0242 0.6649 0.0161 0.3812 0.0260 0.6303 0.0164 0.5700 0.0848 0.6167 0.0523 0.7216
+Llama3.1-8B-Instruct (Dubey et al., 2024) 0.0228 0.7020 0.0160 0.3862 0.0395 0.6541 0.0258 0.6387 0.1300 0.6274 0.0816 0.6905
+Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.0078 0.6732 0.0053 0.2986 0.0102 0.6309 0.0064 0.4524 0.0048 0.6012 0.0029 0.4020
+Qwen2-7B-Instruct (Yang et al., 2024) 0.0110 0.9061 0.0100 0.2622 0.0010 0.8645 0.0090 0.0796 0.0022 0.8601 0.0019 0.0622
+Yi-1.5-9B (Young et al., 2024) 0.0392 0.6848 0.0268 0.6170 0.0208 0.6407 0.0133 0.7072 0.0126 0.6945 0.0088 0.6521
+Chatglm-9B (GLM et al., 2024) 0.0002 0.7483 0.0001 0.2131 0.0254 0.7189 0.0183 0.4682 0.0000 0.6908 0.0000 0.5926
+Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0040 0.6807 0.0027 0.1850 0.0080 0.7465 0.0060 0.2226 0.0008 0.7461 0.0006 0.2818
+MolT5-small (Edwards et al., 2022) 0.0006 0.6586 0.0004 0.6610 0.0064 0.5980 0.0038 0.6202 0.0114 0.5287 0.0060 0.8354
+MolT5-base (Edwards et al., 2022) 0.0008 0.6868 0.0005 0.7560 0.0070 0.6509 0.0046 0.8422 0.0130 0.5464 0.0071 0.8382
+MolT5-large (Edwards et al., 2022) 0.0150 0.7103 0.0107 0.8412 0.0118 0.5611 0.0066 0.8916 0.0382 0.6088 0.0233 0.9406
+BioT5-base (Pei et al., 2023) 0.0118 0.8353 0.0099 0.9950 0.0078 0.6667 0.0052 0.9992 0.0476 0.6792 0.0323 0.9998
+Llama-3.2-1B (OpenMolIns-large) 0.0144 0.6490 0.0093 0.5616 0.0350 0.6150 0.0215 0.6186 0.0252 0.6373 0.0161 0.4412
+Llama-3.1-8B (OpenMolIns-large) 0.0136 0.6634 0.0090 0.7582 0.0544 0.6614 0.0360 0.7456 0.1344 0.6396 0.0860 0.6435
+Galactica-125M (OpenMolIns-light) 0.0044 0.6054 0.0026 0.7930 0.0216 0.5724 0.0124 0.7596 0.0244 0.5756 0.0140 0.8442
+Galactica-125M (OpenMolIns-small) 0.0146 0.6568 0.0096 0.8424 0.0530 0.6365 0.0337 0.7926 0.0570 0.5954 0.0339 0.8874
+Galactica-125M (OpenMolIns-medium) 0.0294 0.6553 0.0193 0.8698 0.0622 0.6473 0.0403 0.7474 0.0882 0.6091 0.0537 0.8932
+Galactica-125M (OpenMolIns-large) 0.0464 0.6729 0.0312 0.9116 0.0716 0.6695 0.0479 0.7374 0.0996 0.6276 0.0625 0.8966
+Galactica-125M (OpenMolIns-xlarge) 0.1862 0.6899 0.1285 0.9308 0.1656 0.6887 0.1140 0.7952 0.2006 0.6445 0.1293 0.9162
+Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).

### v3

@@
-TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation
+Jiatong Li [^1] footnotemark: 1    Junxian Li3    Weida Wang2,4    Yunqing Liu1    Changmeng Zheng1 Dongzhan Zhou [^2] footnotemark: 2    Xiao-yong Wei1    Qing Li1
 Abstract
-In this paper, we propose Text-based Open Molecule Generation Benchmark (TOMG-Bench), the first benchmark to evaluate the open-domain molecule generation capability of LLMs.
-TOMG-Bench encompasses a dataset of three major tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom).
-Each major task further contains three subtasks, while each subtask comprises 5,000 test samples.
-Given the inherent complexity of open molecule generation evaluation, we also developed an automated evaluation system that helps measure both the quality and the accuracy of the generated molecules.
-Our comprehensive benchmarking of 25 LLMs reveals the current limitations as well as potential areas for improvement in text-guided molecule discovery.
-Furthermore, we propose OpenMolIns, a specialized instruction tuning dataset established for solving challenges raised by TOMG-Bench.
-Fine-tuned on OpenMolIns, Llama3.1-8B could outperform all the open-source general LLMs, even surpassing GPT-3.5-turbo by 46.5% on TOMG-Bench.
-Our codes and datasets are available through https://github.com/phenixace/TOMG-Bench.
+Recently, Large Language Models (LLMs) have shown great potential in natural language-driven molecule discovery.
+However, existing datasets and benchmarks for molecule-text alignment are predominantly built on a one-to-one mapping, measuring LLMs’ ability to retrieve a single, pre-defined answer, rather than their creative potential to generate diverse, yet equally valid, molecular candidates.
+To address this critical gap, we propose Speak-to-Structure (S2-Bench), the first benchmark to evaluate LLMs in open-domain natural language-driven molecule generation.
+S2-Bench is specifically designed for one-to-many relationships, challenging LLMs to demonstrate genuine molecular understanding and generation capabilities.
+Our benchmark includes three key tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom), each probing a different aspect of molecule discovery.
+We also introduce OpenMolIns, a large-scale instruction tuning dataset that enables Llama-3.1-8B to surpass the most powerful LLMs like GPT-4o and Claude-3.5 on S2-Bench.
+Our comprehensive evaluation of 28 LLMs shifts the focus from simple pattern recall to realistic molecular design, paving the way for more capable LLMs in natural language-driven molecule discovery.
 1 Introduction
 Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023).
-Normally, molecule discovery is a trial-and-error process (Ekins, 2024), which requires repetitive experimentation and data analysis (Mattern & Grosser, 2023).
-Due to the inefficiency of the traditional techniques, it usually takes more than 10 years to bring a new drug candidate into the market (Lee et al., 2018).
-With the development of machine learning techniques and the advent of Graph Neural Networks (GNNs) (Wang et al., 2023), there has been a significant step forward.
-As molecules could be represented as graphs, GNN-based methods can capture the structural patterns of the molecule and make accurate predictions.
-With the assistance of GNNs, researchers could analyse the properties of molecules (Cai et al., 2022) and generate new molecule candidates (Jin et al., 2018).
-However, challenges still exist.
-GNN-based methods struggle to generalize to different tasks (Chen et al., 2024), necessitating costly data collection and preparation for different downstream tasks.
-Moreover, these methods are constrained in their capacity to generate molecules with specific, customized properties, limiting their flexibility in molecular design (Li et al., 2024c).
-In contrast, Large Language Models (LLMs) have shown their great generalization capability (OpenAI, 2023) and could be easily adapted to different research fields.
-For instance, Cascella et al.
-(2023) utilize ChatGPT for supporting clinical practice and Zhen et al.
-(2024) adopts LLMs as assistants for task planning in the field of Civil Engineering, showing the great potential of LLMs in scientific discovery.
-Figure 1: Comparison of Text-based Targeted Molecule Generation (a) v.s. Text-based Open Molecule Generation (b).
-As molecules can be represented as texts by Simplified Molecular Input Line Entry System (SMILES), a linear notation that encapsulates the structure of a chemical compound (Weininger, 1988), they can be processed and understood by LLMs, bridging the gap between molecules and natural languages.
-With advanced reasoning and in-context learning capabilities, LLMs are particularly adept at generalizing to the molecule domain (Li et al., 2024b).
-The generalization capability makes LLMs a viable option in molecule discovery.
-Furthermore, by aligning molecules with textual data, LLMs can serve as powerful assistants to chemists (Zhang et al., 2024; Li et al., 2024d).
-They could help interpret and generate chemical knowledge, suggest modifications to molecular structures, and even predict the properties and behaviours of compounds, which would potentially streamline the molecule discovery process, leading to breakthroughs in diverse research areas.
-While the integration of LLMs into molecule discovery holds immense promise, the process of aligning molecules with textual data is challenging (Li et al., 2024a).
-A significant challenge lies in the availability and diversity of datasets and benchmarks necessary for training and evaluation.
-Although the molecule-caption translation datasets including ChEBI-20 (Edwards et al., 2022), PubChem324K (Liu et al., 2023) and L+M-24 (Edwards et al., 2024) are crucial for bridging the gap between the molecular and textual domains, they still face several limitations that need to be addressed:
-On the one hand, there is a concern about the generalization of models trained on the molecule-caption translation task (Li et al., 2024b).
-In real-world scenarios, molecule captions that describe molecular structures can be highly ambiguous, with multiple correct interpretations, while the current molecule-caption translation is actually a targeted generation task.
-In this case, these models often struggle to generalize to customized molecules, even for seemingly simple examples (Li et al., 2024b).
-This suggests a fundamental mismatch between the molecular and textual spaces, raising questions about whether this task could truly guide LLMs well.
-On the other hand, a critical issue is the inability to propose new molecule structures.
-The ultimate goal of molecule discovery is not just to understand and describe existing chemical compounds but to innovate and discover new ones, particularly in the context of drug discovery, indicating that the current molecule-caption translation task and the corresponding evaluation metrics fall short in this regard.
-Therefore, addressing these challenges is essential for harnessing the full potential of LLMs in molecule discovery.
-In our efforts to bridge the gap between the natural language and the molecular spaces and to further facilitate LLMs as chemist assistants in molecule discovery, we propose a novel benchmark, Text-based Open Molecule Generation Benchmark (TOMG-Bench).
-TOMG-Bench is designed to evaluate the open-domain generative capabilities of LLMs in the molecular domain through a series of structured instructions for molecule design and operations.
-As shown in Figure 1, different from the previous targeted molecule generation tasks, Text-based Open Molecule Generation does not set a specific target or enables LLMs to generate an exactly matched molecule.
-Instead, we design automatic evaluation processes with chemical toolboxes like RDKit (Landrum, 2013) to test whether the generation meets the requirements.
-In other words, there can be multiple correct answers for a single question, and LLMs are only required to generate one of them.
-Notably, TOMG-Bench is meticulously categorized into three primary tasks, i.e., molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom).
-Each category contains three subtasks, and each subtask is composed of 5,000 test samples, providing a comprehensive assessment of whether LLMs truly grasp the molecular space.
-Meanwhile, we also propose a different set of evaluation metrics to evaluate and rank the performance of LLMs, which considers both the accuracy and quality of the generated molecules.
-To solve the challenges raised by TOMG-Bench, we propose an instruction-tuning dataset, OpenMolIns, by extracting and reformatting molecules with meticulously designed rules from an existing molecule database.
-OpenMolIns is structured across five distinct data levels (i.e., light, small, medium, large, and xlarge) to study the data scaling law of LLMs in the text-based open molecule generation task.
-Our contributions are primarily threefold:
+Traditionally, this is a trial-and-error process (Ekins, 2024) that requires extensive experimentation and data analysis (Mattern & Grosser, 2023), often taking over a decade to bring a new drug candidate to market (Lee et al., 2018).
+Recently, Large Language Models (LLMs) have demonstrated great potential in molecule discovery (Edwards et al., 2021; 2022) by leveraging their powerful language understanding and robust reasoning abilities (OpenAI, 2023).
+Molecules can be represented as SMILES strings (Weininger, 1988), which allows LLMs to process molecules as textual strings, effectively bridging the gap between molecular structures and natural language.
+Meanwhile, by aligning molecules with textual descriptions (Edwards et al., 2022), LLMs can assist chemists in interpreting chemical knowledge, suggesting structural modifications, and predicting compound properties (Li et al., 2024a; Zhang et al., 2024), thereby significantly streamlining the molecule discovery process.
+While the integration of LLMs holds immense promise, a significant challenge lies in the limitations of current datasets and benchmarks.
+For example, existing datasets for aligning molecules and texts like ChEBI-20 (Edwards et al., 2022) and PubChem324K (Liu et al., 2023), are constructed based on a one-to-one mapping assumption, where each textual description is linked to a single, predefined target molecule due to the convenience of acquiring ground truth data from existing databases.
+However, the one-to-one mapping presents a key limitation: it does not fully align with the nature of real-world molecule discovery.
+In practice, multiple distinct molecules can often share the same desired properties or biological activity Petrone et al.
+(2012).
+For instance, a particular pharmaceutical effect is rarely unique to a single compound, and a material with a specific physical property, like tensile strength, can be realized through various molecular structures.
+This mismatch between the evaluation paradigm and the reality of molecule discovery casts doubt on whether these datasets and benchmarks genuinely assess an LLM’s capacity for creative molecular design, or if they instead inadvertently encourage models to rely on memorization and pattern-matching.
+To bridge this critical gap, we introduce a novel benchmark, Speak-to-Structure (S2-Bench), the first benchmark designed to evaluate the open-domain natural language-driven molecule generation capabilities of LLMs.
+S2-Bench is meticulously structured into three primary tasks, as shown in Figure 1, each targeting a specific, real-world capability essential for drug and materials science:
+Molecule Editing (MolEdit) focuses on testing an LLM’s ability to perform precise, localized structural modifications while preserving main structures.
+In drug discovery, this is analogous to Lead Optimization Barcelos et al.
+(2022), which edits the lead compound to create new variants.
+This requires a deep understanding of chemical grammar, including valency, stereochemistry, and ring systems, ensuring that edits could create valid molecules.
+Molecule Optimization (MolOpt) extends the MolEdit task by evaluating a model’s ability to optimize a lead compound under specified property constraints (e.g., increasing solubility or reducing toxicity).
+Its open-ended design requires the model to edit and refine molecular structures in accordance with the desired properties, thereby showcasing its understanding of both structural modifications and property-related characteristics.
+Customized Molecule Generation (MolCustom) requires LLMs to synthesize a novel molecule based on a natural language description of its desired structural components.
+Specifically, it challenges models to generate molecules with a set of quantitative and qualitative constraints, such as a specific number of atoms, bonds, or functional groups, which directly mimics the initial drug design phase where a chemist defines a new compound based on a set of precise requirements for its structure.
+Figure 1: Task illustration of S2-Bench for open domain natural language-driven molecule generation.
+In contrast to text-based target molecule generation, multiple valid molecules may fulfill the textual requirements (right of the arrow).
+To facilitate the evaluation of the above tasks, we also introduce a robust evaluation system with tailored examination process, enabling a comprehensive assessment of LLMs’ performance in open-domain natural language-driven molecule discovery.
+Furthermore, we propose OpenMolIns, a large-scale instruction-tuning dataset, comprising up to 120 million instruction–molecule pairs, enabling Llama3.1-8B to achieve the state-of-the-art performance among 28 powerful LLMs on S2-Bench.
+To summarize, our contributions are primarily threefold:
   • 1.
-    TOMG-Bench introduces the task of Text-based Open Molecule Generation, which consists of three major tasks and their nine subtasks in total to synthetically benchmark the performance of LLMs in this task, demonstrating the limitations of the targeted molecule generation datasets like ChEBI-20.
+    We introduce Speak-to-Structure (S2-Bench), the first benchmark for open-domain natural language-driven molecule generation.
+Moving from one-to-one to one-to-many relationships, S2-Bench better aligns concepts in drug discovery and provides a novel perspective on assessing LLMs’ genuine molecular understanding and design capabilities.
   • 2.
-    We benchmarked a suite of 25 LLMs, providing a brand-new perspective for the evaluation of LLMs.
+    We introduce OpenMolIns, an instruction-tuning dataset with up to 120 million instruction–molecule pairs, enabling Llama3.1-8B to achieve SOTA performance on S2-Bench.
   • 3.
-    We introduce OpenMolIns, an instruction tuning dataset, enabling Llama-3.1-8B-Instruct to outperform all the open-source general LLMs in TOMG-Bench.
-2 Related Work
-2.1 Development of AIs in Molecule Discovery
-Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design (Du et al., 2022).
-Thus, integrating artificial intelligence into molecule discovery has marked a transformative shift in the pharmaceutical landscape, significantly enhancing the efficiency and effectiveness of identifying and developing new therapeutic molecules.
-Recent advancements in machine learning (ML), deep learning (DL), and natural language processing (NLP) have enabled AI systems to analyze complex biological and chemical data more effectively than traditional methods (Wigh et al., 2022; Wu et al., 2018; Zhou et al., 2023).
-For instance, MolReGPT  (Li et al., 2024b) leverages LLMs like ChatGPT to learn molecule SMILES strings representation for molecule-caption translation tasks.
-Moreover, existing studies have explored advanced methods that utilize various AI techniques to further enhance molecule discovery processes, including Convolutional Neural Networks (CNNs) (Peng & Zhao, 2019; Le et al., 2019), Recurrent Neural Networks (RNNs) (Grisoni et al., 2020; Popova et al., 2019), Graph Neural Networks (GNNs) (Wang et al., 2023; Sun et al., 2022), and Transformer-based networks (Xia et al., 2023; Balaji et al., 2023; Edwards et al., 2022).
-2.2 Text-based Molecule Generation
-Text-based Molecule Generation (Edwards et al., 2021) has recently emerged as a transformative task to molecule discovery.
-Previously, this task centres on generating targeted molecules based on natural language descriptions, requiring the creation of paired datasets of molecules and their corresponding textual representations.
-Early approaches leveraged transformer-based models like MolT5 (Edwards et al., 2022), employing self-supervised learning on large datasets to generate high-quality SMILES strings from textual inputs.
-Subsequent advancements, such as KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023), integrated molecular graphs or biochemical text to improve both understanding and generation capabilities.
-3D-MoLM (Li et al., 2024e) further enhanced this by incorporating spatial configurations, leading to more accurate and geometrically valid molecular representations.
-Meanwhile, the application of LLMs like MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) as in-context learners has also shown significant promise.
-These models demonstrate the ability of LLMs to adaptively generate molecules by retrieving and leveraging relevant examples from the provided context.
-Most recently, MolReFlect (Li et al., 2024c) underscored the importance of fine-grained alignment between molecular structures and their textual descriptions, utilizing a teacher-student training paradigm to capture these nuanced relationships effectively.
-3 TOMG-Bench
-In this section, we propose TOMG-Bench to comprehensively assess the performance of LLMs in molecular space.
-Specifically, the benchmark is composed of three basic tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom).
-To ensure the integrity and effectiveness of TOMG-Bench, we have developed a robust set of evaluation metrics for different tasks.
-Additionally, we created OpenMolIns, an instruction-tuning dataset aimed at enhancing the performance and adaptability of LLMs to the challenges presented by this benchmark.
-Figure 2: Task illustrations of TOMG-Bench. Both molecules to the right of the arrow are acceptable solutions.
-3.1 Dataset Categorization
-The categorization of the TOMG-Bench dataset initially considers the inherent characteristics of molecule SMILES representation and the role of LLMs serving as the chemist’s assistant, namely helping the chemists to edit, optimize and customize molecules as they want.
-As shown in Figure 2, following the difficulty of the tasks, we demonstrate the three basic tasks as well as their corresponding subtasks:
-MolEdit emerges as the most straightforward task among the three domains, as an existing molecule is already provided, and LLMs are only required to make modifications to it, which tests the molecular structure knowledge of LLMs.
-In this case, we have crafted three subtasks for MolEdit: AddComponent, DelComponent, and SubComponent.
-In AddComponent, LLMs are instructed to add a specific functional group to the given molecule, and DelComponent challenges LLMs to remove a specified functional group from the provided molecule, while SubComponent is a hybrid of the previous two subtasks, requiring LLMs to first remove a designated functional group and then introduce a new one as specified to the molecule.
-MolOpt challenges LLMs to not only edit molecules but also to discern whether the modification will steer the molecule towards a desired optimization target.
-To assess this capability, we concentrate on three pivotal properties that are vital for molecule discovery and can be effectively validated: LogP (Octanol-water partition coefficient, a metric of lipophilicity), MR (molecular refractivity, a proxy for the molar refractive index), and QED (Quantitative Estimate of Druglikeness, an assessment of drug-like characteristics).
-These metrics offer critical information about the potential pharmacological attributes of the molecule, which could help chemists filter molecules as viable drug candidates.
-MolCustom is the most challenging task, where we have established three subtasks: AtomNum, BondNum, and FunctionalGroup.
-For AtomNum, LLMs are tasked with generating molecules that adhere to a specified count and type of atoms.
-BondNum involves the creation of molecules with a defined number and type of bonds.
-In FunctionalGroup, LLMs must generate a molecule that includes functional groups as specified.
-These subtasks may appear straightforward, yet they are deceptively challenging.
-They demand that LLMs have a sophisticated understanding of molecular syntax to precisely generate molecules that meet the complex criteria set forth.
-Compared to previous text-based target molecule generation datasets, the task design in TOMG-Bench is rather naive, but focuses more on the details of the molecular substructure, which puts even more rigorous requirements on both the molecular understanding and generation capabilities of LLMs.
-Furthermore, the accurate understanding on molecule substructures and editing operations also ensures the explainability and trustworthiness of generated molecules.
-3.2 Dataset Construction
-Previously, molecule-related datasets and tasks have been hindered by the scarcity of human annotations.
-For instance, ChEBI-20  (Edwards et al., 2022), a dataset for the molecule-caption translation task, contains only 26,407 samples for training.
-In contrast, datasets for the text-to-image generation task in the computer vision field like MS COCO  (Chen et al., 2015) have over 1,500,000 annotated captions on more than 330,000 images.
-Meanwhile, the annotation of molecules demands expertise from chemists and can sometimes require wet lab experiments, which are both expensive and time-consuming.
-In contrast, TOMG-Bench, as an open-domain generation task, does not rely on human annotations for construction.
-Instead, we focus on basic molecule structural properties and molecule operations that could be validated by chemical toolboxes, which makes it easy to increase the volume of data, thereby resolving data hunger in chemical corpora.
-In TOMG-Bench, three major tasks consist of nine subtasks in total. Each subtask is allocated 5,000 test samples. Figure 4 illustrates the data construction process.
-For MolEdit and MolOpt, we sample molecules from a specific molecule database.
-We mainly consider two molecule databases: Zinc-250K (Sterling & Irwin, 2015), and PubChem  (Kim et al., 2019).
-Zinc-250K has 250,000 molecules, while PubChem has about 10 million molecules.
-Therefore, the data size of Zinc-250K is more suitable for us to build the test sets, while PubChem could be further sampled to generate the instruction tuning dataset.
-After sampling, we utilize RDKit (Landrum, 2013), a molecular informatics toolbox, to collect basic molecule statistics, including the structural patterns and chemical properties like LogP, MR, and QED values, which can then be integrated into our pre-defined task prompts.
-For the MolCustom task, we randomly generate 5,000 instructions as requests for each subtask, requiring different numbers and collections of atoms, bonds, and functional groups.
-Further details are provided in the Appendix D.
-3.3 Metric Design
-The evaluation of the TOMG-Bench is facilitated through a set of carefully designed metrics tailored to the specific tasks within the benchmark.
+    We provide insightful findings based on our extensive benchmarking of 28 LLMs, revealing the limitations of existing targeted generation datasets and highlighting the potential for LLMs to transition from simple pattern recall to realistic molecular design.
+2 Speak-to-Structure
+In this section, we detail the design philosophy, technical composition, and statistics of Speak-to-Structure (S2-Bench), which is fundamentally structured around the core capabilities required for real-world molecular design: editing, optimizing, and customized generation.
+2.1 Task Composition
+S2-Bench is meticulously designed to move beyond simple pattern-matching and evaluate the genuine molecular understanding and generation capabilities of LLMs.
+We adopt a one-to-many paradigm, where a single language instruction can be fulfilled by multiple valid molecules.
+Our benchmark challenges LLMs to demonstrate a flexible understanding of molecular syntax rather than mere memorization.
+This design philosophy is concretely realized through three tasks of increasing complexity, each mirroring a critical phase of molecule discovery.
+MolEdit assesses an LLM’s foundational molecular syntax and structural modification capabilities.
+In this task, a base molecule is provided, and the model is asked to perform a specific change on the molecular structure.
+In a real-world scenario, this task is analogous to the iterative lead optimization process in drug discovery, where chemists make small, precise changes to a candidate molecule to change its properties.
+We designed three subtasks to probe this capability: AddComponent and DelComponent test the model’s ability to add or remove a specific functional group, respectively, while SubComponent challenges it to perform a combined operation of both.
+MolOpt evaluates an LLM’s ability to perform goal-oriented chemical reasoning.
+The challenge here is twofold: the LLM should not only edit a molecule but also ensure that the modification leads to a desired change in a specific property.
+This task directly mirrors the hit-to-lead optimization stage, where molecules are refined to achieve better pharmacological profiles.
+To assess this, we chose three key properties (LogP, MR, and QED), which are vital for drug discovery and can be reliably and reproducibly calculated.
+By requiring the model to generate a molecule that improves on a given property, we test its ability to understand the complex relationship between molecular structure and function, a capability that cannot be evaluated by tasks with a single, predetermined answer.
+MolCustom serves as the ultimate test of an LLM’s creative chemical design ability.
+MolCustom requires the LLM to generate a novel molecule from a natural language description of its desired structural components.
+This task is not about modifying an existing structure but about synthesizing a new one from scratch, akin to a chemist defining the needs for a new compound at the very beginning of a project.
+To ensure a clear and precise evaluation of this complex task, we focus on generating molecules with specific structural features: a defined number of atoms (AtomNum), a specified number and type of bonds (BondNum), or the inclusion of particular functional groups (FunctionalGroup).
+While these constraints may seem simple, they are deceptively challenging, requiring the model to apply a sophisticated understanding of chemical valence rules to create a valid and novel molecule.
+Compared to targeted molecule generation datasets, these tasks in S2-Bench, while seemingly straightforward in their instructions, impose a more rigorous requirement on LLMs’ ability to perform precise, open-domain molecule generation.
+Our benchmark fundamentally shifts the evaluation from pattern recall to practical chemical design, thereby providing a more accurate measure of an LLM’s potential in molecule discovery and helping to build more explainable and trustworthy models.
+2.2 Data Construction
+Previously, the development of robust datasets for text-based targeted molecule generation has been hindered by the scarcity of high-quality human annotations.
+For example, while text-to-image generation datasets like MS COCO (Chen et al., 2015) contain millions of annotated samples, molecule-caption datasets, such as ChEBI-20 (Edwards et al., 2022) are significantly smaller, often by orders of magnitude.
+This disparity arises because molecular annotation demands specialized expertise, making it both time-intensive and costly.
+The resulting data scarcity further poses a significant challenge for advancing LLMs in text-guided molecule discovery.
+Our approach to data construction fundamentally bypasses this bottleneck.
+S2-Bench is designed for open-domain molecule generation tasks that do not rely on human annotations.
+Instead, we leverage automated chemical toolkits to programmatically construct tasks and evaluate outputs based on objective molecular properties and structural rules.
+This enables us to generate a virtually unlimited volume of data, effectively addressing the “data hunger” that has constrained previous research.
+More importantly, this programmatic approach allows us to create tasks with a one-to-many relationship, where multiple correct answers exist for a single prompt, which better aligns with the complexities of drug discovery, accommodating its inherent variability and diverse outcomes.
+Specifically, we design a systematic and scalable data construction process, as illustrated in Figure 4.
+For MolEdit and MolOpt, we sample molecules from large, publicly available chemical databases.
+We selected Zinc-250K (Sterling & Irwin, 2015) for building our test set due to its manageable size and diversity, while the massive PubChem (Kim et al., 2019) database (with over 10 million molecules) serves as the basis for the instruction-tuning dataset.
+We utilize the RDKit (Landrum, 2013) toolbox to automatically extract key molecular statistics, including structural patterns and properties like LogP, MR, and QED.
+These extracted features are then integrated into our pre-defined, instruction-based prompt templates, which is demonstrated in Appendix D.
+For MolCustom, we move beyond existing databases and programmatically generate instructions to test a model’s ability to create novel molecules.
+For each of the three subtasks, we randomly generate 5,000 instructions that specify a target number and type of atoms, bonds, or functional groups.
+Furthermore, for each subtask, we pre-defined a diverse prompt template pool to ensure that the LLMs are not overfit to a limited set of prompt formats.
+This programmatic construction not only allows for the generation of a much larger data volume but also ensures that our benchmark tests an LLM’s intrinsic chemical reasoning ability, rather than its capacity for memorizing human-annotated patterns.
+By moving beyond the one-to-one paradigm, S2-Bench provides a more robust and realistic evaluation, paving the way for models that can truly innovate and design novel molecules.
+2.3 Evaluation
+The evaluation of S2-Bench is facilitated through a set of carefully designed automated evaluation processes and metrics tailored to the unique nature of our tasks.
+Unlike one-to-one datasets or benchmarks that simply check for exact matches, our metrics are designed to assess an LLM’s ability to produce valid, relevant, and novel outputs within an open-ended framework.
 Task (Weighted) Success Rate Similarity Novelty Validity
@@
 2.3.1 MolEdit & MolOpt Evaluation
-For the MolEdit and MolOpt tasks, as shown in Table 1, the following metrics are adopted:
-Success Rate: We have developed automated testing processes with RDKit, which compare the generated molecule with the original molecule and check if the number of the specified functional group is correct after modification (for MolEdit task) and if the specified chemical property is optimized as required (for MolOpt task).
-Details of the automated testing process are shown in Appendix E.
-The success rate reflects the percentage of molecules that have successfully passed the rigorous testing processes.
-However, as it is impossible to track every change between the generated molecules and the original molecule (See Appendix E for detailed explanation), successfully passing the test processes does not necessarily mean that the generated molecule is modified from the original molecule.
-Similarity: In this case, we need to evaluate the similarity between the original molecule and the generated molecule.
-When removing a specified functional group, the modification should follow the minimal change principle.
-Therefore, a high similarity indicates that the generated molecule can be modified from the original molecule via as few as possible steps, instead of transforming the molecule into an entirely new one or inadvertently removing additional components.
-For implementation, we calculate the Tanimoto Similarity between the generated molecule and the original molecule by transforming them into the Morgan Fingerprints (Butina, 1999), which can be described by the following equation:
+For the MolEdit and MolOpt tasks, as shown in Table 1, we employ a combination of metrics to assess both the correctness of the modification and the rationality of the design.
+Success Rate: This metric evaluates a model’s ability to fulfill the specific molecule generation requirements.
+We design automated evaluation processes to verify if the generated molecule meets the specified criteria (e.g., correct functional group modification for MolEdit, or desired property optimization for MolOpt).
+Detailed Implementations of the automated testing process are shown in Appendix E.
+This metric examines whether the LLM can follow instructions precisely, which is a fundamental requirement for an LLM to serve as a chemist assistant.
+Similarity: In open-domain tasks like MolEdit and MolOpt, a high Success Rate alone is insufficient.
+We must also ensure that the generated molecule is a reasonable modification of the original, rather than a completely new, unrelated structure that happens to satisfy the criteria.
+To address this, we measure the Tanimoto Similarity between the generated and original molecules using Morgan Fingerprints (Butina, 1999).
+A high similarity score indicates that the model has performed a rational, localized edit and has not simply generated a different molecule from scratch.
+The similarity $\delta(m^{g},m^{o})$ between the generated molecule $m^{g}$ and the original molecule $m^{o}$ is calculated as:
 $\delta(m^{g},m^{o})=\frac{\left|fp_{m^{g}}\cap fp_{m^{o}}\right|}{\left|fp_{m^{g}}\cup fp_{m^{o}}\right|},$ (1)
-where $m^{g}$ denotes the generated molecule, $m^{o}$ is the original molecule, and $\delta(m^{g},m^{o})$ is the Tanimoto Similarity between the generated molecule and the original molecule, while $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints.
+where $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints.
 $\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
-Then, the similarity score for a subtask $t$ in MolEdit and MolOpt, denoted as $\delta_{t}$, is determined by taking the average of the similarity scores from all individual test samples.
 Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax.
-We use RDKit to parse the SMILES strings.
-Only the molecules that successfully pass the parser are considered valid molecules.
-A higher validity means that the model is more familiar with the grammar of molecule textual representations.
+Only the molecules that successfully pass the SMILES parser are considered valid molecules.
+A higher validity means that the model is more familiar with the molecule syntax.
 2.3.2 MolCustom Evaluation
-For the MolCustom task, the metrics adopted are also shown in Table 1.
-As LLMs are required to generate the molecule from scratch, we employ a different testing process for the evaluation of success rate and introduce the novelty score instead of similarity:
-Success Rate: For different subtasks in MolCustom, we have designed different testing processes by the chemical toolbox.
-We calculate the number of the specified atoms, functional groups, or bonds for the three subtasks to check if the generation meets the requirements.
+For the MolCustom task, where models are required to generate molecules from scratch based on structural descriptions, we use a different set of metrics to evaluate their creative design capabilities.
+Success Rate: This metric measures how well the generated molecules adhere to the high-level structural constraints provided in the prompt (e.g., number of atoms, bonds, or specific functional groups).
+As the task requires de novo generation, the success rate directly reflects a model’s ability to translate abstract requirements into a valid, concrete molecular structure.
 Detailed testing process of MolCustom could also be found in Appendix E.
-A high success rate indicates that the LLM is effectively generating molecules according to the request.
-Novelty: However, since there may be numerous valid solutions, fulfilling the basic requirements alone does not guarantee the quality of generation.
-To enable the exploration of new molecule structures, we employ novelty as an indication of the quality, which compares the similarity between the newly generated molecules and existing ones in the database.
-The novelty score is particularly important for discovering potential drug or material candidates, where innovation can lead to significant advancements.
-For implementation, we choose the Zinc250K as the database and calculate the average Tanimoto Similarity between the generated molecules and the existing molecules.
-The novelty $n$ for the generated molecule $m^{g}$ can be calculated as follow:
-$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc250K}\delta(m^{g},m^{k})}{\left|Zinc250K\right|},$ (2)
-Then, the novelty score forthe subtask $t$ in MolCustom, denoted as $n_{t}$, is determined by taking the average of the novelty scores from all individual test samples.
-Validity: Similarly, validity is also an important metric for the MolCustom task.
+Novelty: For open-domain generation, fulfilling the requirements is just the first step.
+The true value lies in a model’s ability to generate novel and innovative molecules.
+The novelty score quantifies this by comparing the generated molecules against existing structures in a large database, such as Zinc-250K (Zinc for short).
+A low similarity to known molecules suggests a high degree of novelty, which is a critical indicator of a model’s potential for discovering truly new molecule structures.
+The novelty $n$ for the generated molecule $m^{g}$ can be calculated as:
+$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc}\delta(m^{g},m^{k})}{\left|Zinc\right|},$ (2)
+Validity: As with the other two tasks, validity ensures that all generated molecules are chemically sound and can be interpreted correctly.
 2.3.3 Average Weighted Success Rate
-To comprehensively evaluate the average performance of LLMs on TOMG-Bench, we introduce an average weighted success rate to rank the performance of LLMs.
-Considering that the success rate alone can not ensure the accurate generation for MolEdit and MolOpt tasks and the novelty score for MolCustom is also a critical metric for the quality of molecule generation, we adopt similarity and novelty scores as weights of the success rate:
+To provide a single, comprehensive ranking of LLM performance on S2-Bench, we introduce a weighted success rate.
+This metric combines the core success rate with a quality metric relevant to each task: Similarity for MolEdit/MolOpt and Novelty for MolCustom.
+This approach ensures that a high score reflects not only the ability to follow instructions but also the rationality and creativity of the generation.
+The weighted success rate for a subtask $t$ is defined as:
 $W\!S\!R_{t}=\begin{cases}n_{t}\times S\!R_{t},&t\in\{MolCustom\}\\ \delta_{t}\times S\!R_{t},&t\in\{MolEdit,MolOpt\}\end{cases},$ (3)
@@
 $\overline{W\!S\!R}=\frac{1}{9}\sum_{t}W\!S\!R_{t},$ (4)
-where the average weighted success rate $\overline{W\!S\!R}$ provides a balanced measure of performance that considers both the success rate and quality of the generation.
+where the average weighted success rate $\overline{W\!S\!R}$ provides a balanced measure of a model’s performance across all nine subtasks, offering a robust and nuanced assessment of its capabilities in open-domain natural language-driven molecule generation.
 2.4 OpenMolIns: Instruction Tuning Dataset
-In this section, we introduce OpenMolIns, a specialized dataset derived from the PubChem database to help LLMs get familiar with text-based open molecule generation via instruction tuning.
-This instruction tuning dataset is meticulously designed to ensure that the molecules it contains do not overlap with those in the Zinc-250K dataset to promote the generation of more novel molecular structures and avoid any potential data leakage that could compromise the integrity of the model performance.
-We collect the instruction tuning dataset by constructing the samples of the nine subtasks in equal amounts.
-The RDKit toolbox is applied to construct the instruction tuning dataset.
+To effectively train and evaluate LLMs on the open-ended challenges posed by S2-Bench, we introduce OpenMolIns, a specialized instruction-tuning dataset derived from the PubChem database.
+The core philosophy behind OpenMolIns is to provide models with the kind of flexible, one-to-many training examples they need to learn genuine chemical reasoning, rather than simply memorizing specific input-output pairs.
+To ensure the integrity of our evaluations, we meticulously designed OpenMolIns to have zero overlap with the Zinc-250K, preventing any data leakage that could compromise the validity of our benchmark results.
+We built the instruction-tuning dataset programmatically, creating equal numbers of samples (i.e., 5,000) for all nine subtasks.
+The RDKit toolbox was essential to this process, allowing us to automatically generate data based on objective molecular properties and rules.
+This contrasts sharply with previous methods that rely on scarce and costly human annotations.
 Details of the construction process and the pre-defined prompt templates are shown in Appendix D.
-For the MolEdit and MolOpt domains, we randomly select a functional group from the original molecule for addition or removal.
-Then, similarly, we wrap the original molecule and the edited molecule with a random pre-defined prompt to construct the training sample for MolEdit.
-In the case of MolOpt, we also evaluate the direction of the desired property changes by calculating the properties to determine whether the operation improves or decreases the property value.
-Then, the original molecule, the optimized molecule, and the optimization direction are altogether wrapped by a pre-defined prompt.
-In the MolCustom domain, constructing the instruction tuning samples is rather straightforward.
-We calculate the molecular statistics and encapsulate them within prompts.
-For example, for AtomNum, we count all the atoms we would consider in the molecule, including their types and numbers.
-Then, we wrap these statistics with a random pre-defined prompt to construct the training sample.
-This method allows the generated molecules to better fit the real distribution of molecular space.
-To investigate the impact of data scales on the performance of LLMs, we have established five distinct data levels: light, small, medium, large, and xlarge, as illustrated in Section 3.5.
-Each level represents a different quantity of data, shown in Table 2, allowing researchers to analyze how the amount of training data influences the model’s ability to generate and edit molecules effectively.
-Dataset TOMG-Bench OpenMolIns
+For MolEdit and MolOpt, our goal is to improve the capabilities of LLMs to perform rational, goal-oriented modifications.
+For MolEdit, we start with a molecule, programmatically identify a modifiable functional group, and then perform a change (add, delete, or substitute) to create a new, valid molecule.
+This pair of (original and modified) molecules is then used to construct an instruction.
+This process teaches the model not only a single correct answer, but rather the skill of modifying a molecule in a chemically sound way.
+Similarly, for MolOpt, we calculate properties before and after a modification to identify a specific optimization direction (e.g., increasing LogP or decreasing MR).
+The training sample then explicitly links a desired outcome to a structural change, teaching the LLM to reason about the relationship between structure and property.
+In the MolCustom domain, we aim to train the model to generate molecules from scratch based on structural details.
+Instead of relying on existing molecule-caption pairs, we programmatically analyze molecules from PubChem, extract key structural statistics (e.g., atom counts, bond numbers, and functional group types), and use these as the basis for a training prompt.
+The molecule itself becomes the target output.
+This approach trains the model to synthesize a new molecule that fits a set of high-level, multi-faceted criteria, a skill fundamentally different from recalling a specific, pre-existing structure.
+This method allows the generated molecules to better fit the real distribution of molecular space and prepares the model for de novo design.
+To further investigate the impact of data scales on an LLM’s ability to learn these tasks, we created five distinct data levels: light, small, medium, large, and xlarge, shown in Table 2.
+This tiered structure allows researchers to systematically analyze the data scaling law for open-domain natural language-driven molecule generation, providing valuable insights into how data quantity influences a model’s capacity for chemical reasoning.
+Dataset S2-Bench OpenMolIns
 item each subtask total light small medium large xlarge
 Data Size 5,000 45, 000 4,500 18,000 45,000 90,000 1,200,000
-Table 2: Statisics of TOMG-Bench and OpenMolIns.
+Table 2: Statisics of S2-Bench and OpenMolIns.
 2.5 Statistics
-In this section, we outline the data sizes of TOMG-Bench along with the OpenMolIns dataset, which are shown in Table 2.
-For TOMG-Bench, we have three main tasks with nine subtasks in total, where each subtask contains 5,000 test samples.
-OpenMolIns has five distinct data scales: light, small, medium, large, and xlarge, ranging from 4,500 to 1,200,000 examples, which helps us investigate the data scaling law of applying LLMs to the text-based open molecule generation task.
-The nine subtasks within the TOMG-Bench are uniformly distributed in OpenMolIns.
+In this section, we provide a detailed overview of S2-Bench and OpenMolIns, as summarized in Tables 2 and 3.
+The design and scale of our datasets are not arbitrary; they are meticulously crafted to enable a rigorous and nuanced evaluation of LLMs in open-domain molecule generation, which previous, smaller datasets cannot provide.
+S2-Bench is composed of nine subtasks, with each containing 5,000 test samples, for a total of 45,000 carefully curated test cases.
+This extensive size is critical for providing a statistically robust and comprehensive assessment of model performance, mitigating the risk of misleading results that can arise from smaller test sets.
+More importantly, these samples are designed to test the model’s ability to handle one-to-many relationships, a challenge that is central to true chemical reasoning.
+Meanwhile, to support the development of models capable of succeeding on this benchmark, OpenMolIns is provided in five distinct data scales, ranging from 4,500 to 1,200,000 examples.
+This tiered structure serves a dual purpose: it not only offers ample data to train powerful models but also allows researchers to systematically investigate the data scaling law for this novel task.
+By observing how model performance improves with increasing data size, we can gain valuable insights into the data requirements for aligning molecular space with natural language.
+Benchmarks or Datasets Task Type # Training Data # Test Data Public Access
+PCdes(Zeng et al., 2022) Targeted Generation 10,500 3000 $\times$
+ChEBI-20 (Edwards et al., 2021) Targeted Generation 26,407 3,000
+PubChem (Liu et al., 2023) Targeted Generation 12,000 2,000
+Mol-Instructions (Fang et al., 2024) Targeted Generation 297,319 1000
+L + M - 24 (Edwards et al., 2024) Targeted Generation 160,492 21,839
+S2-Bench & OpenMolIns Open Generation 1,200,000 45,000
+Table 3: Comparison with existing text-based molecule generation benchmarks and datasets. Notably, we denote the statistics of the description guided molecule design part in Mol-Instructions.
+Besides, as shown in Table 3, S2-Bench and OpenMolIns stand out from existing text-based molecule generation datasets.
+S2-Bench is the first to introduce a truly open-domain natural language-driven molecule generation task, moving the field beyond the limitations of one-to-one mapping.
+In terms of data scale, our datasets are also significantly larger.
+OpenMolIns has the largest training set to date, with a volume that is unprecedented in this domain.
+This scale is crucial for enabling LLMs to learn complex chemical rules and generalize beyond the memorization of specific examples.
+Similarly, S2-Bench’s 45,000 test samples make it the largest evaluation benchmark, providing a more reliable and comprehensive measure of a model’s true capabilities.
+This combination of innovative task design and large-scale data represents a new paradigm for evaluating and advancing LLMs in the field of text-guided molecular discovery.
 3 Experiments
@@
 This category includes LLMs that are only accessible via commercial API services.
-In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful language understanding and generation capabilities.
+In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful reasoning and generalization capabilities.
 These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
@@
 This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications.
-Specifically, we benchmark Llama-3-70B-Instruct, Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), and chatglm-9B (GLM et al., 2024).
-These LLMs are top-tier open source general LLMs within our computation budget.
+Specifically, we benchmark Llama3-70B-Instruct, Llama3-8B-Instruct, Llama3.1-8B-Instruct, Llama3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Deepseek-R1-distill-Qwen-7B (Guo et al., 2025), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), chatglm-9B (GLM et al., 2024), and Gemma3-12B (Team et al., 2025).
+These LLMs are top-tier open-source general LLMs within our computation budget.
 Open-source ChEBI-20 Fine-tuned LLMs.
@@
 These models have been the state-of-art models in the molecule-caption translation task.
-OpenMolIns Fine-tuned LLMs We further adopt LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama-3.1-8B-Instruct on OpenMolIns dataset for instruction tuning.
-We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M.
-We select Galatica-125M because it has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023).
+Here, to ensure a fair evaluation of the intrinsic capabilities of LLMs, we exclude models that utilize multi-modal architectures or retrieval-augmented generation.
+OpenMolIns Fine-tuned LLMs.
+We further fine-tune LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama3.1-8B-Instruct on OpenMolIns dataset for comparison.
+We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M because the model has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023).
 Meanwhile, a small size model can also help us study the data scaling law within reasonable budget.
-Besides, Llama3.2-1B-Instruct and Llama-3.1-8B-Instruct are selected due to their advanced capabilities and similar architectures.
-Figure 3: The performance of models benchmarked in TOMG-Bench.
-In TOMG-Bench, LLMs are divided into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs.
+Furthermore, Llama3.2-1B-Instruct and Llama3.1-8B-Instruct are also selected due to their advanced capabilities and similar architectures.
+Figure 2: The performance of LLMs benchmarked in S2-Bench.
+LLMs fall into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs.
 Models of unknown parameters are denoted as horizontal lines.
 3.2 Findings
-Figure 3 presents the performances and model sizes of different models benchmarked on TOMG-Bench, as well as the instruction-tuning performance of Galactica-125M on the five distinct data levels of OpenMolIns.
-More precise details for each subtask are further illustrated in Appendix F.
-Based on the above results, we observe the following key findings:
-Text-based open molecule generation is both promising and challenging for LLMs.
-As illustrated in Figure 3, we have calculated the average weighted success rate across all the nine subtasks.
-Among the LLMs benchmarked, Claude-3.5 stands out as the top performer, achieving an average weighted success rate of 35.92%.
-Gemini-1.5-pro follows closely with an average weighted success rate of 34.80%.
-Meanwhile, fine-tuned on OpenMolIns-large, Llama-3.1-8B-Instruct achieves an average weighted success rate of 27.22%, outperforming all the open-source general LLMs.
-However, these results still underscore the considerable scope for improvement, even among the most advanced proprietary LLMs.
-What’s more, despite LLMs like Claude-3.5 and GPT-4o exhibit relatively strong performance in the MolEdit and MolOpt tasks, the more challenging MolCustom task still remains a challenge.
-In MolCustom tasks, no LLM has managed to achieve a success rate exceeding 25% for a single subtask.
-Meanwhile, as shown in Figure 5, among the three subtasks of MolCustom, BondNum is the most difficult task, as it requires LLMs to understand the connections or edges in the molecule graph merely from the textual representation.
-These observations indicate that the generation of molecules from scratch demands a deep understanding of the molecular structures, an area where current models are still striving to make significant strides.
-Most powerful open-source general LLMs can already outperform GPT-3.5-turbo.
-In TOMG-Bench, Llama-3-70B-Instruct has achieved an impressive average weighted success rate of 23.93%, notably outperforming GPT-3.5-turbo, which scored 18.58%.
-Compared to the proprietary models, open-source general LLMs seem to have been lack of chemical corpora for the pre-training process.
-However, recent development of open-source general LLMs have rapidly bridged the gap.
-The evolution of the Llama series, in particular, has been remarkable, finally surpassing GPT-3.5-turbo and showing the increasing diversity of pre-training corpora.
-More powerful LLMs usually inherit a better performance in TOMG-Bench.
-Across all the LLMs we benchmarked, a clear trend emerged: the more powerful the LLM is, the higher performance it can achieve on the TOMG-Bench.
-For instance, the GPT series has consistently demonstrated improved performance with each new iteration from GPT-3.5-turbo to GPT-4o.
-Similarly, within the Llama-3 series, we could also observe that larger and more powerful models tend to achieve superior results on the TOMG-Bench.
-These findings underscore a strong correlation between an LLM’s capabilities and its performance in text-based open molecule generation.
-However, we encountered an unexpected anomaly with certain open-source LLMs.
-Notably, Qwen2-7B-Instruct, despite its impressive ability to solve mathematical problems and its size of 7 billion parameters, underperformed models with as few as 1 billion parameters.
-This result is particularly striking and suggests that the TOMG-Bench offers a unique and comprehensive evaluation that current races for LLMs may not have adequately addressed.
-This discovery also highlights the significance of the TOMG-Bench as a new benchmark for LLMs.
-It provides a broader and more diverse assessment that exposes potential blind spots in the development of LLMs.
-ChEBI-20 is insufficient for LLMs to master molecular structures and editing operations.
-The ChEBI-20 dataset is designed to bridge the gap between molecules and textual descriptions.
-Despite this intention, LLMs trained on ChEBI-20 have demonstrated limited effectiveness in TOMG-Bench.
-For instance, BioT5-base, which is claimed as the state-of-the-art model for text-based molecule generation on the ChEBI-20 dataset, only achieves an average weighted success rate of 4.21% on the TOMG-Bench.
-In the MolEdit and MolOpt tasks, although these models obtain a high success rate, the generations are actually not modified from the original molecules, leading to poor similarity scores.
-This also shows that the average weighted success rate provides a more accurate reflection of the actual performance.
-Similarly, in the MolCustom task, which closely mirrors the text-based molecule generation task, the performance remains unsatisfactory, with no model achieving a success rate above 5% in a single subtask.
-This performance shortfall highlights a critical limitation of the ChEBI-20 dataset, as it lacks the data quantity and diversity necessary to effectively align molecules with textual descriptions.
-The data scaling law varies among different tasks.
-Across five distinct data scales of OpenMolIns and the experiments on Galactica-125M, we observe a general data scaling law: as the size of the corpus increases, the performance of LLMs also improves.
-As shown in Figure 3 and Table 12, the outcomes were remarkable: Galactica-125M achieved a weighted average score of 25.73% on OpenMolIns-xlarge, surpassing even the 70B Llama-3-70B-Instruct and GPT-3.5-turbo, with only 125 million parameters.
-However, an interesting observation shows that the data scaling law varies among different tasks.
-For instance, in MolCustom tasks, we witnessed a substantial average improvement of 286% when we scaling the data size from large to xlarge, which suggests that, for specific tasks, increasing data size has a beneficial effect on model performance.
-Conversely, in the MolCustom tasks, the performance gains from data scaling were notably modest.
-More concerning, in the MolEdit tasks, we even observed a decline in performance with increased data size, which indicates that the model’s capacity is not solely data-driven but is also constrained by the number of parameters.
-It suggests that for certain subtasks within TOMG-Bench, simply increasing the volume of data may not lead to proportional improvements in performance, and other factors such as model complexity or optimization strategies may need to be considered.
+Based on our comprehensive benchmarking, we have made the following key observations regarding the capabilities of LLMs in open-domain natural language-driven molecule generation.
+F1: Current LLMs show promise but lack the structural understanding necessary for precise molecule generation.
+As shown in Figure 2, top-tier proprietary models such as Claude-3.5 and Gemini-1.5-Pro attain 35.92% and 34.80% in average weighted success rate, respectively.
+These results suggest that general LLMs hold promise for open-ended molecule generation due to their powerful reasoning capability, yet substantial challenges remain in translating textual requirements into chemically valid and structurally precise molecules.
+Notably, within the MolCustom task, no LLM manages to surpass a 25% weighted success rate on any individual subtask.
+The particularly poor performance on MolCustom tasks highlights that while current LLMs can generate chemically plausible molecules in a broad sense, they struggle with satisfying fine-grained structural constraints such as precise atom, bond, or functional group counts, because the precise control of molecular structures requires reasoning over discrete, globally coupled constraints.
+LLMs, even trained on SMILES, primarily learn local token distributions from large-scale pre-training data, which rarely emphasizes exact numeric constraints.
+The precise control, in fact, is crucial for molecular discovery, as many properties of interest, such as solubility, toxicity, and active site composition, depend on specific structural configurations.
+Our finding suggests that open-domain natural language–driven molecule generation task can serve as an effective diagnostic tool, revealing systematic weaknesses in current LLMs and providing rich failure cases that can inform future instruction tuning or reinforcement learning strategies to improve molecule structural constraint adherence.
+F2: While pre-training provides a foundation for LLM performance, instruction tuning is indispensable for guiding and optimizing their capabilities in molecular discovery.
+Open-source general LLMs can often lack chemical corpora during their pre-training stage.
+As a result, models like Qwen2-7B-Instruct and yi-1.5-9B obtained unsatisfactory results.
+The observed shortcoming can be attributed to the lack of chemical data in pre-training, despite their notable proficiency in mathematics and general language understanding.
+In contrast, Llama3.1-8B-Instruct delivers superior performance at a relatively modest model scale, demonstrating a more balanced and comprehensive capability across diverse domains.
+Meanwhile, our experiments with OpenMolIns highlight the effectiveness of instruction tuning.
+Fine-tuned on the largest OpenMolIns scale (i.e., xlarge), Llama3.1-8B-Instruct surpassed all other LLMs with an average weighted success rates of 39.33%.
+More remarkably, the much smaller model, Galactica-125M, within just 125 million parameters, achieved a weighted average success rate of 25.73%, outperforming models two orders of magnitude larger, such as Llama3-70B-Instruct.
+This remarkable outcome indicates that task-specific instruction tuning on a large, high-quality corpus can be highly effective, suggesting a practical avenue for LLMs to efficiently narrow the performance gap in molecular discovery.
+F3: One-to-one mappings between language instructions and molecules make LLMs inherently rely on pattern recognition and recall.
+Our results also substantiate the observation, demonstrating that existing one-to-one mappings are insufficient for evaluating true chemical comprehension of LLMs.
+Notably, ChEBI-20 Fine-tuned LLMs, which are designed to align molecules with texts, perform worse than general LLMs.
+For instance, BioT5-base (Pei et al., 2023), a state-of-the-art model on the ChEBI-20 dataset, achieves an average weighted success rate of only 4.21% on S2-Bench.
+A closer examination of BioT5-base’s performance provides insight into this limitation.
+Although the model achieves relatively high success rates on MolEdit and MolOpt tasks, its generated molecules often exhibit low similarity to the original structures.
+This indicates that the model is not performing rational edits; rather, it generates different but valid molecules that happen to satisfy the criteria.
+In other words, BioT5-base relies more on recalling patterns from its training corpus than on learning to systematically modify molecular structures.
+Consequently, while the ChEBI-20 dataset is valuable for molecule–caption translation, its limited diversity, scale, and one-to-one mapping nature are insufficient for training LLMs to perform nuanced, open-ended molecular generation.
+Figure 3: Task-specific performance scaling with increasing data in S2-Bench. (a) Galactica-125M. (b) Llama-3.1-8B-Instruct.
+F4: The data scaling law does not always hold: task-specific bottlenecks limit gains in molecule optimization and editing.
+As shown in Figure 3, our experiments with Galactica-125M across five data scales of OpenMolIns provide new insights into the data scaling law for molecule generation.
+Overall, we observe that larger datasets generally improve performance, but the effect is highly task-dependent.
+In MolCustom, scaling the dataset from large to xlarge yields a remarkable 286% average performance gain, indicating that tasks involving complex, de novo synthesis strongly benefit from larger and more diverse training data.
+By contrast, Galactica-125M obtains only modest gains in MolOpt, and even negative gains in MolEdit, suggesting that additional data alone cannot overcome existing bottlenecks in these relatively simpler tasks.
+Interestingly, when examining Llama-3.1-8B-Instruct, we find that performance continues to increase when scaling from the large to xlarge data level, except for LogP and MR subtasks.
+This contrast generally highlights that small models like Galactica-125M are capacity-limited, whereas larger models can more effectively leverage additional data.
+To summarize, these findings suggest that while complex, open-ended tasks are data-driven, simpler property optimization or editing tasks may be constrained more by model capacity than by dataset size.
+Thus, merely enlarging training data may not yield proportional benefits.
+Instead, future progress will require a balanced consideration of both model scale and data scale.
 4 Conclusion
-In this study, we introduce TOMG-Bench, the first benchmark designed to assess the capabilities of LLMs in the realm of open-domain molecule generation.
-Benchmarking 25 LLMs, TOMG-Bench highlights the limitations of existing targeted molecule generation tasks and demonstrates the potential of general LLMs in this domain.
-Additionally, we propose an instruction tuning dataset, OpenMolIns, to fully explore the potential of LLMs on the TOMG-Bench, enabling Llama-3.1-8B-Instruct to outperform GPT-3.5-turbo.
-Our contributions not only lie in the development of a novel benchmark for LLMs in molecular discovery but also provide a diverse indication of the LLM evaluation.
-Appendix A Comparison with Existing Text-based Molecule Benchmarks or Datasets
-Benchmarks or Datasets Task Type # Training Data # Test Data Public Access
-PCdes(Zeng et al., 2022) Targeted Generation 10,500 3000 $\times$
-ChEBI-20 (Edwards et al., 2021) Targeted Generation 26,407 3,000
-PubChem (Liu et al., 2023) Targeted Generation 12,000 2,000
-Mol-Instructions (description guided molecule design) (Fang et al., 2024) Targeted Generation 297,319 1000
-TOMG-Bench & OpenMolIns Open Generation 1,200,000 45,000
-Table 3: Comparison with existing text-based molecule generation benchmarks or datasets.
-Table 3 shows the comparison between TOMG-Bench & OpenMolIns and the existing text-based molecule generation benchmarks or datasets.
-TOMG-Bench introduces the brand-new open-ended molecule generation task for LLMs, which boosts the diversity of text-based molecule generation task.
-Compared to the existing benchmarks or datasets, OpenMolIns has the largest training set, which could better aligns the molecular space with the natural language space.
-Meanwhile, TOMG-Bench owns three major tasks and nine subtasks with 45,000 test samples in total, whose test data size is also the largest one among all the existing benchmarks and datasets, enabling more accurate evaluation and comprehensive assessment of LLMs’ performance in text-based molecule generation.
+In this study, we introduced S2-Bench and OpenMolIns, the first benchmark and instruction tuning dataset for evaluating the capabilities of LLMs in open-domain natural language-driven molecule generation.
+By moving beyond traditional one-to-one text-to-molecule mappings, our benchmark accommodates the open-ended nature of molecular discovery.
+S2-Bench focuses on the realistic molecular reasoning and design, rather than simple pattern matching and memorization.
+Our comprehensive benchmarking of 28 LLMs not only highlights the challenges faced by current models and the limitations of existing targeted molecule generation datasets, but also demonstrates the substantial potential of LLMs for natural language–driven molecule discovery.
+Appendix A Related Work
+Molecule discovery is a cornerstone of scientific progress, underpinning advances in both drug development and material design (Du et al., 2022).
+The integration of artificial intelligence into this process has driven a paradigm shift in the pharmaceutical landscape, substantially improving the efficiency and accuracy of identifying and developing novel therapeutic candidates.
+Recent breakthroughs in natural language processing (NLP) have highlighted the potential of Large Language Models (LLMs) to analyze complex biological and chemical data more effectively than traditional computational methods (Zhou et al., 2023).
+Within this context, Text-based Molecule Generation has emerged as a representative task for AI-driven molecule discovery (Edwards et al., 2021).
+This task focuses on generating target molecules directly from natural language descriptions, which requires the construction of paired datasets linking molecular structures with their textual representations.
+Early work in this direction employed transformer-based models such as MolT5 (Edwards et al., 2022), which leveraged large-scale self-supervised learning to produce high-quality SMILES strings from textual prompts.
+Building on this foundation, models like KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023) integrated molecular graphs and biochemical texts to jointly enhance molecular understanding and generation.
+More recently, 3D-MoLM (Li et al., 2024d) advanced the field by incorporating spatial configurations, thereby improving the geometric validity and accuracy of generated molecular structures.
+Parallel to these architecture-driven improvements, LLMs such as MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) have demonstrated strong in-context learning capabilities, enabling adaptive molecule generation by retrieving and utilizing relevant examples from the provided context.
+Besides, MolReFlect (Li et al., 2024c), emphasizes fine-grained alignment between molecular structures and natural language descriptions through a teacher–student paradigm, which effectively captures subtle structural–semantic relationships.
+Moreover, Mol-R1 (Li et al., 2025) takes a step further by introducing DeepSeek-R1 like Long-CoT reasoning in molecule generation, enabling robust and interpretable natural language-driven molecule generation.
 Appendix B Implementation Details
@@
 Appendix C Hyper Parameters
-In this section, we illustrate the detailed parameters adopted in this work, as shown in Table 4.
+[…]
 Item Value
@@
 Instruction Tuning
+epochs(light) 10
+epochs(small, medium) 5
+epochs(large, xlarge) 3
 batchsize 32
 lr 3e-4
+warmup_ratio 0.03
 cutoff_len 1024
@@
 Appendix D Data Construction
-In this section, we introduce the construction details of TOMG-Bench and OpenMolIns dataset, as well as the prompt templates.
-Figure 4 demonstrates the overall data construction process for TOMG-Bench as well as the OpenMolIns.
-Generally, we utilize Zinc250K to construct the test samples in TOMG-Bench, while adopting PubChem for the training samples in OpenMolIns.
+In this section, we introduce the construction details of S2-Bench and OpenMolIns dataset, as well as the prompt templates.
+Figure 4 demonstrates the overall data construction process for S2-Bench as well as the OpenMolIns.
+Generally, we utilize Zinc250K to construct the test samples in S2-Bench, while adopting PubChem for the training samples in OpenMolIns.
 Notably, when constructing the OpenMolIns, we exclude all the molecules in Zinc250K to avoid data leakage and ensure the novelty score of the generated molecules.
-Figure 4: Data construction workflow of TOMG-Bench & OpenMolIns.
+Figure 4: Data construction workflow of S2-Bench & OpenMolIns.
 D.1 MolEdit
@@
 Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time.
-In contrast, TOMG-Bench requires one single LLM to optimize molecules with different metrics and optimization directions.
+In contrast, S2-Bench requires one single LLM to optimize molecules with different metrics and optimization directions.
 In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED.
@@
 A higher QED score suggests a greater likelihood that the molecule will have favourable pharmacological properties.
+Among the three subtasks, the QED optimization, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization.
+In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
 D.3 MolCustom
@@
 0: Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Group-to-add $\alpha$, Group-to-delete $\Delta$
-0:  Pass or Not (Bool) if $t$ is AddComponent then    if $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false else if $t$ is DelComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 then        return true     else        return false else if $t$ is SubComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 and $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false
+0: Pass or Not (Bool) if $t$ is AddComponent then    if $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false else if $t$ is DelComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ - 1 then        return true     else        return false else if $t$ is SubComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ - 1 and $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false
+Algorithm 2 MolOpt Testing Process
+0: Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Optimization Direction Requirement $R$
+0: Pass or Not (Bool) if $t$ is LogP then    if LogPLogP($m^{g}$) >> LogPLogP($m^{o})$ and R is $\uparrow$ then        return true     else if LogPLogP($m^{g}$) << LogPLogP($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is MR then    if MRM\!R($m^{g}$) >> MRM\!R($m^{o})$ and R is $\uparrow$ then        return true     else if MRM\!R($m^{g}$) << MRM\!R($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is QED then    if QEDQ\!E\!D($m^{g}$) >> QEDQ\!E\!D($m^{o})$ and R is $\uparrow$ then        return true     else if QEDQ\!E\!D($m^{g}$) << QEDQ\!E\!D($m^{o})$ and R is $\downarrow$ then        return true     else        return false
 The testing process for MolEdit task is demonstrated in Algorithm 1.
@@
 Therefore, the similarity score needs to be considered to ensure the accuracy and quality of the generation.
-Algorithm 2 MolOpt Testing Process
-0:  : Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Optimization Direction Requirement $R$
-0:  : Pass or Not (Bool) if $t$ is LogP then    if $LogP$($m^{g}$) $>$ $LogP$($m^{o})$ and R is $\uparrow$ then        return true     else if $LogP$($m^{g}$) $<$ $LogP$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is MR then    if $M\!R$($m^{g}$) $>$ $M\!R$($m^{o})$ and R is $\uparrow$ then        return true     else if $M\!R$($m^{g}$) $<$ $M\!R$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is QED then    if $Q\!E\!D$($m^{g}$) $>$ $Q\!E\!D$($m^{o})$ and R is $\uparrow$ then        return true     else if $Q\!E\!D$($m^{g}$) $<$ $Q\!E\!D$($m^{o})$ and R is $\downarrow$ then        return true     else        return false
 For MolOpt, we adopt Algorithm 2 for the testing process. Similarly, merely passing the test does not necessarily mean that the generated molecule follows the principle of minimal change.
 Algorithm 3 MolCustom Testing Process
-0:  : Generated Molecule $m^{g}$, Subtask $t$, Atom List $A$, Bond List $B$, Functional Group List $G$, Requirements $R$
-0:  : Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for $atom$ in $A$ do         if $|atom|$($m^{g}$) $\neq$ $R$[$atom$then            flag = false else if $t$ is BondNum then    for $bond$ in $B$ do         if $|bond|$($m^{g}$) $\neq$ $R$[$bond$then            flag = false else if $t$ is FunctionalGroup then    for $group$ in $G$ do         if $|group|$($m^{g}$) $\neq$ $R$[$group$then            flag = false return flag
+0: Generated Molecule $m^{g}$, Subtask $t$, Atom List AA, Bond List $B$, Functional Group List $G$, Requirements $R$
+0: Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for atomatom in AA do         if |atom||atom|($m^{g}$) $\neq$ RR[atomatomthen            flag = false else if $t$ is BondNum then    for bondbond in $B$ do         if |bond||bond|($m^{g}$) $\neq$ RR[bondbondthen            flag = false else if $t$ is FunctionalGroup then    for groupgroup in $G$ do         if |group||group|($m^{g}$) $\neq$ RR[groupgroupthen            flag = false return flag
 For MolCustom, the testing process is shown in Algorithm 3, which test whether the specified number of the atom, bond, or functional group satisfies the requirements.
 Appendix F Detailed Results
-In this section, we first show the leaderboard of TOMG-Bench in Table 12, where Claude-3.5 achieves first place with an average weighted success rate of 35.92%.
-Notably, via instruction tuning on our OpenMolIns dataset, Llama-3.1-8B achieves 6th place, which outperforms all the existing open-source general LLMs.
-Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
-Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
-Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
-GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
-GPT-4o (OpenAI, 2023) - 49.08 32.29
-Claude-3 (Anthropic, 2024a) - 46.14 30.47
-Llama-3.1-8B (OpenMolIns-large) 8 43.1 27.22
-Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
-Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
-Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
-Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
-GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
-Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
-Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
-Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
-chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
-Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
-Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
-yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
-Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
-BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
-MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
-Llama-3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
-MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
-MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
-Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
-Table 12: Leaderboard of TOMG-Benchmark.
-Then, we compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5.
-Obviously, it can be concluded that Claude-3.5 outperforms almost all the other LLMs in these subtasks, except for the DelComponent task and the BondNum task.
-In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5, while Llama-3.1-8B (OpenMolIns-large) achieves the best performance in the BondNum task.
+In this section, we first show the leaderboard of S2-Bench in Table 12, where Claude-3.5 achieves the second place with an average weighted success rate of 35.92%.
+Notably, via instruction tuning on our OpenMolIns dataset, Llama3.1-8B achieves the first place, which outperforms all the other LLMs.
+Rank Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
+1 Llama3.1-8B (OpenMolIns-xlarge) 8 58.79 39.33
+2 Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
+3 Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
+4 GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
+5 GPT-4o (OpenAI, 2023) - 49.08 32.29
+6 Claude-3 (Anthropic, 2024a) - 46.14 30.47
+7 Llama3.1-8B (OpenMolIns-large) 8 43.1 27.22
+8 Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
+9 Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
+10 Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
+11 Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
+12 GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
+13 Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
+14 Gemma3-12B (Team et al., 2025) 12 26.28 15.00
+15 Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 7 25.07 14.61
+16 Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
+17 Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
+18 chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
+19 Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
+20 Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
+21 yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
+22 Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
+23 BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
+24 MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
+25 Llama3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
+26 MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
+27 MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
+28 Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
+Table 12: Leaderboard of S2-Bench.
+Figure 5: Performance Comparison across several representative LLMs on S2-Bench.
+We also compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5.
+Among the proprietary models, it can be concluded that Claude-3.5 achieves a powerful performance in these subtasks, except for the DelComponent task and the BondNum task.
+In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5.
+More importantly, we can also observe that Llama3.1-8B (OpenMolIns-xlarge) outperforms most of the other LLMs across these nine subtasks, achieving the SOTA performance.
 Meanwhile, these radar maps also imply the difficulties of the subtasks.
@@
 Even the most advanced proprietary LLMs like Claude-3.5 failed to achieve an ideal performance.
-Moreover, in the MolOpt task, QED, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization.
-In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
-Figure 5: The radar maps compare the performance ($\overline{W\!S\!R}$) of some representative LLMs from different model categories in each subtask of TOMG-Bench.
+Finally, we also include a reasoning LLM, Deepseek-R1-distill-Qwen-7B (Guo et al., 2025), as shown in Table 12, which achieves the best performance among 7 and 8B level open-source general LLMs, showing the potential of LLM reasoning in open-domain natural language-driven molecule generation tasks.
 Finally, we present the detailed experimental results of all the subtasks in Table 13, 14, and 15.
+Note that we do not provide results of vanilla galactica-125M, for it’s not instruction tuned and cannot respond to human instructions.
+All results of vanilla galactica-125M should be close to zeros.
 Models AddComponent DelComponent SubComponent
@@
 Chatglm-9B (GLM et al., 2024) 0.2932 0.7622 0.2235 0.5686 0.2956 0.7494 0.2215 0.6914 0.1498 0.7150 0.1071 0.5084
-Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0374 0.5343 0.0200 0.1982 0.0768 0.5750 0.0442 0.3028 0.0102 0.3671 0.0037 0.1468
+Llama3.2-1B-Instruct (Dubey et al., 2024) 0.0374 0.5343 0.0200 0.1982 0.0768 0.5750 0.0442 0.3028 0.0102 0.3671 0.0037 0.1468
+Gemma3-12B (Team et al., 2025) 0.4300 0.7081 0.3045 0.6324 0.4948 0.5158 0.2552 0.7168 0.4208 0.7154 0.3010 0.7616
+Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.2667 0.6904 0.1841 0.0090 0.6367 0.5921 0.3770 0.0736 0.2156 0.7380 0.1591 0.0356
 MolT5-small (Edwards et al., 2022) 0.1220 0.1027 0.0125 0.4490 0.1598 0.1125 0.0180 0.4504 0.0708 0.1029 0.0073 0.4876
@@
 BioT5-base (Pei et al., 2023) 0.3462 0.1567 0.0542 1.0000 0.1668 0.1597 0.0266 1.0000 0.0684 0.1576 0.0108 0.9998
-Llama-3.2-1B (OpenMolIns-large) 0.1756 0.5676 0.0997 0.3216 0.1816 0.4963 0.0901 0.2466 0.0844 0.5415 0.0457 0.2958
-Llama-3.1-8B (OpenMolIns-large) 0.5822 0.6541 0.3808 0.6730 0.5104 0.5074 0.2590 0.6896 0.5440 0.6258 0.3404 0.8400
+Llama3.2-1B (OpenMolIns-large) 0.1756 0.5676 0.0997 0.3216 0.1816 0.4963 0.0901 0.2466 0.0844 0.5415 0.0457 0.2958
+Llama3.1-8B (OpenMolIns-large) 0.5822 0.6541 0.3808 0.6730 0.5104 0.5074 0.2590 0.6896 0.5440 0.6258 0.3404 0.8400
+Llama3.1-8B (OpenMolIns-xlarge) 0.7790 0.6769 0.5273 0.9468 0.8640 0.6166 0.5327 0.9132 0.6100 0.7434 0.4535 0.9596
 Galactica-125M (OpenMolIns-light) 0.3786 0.5958 0.2256 0.6842 0.2062 0.6521 0.1345 0.7048 0.3102 0.5879 0.1824 0.6674
@@
 Galactica-125M (OpenMolIns-xlarge) 0.5842 0.5859 0.3423 0.8438 0.6526 0.5084 0.3318 0.8286 0.1872 0.6024 0.1128 0.8538
-Table 13: Detailed results on MolEdit. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
+Table 13: Detailed results on MolEdit. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
 Models LogP MR QED
@@
 Chatglm-9B (GLM et al., 2024) 0.3666 0.6902 0.2530 0.4736 0.3514 0.6820 0.2397 0.5000 0.1832 0.6506 0.1192 0.4342
-Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0644 0.5055 0.0326 0.1664 0.0822 0.4410 0.0363 0.1604 0.0714 0.4757 0.0340 0.1796
+Llama3.2-1B-Instruct (Dubey et al., 2024) 0.0644 0.5055 0.0326 0.1664 0.0822 0.4410 0.0363 0.1604 0.0714 0.4757 0.0340 0.1796
+Gemma3-12B (Team et al., 2025) 0.4528 0.694 0.3142 0.5912 0.3256 0.3548 0.1155 0.4478 0.1666 0.0869 0.0145 0.3322
+Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.4048 0.4416 0.1788 0.5622 0.2896 0.4535 0.1313 0.2034 0.0835 0.4556 0.0380 0.6074
 MolT5-small (Edwards et al., 2022) 0.2158 0.1052 0.0227 0.4302 0.2316 0.1011 0.0234 0.4420 0.2214 0.1031 0.0228 0.4326
@@
 BioT5-base (Pei et al., 2023) 0.5158 0.1526 0.0787 1.0000 0.5060 0.1597 0.0808 1.0000 0.5068 0.1580 0.0801 1.0000
-Llama-3.2-1B (OpenMolIns-large) 0.2898 0.5951 0.1725 0.3850 0.2644 0.5956 0.1575 0.3678 0.1996 0.5849 0.1167 0.3490
-Llama-3.1-8B (OpenMolIns-large) 0.8054 0.6678 0.5378 0.8720 0.7122 0.6548 0.4663 0.8514 0.5224 0.6398 0.3342 0.8802
+Llama3.2-1B (OpenMolIns-large) 0.2898 0.5951 0.1725 0.3850 0.2644 0.5956 0.1575 0.3678 0.1996 0.5849 0.1167 0.3490
+Llama3.1-8B (OpenMolIns-large) 0.8054 0.6678 0.5378 0.8720 0.7122 0.6548 0.4663 0.8514 0.5224 0.6398 0.3342 0.8802
+Llama3.1-8B (OpenMolIns-xlarge) 0.8822 0.6662 0.5877 0.9314 0.6982 0.6693 0.4673 0.9422 0.8648 0.6736 0.5825 0.9310
 Galactica-125M (OpenMolIns-light) 0.3202 0.6547 0.2096 0.6416 0.3508 0.6435 0.2257 0.6358 0.2690 0.6521 0.1754 0.6380
@@
 Galactica-125M (OpenMolIns-xlarge) 0.7362 0.5744 0.4229 0.8902 0.7124 0.5697 0.4059 0.8612 0.5786 0.5677 0.3285 0.8626
-Table 14: Detailed results on MolOpt. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
+Table 14: Detailed results on MolOpt. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
 Models AtomNum BondNum FunctionalGroup
@@
 Chatglm-9B (GLM et al., 2024) 0.0002 0.7483 0.0001 0.2131 0.0254 0.7189 0.0183 0.4682 0.0000 0.6908 0.0000 0.5926
-Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0040 0.6807 0.0027 0.1850 0.0080 0.7465 0.0060 0.2226 0.0008 0.7461 0.0006 0.2818
+Llama3.2-1B-Instruct (Dubey et al., 2024) 0.0040 0.6807 0.0027 0.1850 0.0080 0.7465 0.0060 0.2226 0.0008 0.7461 0.0006 0.2818
+Gemma3-12B (Team et al., 2025) 0.0460 0.6017 0.0277 0.3414 0.0280 0.5973 0.0167 0.3318 0.0006 0.5940 0.0004 0.3224
+Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.1574 0.6611 0.1041 0.4815 0.1631 0.7264 0.1185 0.5758 0.0385 0.6248 0.0241 0.5634
 MolT5-small (Edwards et al., 2022) 0.0006 0.6586 0.0004 0.6610 0.0064 0.5980 0.0038 0.6202 0.0114 0.5287 0.0060 0.8354
@@
 BioT5-base (Pei et al., 2023) 0.0118 0.8353 0.0099 0.9950 0.0078 0.6667 0.0052 0.9992 0.0476 0.6792 0.0323 0.9998
-Llama-3.2-1B (OpenMolIns-large) 0.0144 0.6490 0.0093 0.5616 0.0350 0.6150 0.0215 0.6186 0.0252 0.6373 0.0161 0.4412
-Llama-3.1-8B (OpenMolIns-large) 0.0136 0.6634 0.0090 0.7582 0.0544 0.6614 0.0360 0.7456 0.1344 0.6396 0.0860 0.6435
+Llama3.2-1B (OpenMolIns-large) 0.0144 0.6490 0.0093 0.5616 0.0350 0.6150 0.0215 0.6186 0.0252 0.6373 0.0161 0.4412
+Llama3.1-8B (OpenMolIns-large) 0.0136 0.6634 0.0090 0.7582 0.0544 0.6614 0.0360 0.7456 0.1344 0.6396 0.0860 0.6435
+Llama3.1-8B (OpenMolIns-xlarge) 0.1186 0.6834 0.0811 0.8540 0.1196 0.6746 0.0807 0.9000 0.3548 0.6393 0.2268 0.9492
 Galactica-125M (OpenMolIns-light) 0.0044 0.6054 0.0026 0.7930 0.0216 0.5724 0.0124 0.7596 0.0244 0.5756 0.0140 0.8442
@@
 Galactica-125M (OpenMolIns-xlarge) 0.1862 0.6899 0.1285 0.9308 0.1656 0.6887 0.1140 0.7952 0.2006 0.6445 0.1293 0.9162
-Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
+Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
+Appendix G Limitations
+Although S2-Bench is carefully designed and well-validated through our experiments, we still observe several limitations:
+Data Distribution.
+In our data construction process, we allocate distributions to atoms, bonds, and functional groups with the aim of making our benchmark more reflective of real-world distributions.
+Nevertheless, the distribution we use is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios accurately.
+The Selection of Subtasks.
+For the selection of subtasks, there are various choices in MolOpt, we compeletly understand that there are multiple viable options, especially when there are enormous metrics for evaluating molecular properties.
+However, we need to 1) consider the consistency between the three major tasks in S2-Bench and balance their weights (which means we can not select too many metrics for MolOpt, otherwise it would make the benchmark solely an optimization benchmark), and 2) ensure that the metrics can be efficiently and robustly calculated.
+Therefore, we only select the three common metrics, LogP, MR, and QRD, for MolOpt.
+Appendix H Case Study
+Finally, we include the case study to better illustrate the performance gap among different models.
+Figure 6: Case study of AddComponent.
+In this example, we require LLMs to add a hydroxyl to the given molecule.
+Here, GPT-4o and Claude-3.5 barely modified the molecule and failed to add a hydroxyl to the given molecule, while other four models successfully passed the testing process.
+However, it can be clearly noted that although BioT5-base passed the testing process, the generated molecule of BioT5-base could be quite different from the original molecule.
+Figure 7: Case study of DelComponent.
+In this example, we require LLMs to remove an amine from the given molecule.
+Here, GPT-4o, Claude-3.5, and BioT5-base all removed two amines, while the remaining models successfully passed the testing process.
+Figure 8: Case study of SubComponent.
+In this example, we require LLMs to substitute the hydroxyl from the given molecule by a aldehyde.
+Here, although BioT5-base added a aldehyde, it failed to remove the hydroxyl.
+In contrast, the remaining models successfully passed the testing process.
+Figure 9: Case study of LogP.
+In this example, we require LLMs to modfify the molecule to have a higher LogP value.
+Here, after the calculation of LogP values, we found all the generated molecules satisfy the requirement.
+Figure 10: Case study of MR.
+In this example, we require LLMs to modify the molecule to have a higher MR value.
+Here, Llama3.1-8B-Instruct and Llama3.1-8B-Instruct (OpenMolIns-large) surprisingly failed the testing process.
+From the two errors, we could observe that the two models were trying to introduce extra rings to increase the MR value, while the ring numbers were misassigned.
+Figure 11: Case study of QED.
+In this example, we require LLMs to lower the QED value of the given molecule.
+Compared to LogP and MR, QED is more difficult to predict.
+We could observe that GPT-4o, BioT5-base, Llama3.1-8B-Instruct, and Llama3.1-8B-Instruct (OpenMolIns-large) all increased the QED value of the molecule, while Claude-3.5 and Galactica-125M (OpenMolIns-xlarge) successfully captured the requirement and lowered the QED value.
+Figure 12: Case study of AtomNum.
+In this example, we require LLMs to generate a molecule with 15 carbon atoms and 2 bromine atoms.
+Here, GPT-4o, Claude-3.5, BioT5-base, and Llama3.1-8B-Instruct all generated incorrect numbers of the specified atoms.
+Figure 13: Case study of BondNum.
+In this example, we require LLMs to generate a molecule with 6 aromatic bonds, which normally exist in benzene rings.
+Here, GPT-4o, Claude-3.5, Llama3.1-8B-Instruct (OpenMolIns-large), and Galactica-125M (OpenMolIns-xlarge) captured the concept and generated molecules with a benzene ring in their structures.
+In contrast, BioT5-base failed to identify the requirement and Llama3.1-8B-Instruct generated too may aromatic rings.
+Figure 14: Case study of FunctionalGroup. In this example, we require LLMs to generate a molecule with 2 benzene rings. Here, only Claude-3.5 failed to generate the correct number of benzene rings.

### v4

@@
-Jiatong Li [^1] footnotemark: 1    Junxian Li3    Weida Wang2,4    Yunqing Liu1    Changmeng Zheng1 Dongzhan Zhou [^2] footnotemark: 2    Xiao-yong Wei1    Qing Li1
-Abstract
-Recently, Large Language Models (LLMs) have shown great potential in natural language-driven molecule discovery.
-However, existing datasets and benchmarks for molecule-text alignment are predominantly built on a one-to-one mapping, measuring LLMs’ ability to retrieve a single, pre-defined answer, rather than their creative potential to generate diverse, yet equally valid, molecular candidates.
+Speak-to-Structure: Evaluating LLMs in Open-domain Natural Language-Driven Molecule Generation
+Abstract.
+Recently, Large Language Models (LLMs) have demonstrated great potential in natural language-driven molecule discovery.
+However, existing datasets and benchmarks for molecule-text alignment are predominantly built on one-to-one mappings, measuring LLMs’ ability to retrieve a single, pre-defined answer, rather than their creative potential to generate diverse, yet equally valid, molecular candidates.
 To address this critical gap, we propose Speak-to-Structure (S2-Bench), the first benchmark to evaluate LLMs in open-domain natural language-driven molecule generation.
-S2-Bench is specifically designed for one-to-many relationships, challenging LLMs to demonstrate genuine molecular understanding and generation capabilities.
+S2-Bench is specifically designed for one-to-many relationships, challenging LLMs to exhibit genuine molecular understanding and open-ended generation capabilities.
 Our benchmark includes three key tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom), each probing a different aspect of molecule discovery.
-We also introduce OpenMolIns, a large-scale instruction tuning dataset that enables Llama-3.1-8B to surpass the most powerful LLMs like GPT-4o and Claude-3.5 on S2-Bench.
-Our comprehensive evaluation of 28 LLMs shifts the focus from simple pattern recall to realistic molecular design, paving the way for more capable LLMs in natural language-driven molecule discovery.
-1 Introduction
-Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023).
-Traditionally, this is a trial-and-error process (Ekins, 2024) that requires extensive experimentation and data analysis (Mattern & Grosser, 2023), often taking over a decade to bring a new drug candidate to market (Lee et al., 2018).
-Recently, Large Language Models (LLMs) have demonstrated great potential in molecule discovery (Edwards et al., 2021; 2022) by leveraging their powerful language understanding and robust reasoning abilities (OpenAI, 2023).
-Molecules can be represented as SMILES strings (Weininger, 1988), which allows LLMs to process molecules as textual strings, effectively bridging the gap between molecular structures and natural language.
-Meanwhile, by aligning molecules with textual descriptions (Edwards et al., 2022), LLMs can assist chemists in interpreting chemical knowledge, suggesting structural modifications, and predicting compound properties (Li et al., 2024a; Zhang et al., 2024), thereby significantly streamlining the molecule discovery process.
+We also introduce OpenMolIns, a large-scale instruction tuning dataset that enables Llama3.1-8B to surpass the most powerful LLMs like GPT-4o and Claude-3.5 on S2-Bench.
+Our comprehensive evaluation of 31 LLMs shifts the focus from simple pattern recall to realistic molecular design, paving the way for more capable LLMs in natural language-driven molecule discovery.
+Our codes and datasets are fully accessible through the Github Repository: https://github.com/phenixace/S2-TOMG-Bench and Huggingface Datasets: https://huggingface.co/datasets/phenixace/S2-TOMG-Bench.
+Large Language Models, Benchmark, Text-based Molecule Generation, Molecular Design, Open-ended Generation, Instruction Tuning, Bioinformatics
+copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 9–13, 2026; Jeju Island, Republic of Korea.booktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD 2026), August 9–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770855.3817473isbn: 979-8-4007-2259-2/2026/08ccs: Applied computing Bioinformaticsccs: Applied computing Molecular sequence analysisccs: Computing methodologies Natural language processing
+1. Introduction
+Molecule discovery plays a pivotal role in various scientific research fields (Liu et al., 2026, 2025; Zhang et al., 2025; Cao et al., 2022; Li et al., 2026b), from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023).
+Traditionally, this is a trial-and-error process (Ekins, 2024) that requires extensive experimentation and data analysis (Mattern and Grosser, 2023), often taking over a decade to bring a new drug candidate to market (Lee et al., 2018).
+Recently, Large Language Models (LLMs) have demonstrated great potential in molecule discovery (Edwards et al., 2021, 2022; Chitsaz et al., 2025; Wang et al., 2025; Li et al., 2025c) by leveraging their powerful language understanding and robust reasoning abilities (OpenAI, 2023).
+Molecules can be represented as linearized representations like SMILES (Weininger, 1988) and SELFIES (Krenn et al., 2022), which allows LLMs to process molecules as textual strings, effectively bridging the gap between molecular structures and natural language.
+Meanwhile, by aligning molecules with textual descriptions (Edwards et al., 2022; Chen et al., 2025), LLMs can assist chemists in interpreting chemical knowledge, suggesting structural modifications, and predicting compound properties (Li et al., 2025a; Zhang et al., 2024), thereby significantly streamlining the molecule discovery process.
 While the integration of LLMs holds immense promise, a significant challenge lies in the limitations of current datasets and benchmarks.
-For example, existing datasets for aligning molecules and texts like ChEBI-20 (Edwards et al., 2022) and PubChem324K (Liu et al., 2023), are constructed based on a one-to-one mapping assumption, where each textual description is linked to a single, predefined target molecule due to the convenience of acquiring ground truth data from existing databases.
+For example, existing datasets for aligning molecules and texts like ChEBI-20 (Edwards et al., 2022), PubChem324K (Liu et al., 2023), and M3-20M (Guo et al., 2025b), are constructed based on the one-to-one mapping assumption, where each textual description is linked to a single, predefined target molecule due to the convenience of acquiring ground truth data from existing databases.
 However, the one-to-one mapping presents a key limitation: it does not fully align with the nature of real-world molecule discovery.
-In practice, multiple distinct molecules can often share the same desired properties or biological activity Petrone et al.
-(2012).
+In practice, multiple distinct molecules can often share the same desired properties or biological activity (Petrone et al., 2012).
 For instance, a particular pharmaceutical effect is rarely unique to a single compound, and a material with a specific physical property, like tensile strength, can be realized through various molecular structures.
-This mismatch between the evaluation paradigm and the reality of molecule discovery casts doubt on whether these datasets and benchmarks genuinely assess an LLM’s capacity for creative molecular design, or if they instead inadvertently encourage models to rely on memorization and pattern-matching.
+This mismatch between the evaluation paradigm and the reality of molecule discovery casts doubt on whether these datasets and benchmarks genuinely assess the capacity of LLMs for creative molecular design, or if they instead inadvertently encourage models to rely on memorization and pattern-matching.
 To bridge this critical gap, we introduce a novel benchmark, Speak-to-Structure (S2-Bench), the first benchmark designed to evaluate the open-domain natural language-driven molecule generation capabilities of LLMs.
-S2-Bench is meticulously structured into three primary tasks, as shown in Figure 1, each targeting a specific, real-world capability essential for drug and materials science:
+S2-Bench is well structured into three primary tasks, as shown in Figure 1, each targeting a specific, real-world capability essential for drug and materials science:
 Molecule Editing (MolEdit) focuses on testing an LLM’s ability to perform precise, localized structural modifications while preserving main structures.
-In drug discovery, this is analogous to Lead Optimization Barcelos et al.
-(2022), which edits the lead compound to create new variants.
+In drug discovery, this is analogous to Lead Optimization (Barcelos et al., 2022), which edits the lead compound to create new variants.
 This requires a deep understanding of chemical grammar, including valency, stereochemistry, and ring systems, ensuring that edits could create valid molecules.
@@
 Specifically, it challenges models to generate molecules with a set of quantitative and qualitative constraints, such as a specific number of atoms, bonds, or functional groups, which directly mimics the initial drug design phase where a chemist defines a new compound based on a set of precise requirements for its structure.
-Figure 1: Task illustration of S2-Bench for open domain natural language-driven molecule generation.
+Figure 1.
+Task illustration of S2-Bench for open domain natural language-driven molecule generation.
 In contrast to text-based target molecule generation, multiple valid molecules may fulfill the textual requirements (right of the arrow).
-To facilitate the evaluation of the above tasks, we also introduce a robust evaluation system with tailored examination process, enabling a comprehensive assessment of LLMs’ performance in open-domain natural language-driven molecule discovery.
-Furthermore, we propose OpenMolIns, a large-scale instruction-tuning dataset, comprising up to 120 million instruction–molecule pairs, enabling Llama3.1-8B to achieve the state-of-the-art performance among 28 powerful LLMs on S2-Bench.
+To facilitate the evaluation of the above tasks, we introduce an automated evaluation system tailored to open-domain molecule generation, where exact string matching is neither appropriate nor sufficient.
+Specifically, we use cheminformatics toolkits to (i) verify validity by parsing the generated outputs; (ii) check constraint satisfaction (e.g., functional-group edits for MolEdit, property direction for MolOpt, and atom/bond/functional-group constraints for MolCustom); and (iii) quantify quality via task-specific metrics, including structural similarity to the source molecule for MolEdit/MolOpt and novelty against a reference database for MolCustom.
+This design enables a scalable and objective assessment of whether a model follows instructions, performs rational local edits, and generates chemically meaningful candidates.
+Furthermore, we propose OpenMolIns, a large-scale instruction-tuning dataset comprising up to 120 million instruction-molecule pairs constructed programmatically from large chemical databases.
+OpenMolIns mirrors the three tasks in S2-Bench and provides diverse prompt templates to improve robustness to phrasing variations.
 To summarize, our contributions are primarily threefold:
@@
   • 2.
-    We introduce OpenMolIns, an instruction-tuning dataset with up to 120 million instruction–molecule pairs, enabling Llama3.1-8B to achieve SOTA performance on S2-Bench.
+    We introduce OpenMolIns, an instruction-tuning dataset with up to 1.2 million instruction–molecule pairs, enabling Llama3.1-8B to achieve SOTA performance on S2-Bench.
   • 3.
-    We provide insightful findings based on our extensive benchmarking of 28 LLMs, revealing the limitations of existing targeted generation datasets and highlighting the potential for LLMs to transition from simple pattern recall to realistic molecular design.
-2 Speak-to-Structure
+    We provide insightful findings based on our extensive benchmarking of 31 LLMs, revealing the limitations of existing targeted generation datasets and highlighting the potential for LLMs to transition from simple pattern recall to realistic molecular design.
+2. Related Work
+Recent breakthroughs in natural language processing (NLP) have highlighted the potential of Large Language Models (LLMs) to analyze complex biological and chemical data more effectively than traditional computational methods (Zhou et al., 2023).
+Within this context, Text-based Molecule Generation has emerged as a representative task for AI-driven molecule discovery (Edwards et al., 2021).
+This task focuses on generating target molecules directly from natural language descriptions, which requires the construction of paired datasets linking molecular structures with their textual representations.
+Early work in this direction employed transformer-based models such as MolT5 (Edwards et al., 2022), which leveraged large-scale self-supervised learning to produce high-quality SMILES strings from textual prompts.
+Building on this foundation, models like KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023) integrated molecular graphs and biochemical texts to jointly enhance molecular understanding and generation.
+Parallel to these architecture-driven improvements, LLMs such as MolReGPT (Li et al., 2024) and ICMA (Li et al., 2025a) have demonstrated strong in-context learning capabilities, enabling adaptive molecule generation by retrieving and utilizing relevant examples from the provided context.
+Besides, MolReFlect (Li et al., 2026a), emphasizes fine-grained alignment between molecular structures and natural language descriptions through a teacher–student paradigm, which effectively captures subtle structural–semantic relationships.
+Moreover, Mol-R1 (Li et al., 2025b) takes a step further by introducing DeepSeek-R1 like Long-CoT reasoning in molecule generation, enabling robust and interpretable natural language-driven molecule generation.
+Meanwhile, Nicolaou and Brown (2013) propose the task of multi-objective molecular optimization and Wu et al.
+(2024) further extend it to textual requirements by using prompt engineering to guide LLM-based molecule optimization.
+In addition, Ye et al.
+(2025) incorporate explicit structural constraints into the optimization process, while InstructMol (Zhuang et al., 2025) further proposes an LLM that can understand and design biomolecules following human instructions.
+3. Speak-to-Structure (S2-Bench)
 In this section, we detail the design philosophy, technical composition, and statistics of Speak-to-Structure (S2-Bench), which is fundamentally structured around the core capabilities required for real-world molecular design: editing, optimizing, and customized generation.
-2.1 Task Composition
+3.1. Task Composition
 S2-Bench is meticulously designed to move beyond simple pattern-matching and evaluate the genuine molecular understanding and generation capabilities of LLMs.
@@
 In a real-world scenario, this task is analogous to the iterative lead optimization process in drug discovery, where chemists make small, precise changes to a candidate molecule to change its properties.
-We designed three subtasks to probe this capability: AddComponent and DelComponent test the model’s ability to add or remove a specific functional group, respectively, while SubComponent challenges it to perform a combined operation of both.
+We designed three subtasks to probe this capability: AddComponent and DelComponent test the model’s ability to add or remove a specific functional group, respectively, while SubComponent challenges a combined operation of both.
 MolOpt evaluates an LLM’s ability to perform goal-oriented chemical reasoning.
@@
 This task directly mirrors the hit-to-lead optimization stage, where molecules are refined to achieve better pharmacological profiles.
-To assess this, we chose three key properties (LogP, MR, and QED), which are vital for drug discovery and can be reliably and reproducibly calculated.
+To assess this, we choose three key properties (LogP, MR, and QED), which are vital for drug discovery and can be reliably and reproducibly calculated.
 By requiring the model to generate a molecule that improves on a given property, we test its ability to understand the complex relationship between molecular structure and function, a capability that cannot be evaluated by tasks with a single, predetermined answer.
@@
 Our benchmark fundamentally shifts the evaluation from pattern recall to practical chemical design, thereby providing a more accurate measure of an LLM’s potential in molecule discovery and helping to build more explainable and trustworthy models.
-2.2 Data Construction
+3.2. Data Construction
 Previously, the development of robust datasets for text-based targeted molecule generation has been hindered by the scarcity of high-quality human annotations.
@@
 More importantly, this programmatic approach allows us to create tasks with a one-to-many relationship, where multiple correct answers exist for a single prompt, which better aligns with the complexities of drug discovery, accommodating its inherent variability and diverse outcomes.
-Specifically, we design a systematic and scalable data construction process, as illustrated in Figure 4.
+Specifically, we design a systematic and scalable data construction process, as shown in Figure 2.
 For MolEdit and MolOpt, we sample molecules from large, publicly available chemical databases.
-We selected Zinc-250K (Sterling & Irwin, 2015) for building our test set due to its manageable size and diversity, while the massive PubChem (Kim et al., 2019) database (with over 10 million molecules) serves as the basis for the instruction-tuning dataset.
+We selected Zinc-250K (Sterling and Irwin, 2015) for building our test set due to its manageable size and diversity, while the massive PubChem (Kim et al., 2019) database (with over 10 million molecules) serves as the basis for the instruction-tuning dataset.
 We utilize the RDKit (Landrum, 2013) toolbox to automatically extract key molecular statistics, including structural patterns and properties like LogP, MR, and QED.
-These extracted features are then integrated into our pre-defined, instruction-based prompt templates, which is demonstrated in Appendix D.
-For MolCustom, we move beyond existing databases and programmatically generate instructions to test a model’s ability to create novel molecules.
+These extracted features are then integrated into our pre-defined, instruction-based prompt templates.
+For MolCustom, we move beyond existing databases and generate instructions to test a model’s ability to create novel molecules.
 For each of the three subtasks, we randomly generate 5,000 instructions that specify a target number and type of atoms, bonds, or functional groups.
@@
 By moving beyond the one-to-one paradigm, S2-Bench provides a more robust and realistic evaluation, paving the way for models that can truly innovate and design novel molecules.
-2.3 Evaluation
+3.3. Evaluation
 The evaluation of S2-Bench is facilitated through a set of carefully designed automated evaluation processes and metrics tailored to the unique nature of our tasks.
 Unlike one-to-one datasets or benchmarks that simply check for exact matches, our metrics are designed to assess an LLM’s ability to produce valid, relevant, and novel outputs within an open-ended framework.
+Table 1. All used metrics are listed here. ’✓’ means that a metric is calculated on a task.
 Task (Weighted) Success Rate Similarity Novelty Validity
 MolEdit
+MolOpt
 MolCustom
-MolOpt
-Table 1: All used metrics are listed here. ’✓’ means that a metric is calculated on a task.
-2.3.1 MolEdit & MolOpt Evaluation
+3.3.1. MolEdit & MolOpt Evaluation
 For the MolEdit and MolOpt tasks, as shown in Table 1, we employ a combination of metrics to assess both the correctness of the modification and the rationality of the design.
-Success Rate: This metric evaluates a model’s ability to fulfill the specific molecule generation requirements.
+Success Rate: This metric evaluates a model’s ability to fulfill the specific molecule generation requirements, with values ranging from 0 to 1.
 We design automated evaluation processes to verify if the generated molecule meets the specified criteria (e.g., correct functional group modification for MolEdit, or desired property optimization for MolOpt).
-Detailed Implementations of the automated testing process are shown in Appendix E.
 This metric examines whether the LLM can follow instructions precisely, which is a fundamental requirement for an LLM to serve as a chemist assistant.
@@
 A high similarity score indicates that the model has performed a rational, localized edit and has not simply generated a different molecule from scratch.
-The similarity $\delta(m^{g},m^{o})$ between the generated molecule $m^{g}$ and the original molecule $m^{o}$ is calculated as:
-$\delta(m^{g},m^{o})=\frac{\left|fp_{m^{g}}\cap fp_{m^{o}}\right|}{\left|fp_{m^{g}}\cup fp_{m^{o}}\right|},$ (1)
-where $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints.
+The similarity $\delta(m^{g},m^{o})\in[0,1]$ between the generated molecule $m^{g}$ and the original molecule $m^{o}$ is calculated as:
+(1) $\delta(m^{g},m^{o})=\frac{\left|fp_{m^{g}}\cap fp_{m^{o}}\right|}{\left|fp_{m^{g}}\cup fp_{m^{o}}\right|},$
+where $fp_{m^{g}}$ and $fp_{m^{o}}$ represent their corresponding Morgan Fingerprints.
 $\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
-Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax.
-Only the molecules that successfully pass the SMILES parser are considered valid molecules.
-A higher validity means that the model is more familiar with the molecule syntax.
-2.3.2 MolCustom Evaluation
+Validity: This metric evaluates the percentage of the generated molecules that are chemically valid and follow the rules of molecular syntax.
+Molecules need to successfully pass the SMILES parser, and a higher validity means that the model is more familiar with the molecule syntax.
+3.3.2. MolCustom Evaluation
 For the MolCustom task, where models are required to generate molecules from scratch based on structural descriptions, we use a different set of metrics to evaluate their creative design capabilities.
@@
 As the task requires de novo generation, the success rate directly reflects a model’s ability to translate abstract requirements into a valid, concrete molecular structure.
-Detailed testing process of MolCustom could also be found in Appendix E.
 Novelty: For open-domain generation, fulfilling the requirements is just the first step.
@@
 The novelty $n$ for the generated molecule $m^{g}$ can be calculated as:
-$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc}\delta(m^{g},m^{k})}{\left|Zinc\right|},$ (2)
+(2) $n(m^{g})=1-\max_{m^{k}\in\text{Zinc}}\delta(m^{g},m^{k})$
 Validity: As with the other two tasks, validity ensures that all generated molecules are chemically sound and can be interpreted correctly.
-2.3.3 Average Weighted Success Rate
+3.3.3. Average Weighted Success Rate
 To provide a single, comprehensive ranking of LLM performance on S2-Bench, we introduce a weighted success rate.
@@
 The weighted success rate for a subtask $t$ is defined as:
-$W\!S\!R_{t}=\begin{cases}n_{t}\times S\!R_{t},&t\in\{MolCustom\}\\ \delta_{t}\times S\!R_{t},&t\in\{MolEdit,MolOpt\}\end{cases},$ (3)
+(3) $W\!S\!R_{t}=\begin{cases}n_{t}\times S\!R_{t},&t\in\{MolCustom\}\\ \delta_{t}\times S\!R_{t},&t\in\{MolEdit,MolOpt\}\end{cases},$
 where $W\!S\!R_{t}$ denotes the weighted success rate for a subtask $t$, while $\delta_{t}$ is the similarity score for the MolEdit and MolOpt tasks, $n_{t}$ represents the novelty score for the MolCustom tasks, and $S\!R_{t}$ is the corresponding success rate of the subtask.
 Then, the average weighted success rate $\overline{W\!S\!R}$ could evaluate the synthetic performance of LLMs among all the nine subtasks:
-$\overline{W\!S\!R}=\frac{1}{9}\sum_{t}W\!S\!R_{t},$ (4)
+(4) $\overline{W\!S\!R}=\frac{1}{9}\sum_{t}W\!S\!R_{t},$
 where the average weighted success rate $\overline{W\!S\!R}$ provides a balanced measure of a model’s performance across all nine subtasks, offering a robust and nuanced assessment of its capabilities in open-domain natural language-driven molecule generation.
-2.4 OpenMolIns: Instruction Tuning Dataset
+3.4. OpenMolIns: Instruction Tuning Dataset
+Figure 2. Data construction workflow of S2-Bench & OpenMolIns.
 To effectively train and evaluate LLMs on the open-ended challenges posed by S2-Bench, we introduce OpenMolIns, a specialized instruction-tuning dataset derived from the PubChem database.
 The core philosophy behind OpenMolIns is to provide models with the kind of flexible, one-to-many training examples they need to learn genuine chemical reasoning, rather than simply memorizing specific input-output pairs.
-To ensure the integrity of our evaluations, we meticulously designed OpenMolIns to have zero overlap with the Zinc-250K, preventing any data leakage that could compromise the validity of our benchmark results.
+To ensure the integrity of our evaluations, we meticulously designed OpenMolIns to have zero overlap with Zinc-250K, preventing any data leakage that could compromise the validity of our benchmark results.
+Detailed Comparison of the data construction process between S2-Bench and OpenMolIns is shown in Figure 2.
 We built the instruction-tuning dataset programmatically, creating equal numbers of samples (i.e., 5,000) for all nine subtasks.
@@
 This contrasts sharply with previous methods that rely on scarce and costly human annotations.
-Details of the construction process and the pre-defined prompt templates are shown in Appendix D.
 For MolEdit and MolOpt, our goal is to improve the capabilities of LLMs to perform rational, goal-oriented modifications.
@@
 This tiered structure allows researchers to systematically analyze the data scaling law for open-domain natural language-driven molecule generation, providing valuable insights into how data quantity influences a model’s capacity for chemical reasoning.
+Table 2. Statistics of S2-Bench and OpenMolIns.
 Dataset S2-Bench OpenMolIns
@@
 Data Size 5,000 45, 000 4,500 18,000 45,000 90,000 1,200,000
-Table 2: Statisics of S2-Bench and OpenMolIns.
-2.5 Statistics
-In this section, we provide a detailed overview of S2-Bench and OpenMolIns, as summarized in Tables 2 and 3.
+3.5. Statistics
+Here, we provide a detailed overview of S2-Bench and OpenMolIns, as summarized in Tables 2 and 3.
 The design and scale of our datasets are not arbitrary; they are meticulously crafted to enable a rigorous and nuanced evaluation of LLMs in open-domain molecule generation, which previous, smaller datasets cannot provide.
@@
 By observing how model performance improves with increasing data size, we can gain valuable insights into the data requirements for aligning molecular space with natural language.
+Table 3.
+Comparison with existing text-based molecule generation benchmarks and datasets.
+Notably, we denote the statistics of the description-guided molecule design part in Mol-Instructions (Fang et al., 2024).
 Benchmarks or Datasets Task Type # Training Data # Test Data Public Access
 PCdes(Zeng et al., 2022) Targeted Generation 10,500 3000 $\times$
-ChEBI-20 (Edwards et al., 2021) Targeted Generation 26,407 3,000
-PubChem (Liu et al., 2023) Targeted Generation 12,000 2,000
+ChEBI-20 (Edwards et al., 2022) Targeted Generation 26,407 3,000
+PubChem324K (Liu et al., 2023) Targeted Generation 12,000 2,000
 Mol-Instructions (Fang et al., 2024) Targeted Generation 297,319 1000
@@
 S2-Bench & OpenMolIns Open Generation 1,200,000 45,000
-Table 3: Comparison with existing text-based molecule generation benchmarks and datasets. Notably, we denote the statistics of the description guided molecule design part in Mol-Instructions.
 Besides, as shown in Table 3, S2-Bench and OpenMolIns stand out from existing text-based molecule generation datasets.
@@
 This combination of innovative task design and large-scale data represents a new paradigm for evaluating and advancing LLMs in the field of text-guided molecular discovery.
-3 Experiments
-3.1 Models
+We also report molecular distribution statistics for Zinc-250K, PubChem, and our pre-selected samples used in MolEdit and MolOpt.
+Specifically, we compare the average (i) atom count, (ii) ring count, (iii) branch count, and (iv) path length across datasets.
+Table 4. Molecular Distribution Statistics.
+Dataset # Samples Avg Atom Count Avg Ring Count Avg Branch Count Avg Path Length
+Zinc-250K 250,000 23.15 2.76 7.31 12.48
+PubChem 9,000,000 25.18 2.8 7.87 13.07
+S²-Bench’s sampling 30,000 23.17 2.74 7.34 12.47
+As shown in Table 4, Zinc-250K exhibits broadly similar statistical trends to PubChem, suggesting that Zinc-250K captures the core characteristics of drug-like chemistry and remains a reasonable proxy for discovery-scale small molecules.
+Meanwhile, the distribution of our pre-selected samples aligns closely with the original Zinc-250K, suggesting that our pre-selection does not substantially distort the underlying molecular distribution.
+4. Experiments
+4.1. Models
 The models benchmarked are categorized into four groups: proprietary models, open-source general LLMs, open-source ChEBI-20 fine-tuned LLMs, and OpenMolIns fine-tuned LLMs.
@@
 This category includes LLMs that are only accessible via commercial API services.
-In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful reasoning and generalization capabilities.
+In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024a), Claude-3 (Anthropic, 2024b), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful reasoning and generalization capabilities.
 These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
-Open-source General LLMs.
-This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications.
-Specifically, we benchmark Llama3-70B-Instruct, Llama3-8B-Instruct, Llama3.1-8B-Instruct, Llama3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Deepseek-R1-distill-Qwen-7B (Guo et al., 2025), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), chatglm-9B (GLM et al., 2024), and Gemma3-12B (Team et al., 2025).
-These LLMs are top-tier open-source general LLMs within our computation budget.
+Open-source General & Chemical LLMs.
+This group contains open-source general LLMs with instruction following capability for a wide range of tasks, including Llama3-70B-Instruct, Llama3-8B-Instruct, Llama3.1-8B-Instruct, Llama3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), chatglm-9B (GLM et al., 2024), Gemma3-12B (Team et al., 2025).
+Additionally, domain-specific LLMs like ChemLLM-20B (Zhang et al., 2024), SmileyLlama (Cavanagh et al., 2026),and ChemDFM-v1.5-8B (Zhao et al., 2025) are also considered.
 Open-source ChEBI-20 Fine-tuned LLMs.
@@
 We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M because the model has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023).
-Meanwhile, a small size model can also help us study the data scaling law within reasonable budget.
+Meanwhile, a small-sized model can also help us study the data scaling law within a reasonable budget.
 Furthermore, Llama3.2-1B-Instruct and Llama3.1-8B-Instruct are also selected due to their advanced capabilities and similar architectures.
-Figure 2: The performance of LLMs benchmarked in S2-Bench.
-LLMs fall into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs.
-Models of unknown parameters are denoted as horizontal lines.
-3.2 Findings
-Based on our comprehensive benchmarking, we have made the following key observations regarding the capabilities of LLMs in open-domain natural language-driven molecule generation.
-F1: Current LLMs show promise but lack the structural understanding necessary for precise molecule generation.
-As shown in Figure 2, top-tier proprietary models such as Claude-3.5 and Gemini-1.5-Pro attain 35.92% and 34.80% in average weighted success rate, respectively.
-These results suggest that general LLMs hold promise for open-ended molecule generation due to their powerful reasoning capability, yet substantial challenges remain in translating textual requirements into chemically valid and structurally precise molecules.
-Notably, within the MolCustom task, no LLM manages to surpass a 25% weighted success rate on any individual subtask.
-The particularly poor performance on MolCustom tasks highlights that while current LLMs can generate chemically plausible molecules in a broad sense, they struggle with satisfying fine-grained structural constraints such as precise atom, bond, or functional group counts, because the precise control of molecular structures requires reasoning over discrete, globally coupled constraints.
-LLMs, even trained on SMILES, primarily learn local token distributions from large-scale pre-training data, which rarely emphasizes exact numeric constraints.
-The precise control, in fact, is crucial for molecular discovery, as many properties of interest, such as solubility, toxicity, and active site composition, depend on specific structural configurations.
-Our finding suggests that open-domain natural language–driven molecule generation task can serve as an effective diagnostic tool, revealing systematic weaknesses in current LLMs and providing rich failure cases that can inform future instruction tuning or reinforcement learning strategies to improve molecule structural constraint adherence.
-F2: While pre-training provides a foundation for LLM performance, instruction tuning is indispensable for guiding and optimizing their capabilities in molecular discovery.
-Open-source general LLMs can often lack chemical corpora during their pre-training stage.
-As a result, models like Qwen2-7B-Instruct and yi-1.5-9B obtained unsatisfactory results.
-The observed shortcoming can be attributed to the lack of chemical data in pre-training, despite their notable proficiency in mathematics and general language understanding.
-In contrast, Llama3.1-8B-Instruct delivers superior performance at a relatively modest model scale, demonstrating a more balanced and comprehensive capability across diverse domains.
-Meanwhile, our experiments with OpenMolIns highlight the effectiveness of instruction tuning.
-Fine-tuned on the largest OpenMolIns scale (i.e., xlarge), Llama3.1-8B-Instruct surpassed all other LLMs with an average weighted success rates of 39.33%.
-More remarkably, the much smaller model, Galactica-125M, within just 125 million parameters, achieved a weighted average success rate of 25.73%, outperforming models two orders of magnitude larger, such as Llama3-70B-Instruct.
-This remarkable outcome indicates that task-specific instruction tuning on a large, high-quality corpus can be highly effective, suggesting a practical avenue for LLMs to efficiently narrow the performance gap in molecular discovery.
-F3: One-to-one mappings between language instructions and molecules make LLMs inherently rely on pattern recognition and recall.
-Our results also substantiate the observation, demonstrating that existing one-to-one mappings are insufficient for evaluating true chemical comprehension of LLMs.
-Notably, ChEBI-20 Fine-tuned LLMs, which are designed to align molecules with texts, perform worse than general LLMs.
-For instance, BioT5-base (Pei et al., 2023), a state-of-the-art model on the ChEBI-20 dataset, achieves an average weighted success rate of only 4.21% on S2-Bench.
-A closer examination of BioT5-base’s performance provides insight into this limitation.
-Although the model achieves relatively high success rates on MolEdit and MolOpt tasks, its generated molecules often exhibit low similarity to the original structures.
-This indicates that the model is not performing rational edits; rather, it generates different but valid molecules that happen to satisfy the criteria.
-In other words, BioT5-base relies more on recalling patterns from its training corpus than on learning to systematically modify molecular structures.
-Consequently, while the ChEBI-20 dataset is valuable for molecule–caption translation, its limited diversity, scale, and one-to-one mapping nature are insufficient for training LLMs to perform nuanced, open-ended molecular generation.
-Figure 3: Task-specific performance scaling with increasing data in S2-Bench. (a) Galactica-125M. (b) Llama-3.1-8B-Instruct.
-F4: The data scaling law does not always hold: task-specific bottlenecks limit gains in molecule optimization and editing.
-As shown in Figure 3, our experiments with Galactica-125M across five data scales of OpenMolIns provide new insights into the data scaling law for molecule generation.
-Overall, we observe that larger datasets generally improve performance, but the effect is highly task-dependent.
-In MolCustom, scaling the dataset from large to xlarge yields a remarkable 286% average performance gain, indicating that tasks involving complex, de novo synthesis strongly benefit from larger and more diverse training data.
-By contrast, Galactica-125M obtains only modest gains in MolOpt, and even negative gains in MolEdit, suggesting that additional data alone cannot overcome existing bottlenecks in these relatively simpler tasks.
-Interestingly, when examining Llama-3.1-8B-Instruct, we find that performance continues to increase when scaling from the large to xlarge data level, except for LogP and MR subtasks.
-This contrast generally highlights that small models like Galactica-125M are capacity-limited, whereas larger models can more effectively leverage additional data.
-To summarize, these findings suggest that while complex, open-ended tasks are data-driven, simpler property optimization or editing tasks may be constrained more by model capacity than by dataset size.
-Thus, merely enlarging training data may not yield proportional benefits.
-Instead, future progress will require a balanced consideration of both model scale and data scale.
-4 Conclusion
-In this study, we introduced S2-Bench and OpenMolIns, the first benchmark and instruction tuning dataset for evaluating the capabilities of LLMs in open-domain natural language-driven molecule generation.
-By moving beyond traditional one-to-one text-to-molecule mappings, our benchmark accommodates the open-ended nature of molecular discovery.
-S2-Bench focuses on the realistic molecular reasoning and design, rather than simple pattern matching and memorization.
-Our comprehensive benchmarking of 28 LLMs not only highlights the challenges faced by current models and the limitations of existing targeted molecule generation datasets, but also demonstrates the substantial potential of LLMs for natural language–driven molecule discovery.
-Appendix A Related Work
-Molecule discovery is a cornerstone of scientific progress, underpinning advances in both drug development and material design (Du et al., 2022).
-The integration of artificial intelligence into this process has driven a paradigm shift in the pharmaceutical landscape, substantially improving the efficiency and accuracy of identifying and developing novel therapeutic candidates.
-Recent breakthroughs in natural language processing (NLP) have highlighted the potential of Large Language Models (LLMs) to analyze complex biological and chemical data more effectively than traditional computational methods (Zhou et al., 2023).
-Within this context, Text-based Molecule Generation has emerged as a representative task for AI-driven molecule discovery (Edwards et al., 2021).
-This task focuses on generating target molecules directly from natural language descriptions, which requires the construction of paired datasets linking molecular structures with their textual representations.
-Early work in this direction employed transformer-based models such as MolT5 (Edwards et al., 2022), which leveraged large-scale self-supervised learning to produce high-quality SMILES strings from textual prompts.
-Building on this foundation, models like KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023) integrated molecular graphs and biochemical texts to jointly enhance molecular understanding and generation.
-More recently, 3D-MoLM (Li et al., 2024d) advanced the field by incorporating spatial configurations, thereby improving the geometric validity and accuracy of generated molecular structures.
-Parallel to these architecture-driven improvements, LLMs such as MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) have demonstrated strong in-context learning capabilities, enabling adaptive molecule generation by retrieving and utilizing relevant examples from the provided context.
-Besides, MolReFlect (Li et al., 2024c), emphasizes fine-grained alignment between molecular structures and natural language descriptions through a teacher–student paradigm, which effectively captures subtle structural–semantic relationships.
-Moreover, Mol-R1 (Li et al., 2025) takes a step further by introducing DeepSeek-R1 like Long-CoT reasoning in molecule generation, enabling robust and interpretable natural language-driven molecule generation.
-Appendix B Implementation Details
-We implement various scripts to facilitate the testing of the aforementioned models.
-For proprietary models, we adopt the OpenAI API [^1] https://platform.openai.com/docs/ framework.
-For open-source general LLMs, we utilize both the VLLM [^2] https://docs.vllm.ai framework and the OpenAI framework.
-For the remaining LLMs, we adopt the Hugging Face transformers library [^3] https://huggingface.co/docs/ for inference.
-Furthermore, BioT5 is designed to use SELFIES, an alternative string representation of molecules by using special tokens to ensure valid generations (Krenn et al., 2019), as input instead of SMILES.
-Consequently, we convert the molecule SMILES strings into SELFIES format on BioT5.
-Appendix C Hyper Parameters
+4.2. Hyperparameters
 […]
+Table 5. Hyper-parameters.
 Item Value
@@
 dropout 0.1
-Table 4: Hyper-parameters.
-Appendix D Data Construction
-In this section, we introduce the construction details of S2-Bench and OpenMolIns dataset, as well as the prompt templates.
-Figure 4 demonstrates the overall data construction process for S2-Bench as well as the OpenMolIns.
-Generally, we utilize Zinc250K to construct the test samples in S2-Bench, while adopting PubChem for the training samples in OpenMolIns.
-Notably, when constructing the OpenMolIns, we exclude all the molecules in Zinc250K to avoid data leakage and ensure the novelty score of the generated molecules.
-Figure 4: Data construction workflow of S2-Bench & OpenMolIns.
-D.1 MolEdit
-For the MolEdit task, we consider the common operations on modifying functional groups in a given molecule (i.e., add, drop, and substitute), which are simple tasks for human experts but challenging to LLMs.
-In this case, we further develop three corresponding subtasks: AddComponent, DelComponent, and SubComponent.
-Prompt templates for MolEdit are shown in Table 5.
-However, there are different kinds of functional groups, and some functional groups can play an important role in the molecule structure, such as connecting two separate parts of the molecule, which makes them unsuitable for these operations above as these operations will entirely change the structure of the molecule.
-In this case, we aim to make a slight change in the molecule structure and limit most of the functional groups we choose within the end groups.
-Prompt Templates for MolEdit
-AddComponent
-Please add a {} to the molecule {}.
-Modify the molecule {} by adding a {}.
-Add a {} to the molecule {}.
-DelComponent
-Please remove a {} from the molecule {}.
-Modify the molecule {} by removing a {}.
-Remove a {} from the molecule {}.
-SubComponent
-Please substitute a {} in the molecule {} by {}.
-Modify the molecule {} by replacing a {} by {}.
-Replace a {} in the molecule {} by {}.
-Please replace a {} in the molecule {} with {}.
-Modify the molecule {} by substituting a {} with {}.
-Substitute a {} in the molecule {} with {}.
-Table 5: Prompt Templates for MolEdit.
-Table 6 presents the functional groups that are taken into account for AddComponent and DelComponent, along with their respective selection weights.
-Given that certain functional groups are encountered less frequently in real-world scenarios, we have implemented a weighted random selection process for AddComponent, which ensures that less common functional groups are assigned a lower probability to be chosen.
-Functional Group benzene ring hydroxyl aldehyde carboxyl amide
-Weights 15 15 5 5 10
-Functional Group amine nitro halo nitrile thiol
-Weights 5 5 5 1 1
-Table 6: Functional Groups that are considered in AddComponent and DelComponent, as well as their weights to be selected in AddComponent.
-For SubComponent, our focus is exclusively on end groups for simplification, which include hydroxyl, aldehyde, carboxyl, nitro, halo, nitrile, and thiol, which ensures that the editing operations are confined to substituting the existing functional group with another from this list, thereby maintaining the integrity of the molecule’s overall structure without altering it fundamentally.
-D.2 MolOpt
-MolOpt is designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task.
-Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time.
-In contrast, S2-Bench requires one single LLM to optimize molecules with different metrics and optimization directions.
-In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED.
-The prompt templates for MolOpt are listed in Table 7.
-Prompt Templates for MolOpt
-LogP
-Please optimize the molecule {} to have a lower/higher LogP value.
-Modify the molecule {} to decrease/increase its LogP value.
-Optimize the molecule {} to have a lower/higher LogP value.
-Please modify the molecule {} to decrease/increase its LogP value.
-Modify the molecule {} to have a lower/higher LogP value.
-MR
-Please optimize the molecule {} to have a lower/higher MR value.
-Modify the molecule {} to decrease/increase its MR value.
-Optimize the molecule {} to have a lower/higher MR value.
-Please modify the molecule {} to decrease/increase its MR value.
-Modify the molecule {} to have a lower/higher MR value.
-QED
-Please optimize the molecule {} to have a lower/higher QED value.
-Modify the molecule {} to decrease/increase its QED value.
-Optimize the molecule {} to have a lower/higher QED value.
-Please modify the molecule {} to decrease/increase its QED value.
-Modify the molecule {} to have a lower/higher QED value.
-Table 7: Prompt Templates for MolOpt.
-Below are the detailed explanations of the chemical properties adopted in the three subtasks of MolOpt:
-LogP refers to the logarithm of the partition coefficient, which is a measure of a molecule’s hydrophilicity or lipophilicity.
-It is an important factor in determining a compound’s bioavailability and membrane permeability.
-Molecular Refractivity (MR) is a measure of the molar refractive index, which provides insight into the molecular size and the degree of molecular branching.
-It is used to assess the overall shape and bulk of a molecule.
-Quantitative Estimation of Drug-Likeness (QED) is a computational metric that evaluates the drug-likeness of a molecule based on a set of predefined rules.
-A higher QED score suggests a greater likelihood that the molecule will have favourable pharmacological properties.
-Among the three subtasks, the QED optimization, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization.
-In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
-D.3 MolCustom
-To enable customized design of molecules, we consider three fundamental features for describing the molecule, including atoms, bonds, and functional groups.
-Given the specified category and number of atoms, bonds, and functional groups, LLMs should generate the molecule as we request.
-The prompt templates for MolCustom are shown in Table 8.
-Below, we present the construction details of the three subtasks for MolCustom:
-Prompt Templates for MolCustom
-AtomNum
-Please generate a molecule with {} atom(s).
-Please generate a molecule composed of {} atom(s).
-Please generate a molecule consisting {} atom(s).
-The molecule has {} atom(s).
-The molecule is composed of {} atom(s).
-The molecule consists of {} atom(s).
-There is a molecule with {} atom(s).
-There is a molecule composed of {} atom(s).
-There is a molecule consisting of {} atom(s).
-The molecule contains {} atom(s).
-BondNum
-Please generate a molecule with {} bond(s).
-Please generate a molecule composed of {} bond(s).
-Please generate a molecule consisting {} bond(s).
-The molecule has {} bond(s).
-The molecule is composed of {} bond(s).
-The molecule consists of {} bond(s).
-There is a molecule with {} bond(s).
-There is a molecule composed of {} bond(s).
-There is a molecule consisting of {} bond(s).
-The molecule contains {} bond(s).
-FunctionalGroup
-Please generate a molecule with {} group(s).
-Please generate a molecule composed of {} group(s).
-Please generate a molecule consisting {} group(s).
-The molecule has {} group(s).
-The molecule is composed of {} group(s).
-The molecule consists of {} group(s).
-There is a molecule with {} group(s).
-There is a molecule composed of {} group(s).
-There is a molecule consisting of {} group(s).
-The molecule contains {} group(s).
-Table 8: Prompt Templates for MolCustom.
-Atom carbon oxygen nitrogen sulfur fluorine chlorine bromine iodine phosphorus
-Weights [Mandatory] 5 3 3 2 2 2 2 1
-Atom boron silicon selenium tellurium arsenic antimony bismuth polonium
-Weights 1 1 1 1 1 1 1 1
-Table 9: Atoms that are considered in AtomNum, as well as their weights to be selected.
-Higher weights indicate that they are more likely to be selected.
-For simplification, we only consider normal atoms in the molecule structures.
-AtomNum.
-Table 9 shows the atoms we consider in AtomNum, as well as their weights to be selected.
-Notably, carbon, as the basic unit in organic chemicals, is a mandatory option.
-The number of carbon atoms ranges from 1 to 40, while the number of other selected atoms ranges from 1 to 5.
-This setting relieves the difficulty for generation, as LLMs could generate a carbon backbone first and attach the remaining atoms to the backbone one by one.
-Bond single double triple rotatable aromatic
-Weights 5 4 3 1 1
-Table 10: Chemical bonds that are considered in BondNum, as well as their weights to be selected.
-Higher weights indicate that they are more likely to be selected.
-For simplification, we only consider normal bonds in the molecule structures.
-BondNum.
-Similarly, we select five different kinds of chemical bonds: single, double, triple, rotatable, and aromatic, as shown in Table 10.
-For the single bond, if selected, the number can vary from 1 to 50.
-For the aromatic bond, the number follows the rules of the formation of aromatic bonds, varying from 5 to 20.
-Moreover, the number of these remaining bonds, if selected, is specified from 1 to 5.
-Functional Group benzene ring hydroxyl anhydride aldehyde ketone carboxyl ester amide amine nitro
-Weights 15 15 2 5 5 10 5 5 5 2
-Functional Group halo thioether nitrile thiol sulfide disulfide sulfoxide sulfone borane
-Weights 2 1 1 1 1 1 1 1 1
-Table 11: Functional Groups that are considered in FunctionalGroup, as well as their weights to be selected.
-Higher weights indicate that they are more likely to be selected.
-For simplification, we only consider normal and edge functional groups in the molecule structures.
-FunctionalGroup. Lastly, we also specify functional groups in the molecule structure. Table 11 shows the range of functional groups and their weights that are taken into consideration.
-Notably, in MolCustom, if not required, LLMs can generate any number of these unspecified atoms, bonds, and functional groups.
-However, for these specified items, LLMs should strictly follow the requirements.
-Appendix E Testing Process
+Figure 3.
+The performance of LLMs evaluated in S2-Bench.
+LLMs fall into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs.
+Models with unknown parameters are denoted as horizontal lines.
+4.3. Findings
+Based on our comprehensive benchmarking, we have made the following key observations regarding the capabilities of LLMs in open-domain natural language-driven molecule generation.
+F1: Current LLMs show promise but lack the structural understanding necessary for precise molecule generation.
+As shown in Figure 3, top-tier proprietary models such as Claude-3.5 and Gemini-1.5-Pro attain 35.92% and 34.80% in average weighted success rate, respectively.
+These results suggest that general LLMs hold promise for open-ended molecule generation due to their powerful reasoning capability, yet substantial challenges remain in translating textual requirements into chemically valid and structurally precise molecules.
+Notably, within the MolCustom task, no LLM manages to surpass a 25% weighted success rate on any individual subtask.
+The particularly poor performance on MolCustom tasks highlights that while current LLMs can generate chemically plausible molecules in a broad sense, they struggle with satisfying fine-grained structural constraints such as precise atom, bond, or functional group counts, because the precise control of molecular structures requires reasoning over discrete, globally coupled constraints.
+LLMs, even trained on SMILES, primarily learn local token distributions from large-scale pre-training data, which rarely emphasizes exact numeric constraints.
+The precise control, in fact, is crucial for molecular discovery, as many properties of interest, such as solubility, toxicity, and active site composition, depend on specific structural configurations.
+Our findings suggest that open-domain natural language–driven molecule generation task can serve as an effective diagnostic tool, revealing systematic weaknesses in current LLMs and providing rich failure cases that can inform future instruction tuning or reinforcement learning strategies to improve molecule structural constraint adherence.
+F2: While pre-training provides a foundation for LLM performance, instruction tuning is indispensable for guiding and optimizing their capabilities in molecular discovery.
+Open-source general LLMs can often lack chemical corpora during their pre-training stage.
+As a result, models like Qwen2-7B-Instruct and yi-1.5-9B obtained unsatisfactory results.
+The observed shortcoming can be attributed to the lack of chemical data in pre-training, despite their notable proficiency in mathematics and general language understanding.
+Surprisingly, open-source chemical LLMs such as ChemLLM-20B and ChemDFM-v1.5-8B achieve relatively poor performance, indicating that current chemical LLM pre-training still fails to capture the molecular structural details.
+In contrast, Llama3.1-8B-Instruct delivers superior performance at a relatively modest model scale, demonstrating a more balanced and comprehensive capability across diverse domains.
+Meanwhile, our experiments with OpenMolIns highlight the effectiveness of instruction tuning.
+Fine-tuned on the largest OpenMolIns scale (i.e., xlarge), Llama3.1-8B-Instruct surpassed all other LLMs with an average weighted success rate of 39.33%.
+More remarkably, the much smaller model, Galactica-125M, within just 125 million parameters, achieved an average weighted success rate of 25.73%, outperforming models two orders of magnitude larger, such as Llama3-70B-Instruct.
+This remarkable outcome indicates that task-specific instruction tuning on a large, high-quality corpus can be highly effective, suggesting a practical avenue for LLMs to efficiently narrow the performance gap in molecular discovery.
+F3: One-to-one mappings between language instructions and molecules make LLMs inherently rely on pattern recognition and recall.
+Our results also substantiate the observation, demonstrating that existing one-to-one mappings are insufficient for evaluating true chemical comprehension of LLMs.
+Notably, ChEBI-20 Fine-tuned LLMs, which are designed to align molecules with texts, perform worse than general LLMs.
+For instance, BioT5-base (Pei et al., 2023), a state-of-the-art model on the ChEBI-20 dataset, achieves an average weighted success rate of only 4.21% on S2-Bench.
+A closer examination of BioT5-base’s performance provides insight into this limitation.
+Although the model achieves relatively high success rates on MolEdit and MolOpt tasks, its generated molecules often exhibit low similarity to the original structures.
+This indicates that the model is not performing rational edits; rather, it generates different but valid molecules that happen to satisfy the criteria.
+In other words, BioT5-base relies more on recalling patterns from its training corpus than on learning to systematically modify molecular structures.
+Consequently, while the ChEBI-20 dataset is valuable for molecule–caption translation, its limited diversity, scale, and one-to-one mapping nature are insufficient for training LLMs to perform nuanced, open-ended molecular generation.
+Table 6.
+Comparing WSR and SR with Human Expert Evaluation across BioT5-base and Galactica-125M.
+WSR consistently achieves stronger alignment with expert judgment than SR alone.
+Spearman correlations are computed over the six subtasks per model; $^{*}p<0.05$.
+BioT5-base Galactica-125M Human Expert Eval.
+Task Subtask SR WSR SR WSR BioT5-base Galactica-125M
+AddComponent 0.3462 0.0542 0.5842 0.3423 0.64 3.57
+MolEdit DelComponent 0.1668 0.0266 0.6526 0.3318 0.22 3.00
+SubComponent 0.0684 0.0108 0.1872 0.1128 0.19 2.60
+LogP 0.5158 0.0787 0.7362 0.4429 0.58 3.74
+MolOpt MR 0.5060 0.0808 0.7124 0.4059 0.67 3.47
+QED 0.5068 0.0801 0.5786 0.3285 0.64 3.40
+Spearman $\rho$ (vs. Human Expert Eval.) 0.551 $\mathbf{0.899}^{*}$ 0.714 $\mathbf{0.886}^{*}$
+p-value 0.257 0.015 0.111 0.019
+Figure 4. Task-specific performance scaling with increasing data in S2-Bench. (a) Galactica-125M. (b) Llama3.1-8B-Instruct.
+F4: The data scaling law does not always hold: task-specific bottlenecks limit gains in molecule optimization and editing.
+As shown in Figure 4, our experiments with Galactica-125M across the five data scales provide new insights into the data scaling law for molecule generation.
+Overall, we observe that larger datasets generally improve performance, but the effect is highly task-dependent.
+In MolCustom, scaling the dataset from large to xlarge yields a remarkable 286% average performance gain, while there is limited performance improvement from light to large, indicating that tasks involving complex, de novo synthesis strongly benefits from larger and more diverse training data.
+Conversely, Galactica-125M obtains only modest gains in MolOpt, and even negative gains in MolEdit, suggesting that additional data alone cannot overcome existing bottlenecks in these relatively simpler tasks.
+Interestingly, when examining Llama3.1-8B-Instruct, we find that performance continues to increase when scaling from the large to xlarge data level, except for LogP and MR subtasks.
+This contrast generally highlights that small models like Galactica-125M are capacity-limited, whereas larger models can more effectively leverage additional data.
+To summarize, these findings suggest that while complex, open-ended tasks are data-driven, simpler property optimization or editing tasks may be constrained more by model capacity than by dataset size.
+Thus, merely enlarging training data may not yield proportional benefits.
+Instead, future progress will require a balanced consideration of both model scale and data scale.
+4.4. Validating WSR with Human Evaluation
+We conducted an extra correlation analysis to justify our design for using similarity in Weighted Success Rate (WSR).
+Here, we focused on the unusual performance of BioT5-base in MolEdit and MolOpt, where the model achieves high success rates but performs poorly in the quality term, indicating that the generated molecules are not modified from the original molecule but from scratch.
+For this analysis, we first randomly sampled 100 examples from each subtask and then asked human experts to evaluate the correctness and quality of the generated molecules.
+Each example is assigned an expert score from 0 to 5, where:
+  • 5 denotes a perfect success,
+  • 4 means success but with redundant operations (a little different structure),
+  • 3 shows failure but modified from the original molecule (similar structure),
+  • 2 denotes failure and a little different structure,
+  • 1 indicates success but an unrelated/different structure,
+  • 0 denotes failure and an unrelated/different structure.
+We calculate the average score of expert evaluation.
+Table 6 presents per-subtask SR, WSR, and expert scores for BioT5-base and Galactica-125M, with Spearman correlations reported at the bottom.
+WSR achieves significantly stronger alignment with expert judgment than SR alone across both models: for BioT5-base, $\rho_{\mathrm{WSR}}=0.899$ ($p=0.015$) versus $\rho_{\mathrm{SR}}=0.551$ ($p=0.257$); for Galactica-125M, $\rho_{\mathrm{WSR}}=0.886$ ($p=0.019$) versus $\rho_{\mathrm{SR}}=0.714$ ($p=0.111$).
+In both cases WSR reaches statistical significance while SR does not.
+Therefore, it is essential to apply similarities as weights of Success Rate to reflect the genuine capabilities of LLMs more faithfully.
+The results confirm that SR is a noisy proxy for quality: as a binary criterion, it rewards any molecule satisfying the target constraint irrespective of structural distance from the input, allowing models to score high SR by generating chemically valid but structurally unrelated outputs, a pattern human experts consistently penalize.
+WSR corrects for this by down-weighting successes achieved through large structural departures, and the significantly higher Spearman correlations confirm that this re-weighting tracks human judgment more faithfully.
+4.5. Extending MolOpt to Additional Properties
+The choice of LogP, MR, and QED in MolOpt was motivated by the need for automated, reproducible, and unambiguous evaluation at scale, which currently rules out many complex properties.
+To address the concern that this set may be overly simplistic, we extend MolOpt with two more challenging properties: Topological Polar Surface Area (TPSA), widely used to estimate drug permeability, and HOMO-LUMO Gap (HLG), a fundamental quantum-chemical descriptor of molecular reactivity.
+Due to resource constraints, we report results on a set of representative models in Table 7.
+The overall success rates on TPSA and HLG are noticeably lower than those on LogP/MR/QED, confirming that these properties pose a harder optimization challenge and that MolOpt is far from saturated; meanwhile, no single model dominates across both properties, suggesting complementary strengths on different property axes and reinforcing the value of MolOpt as a discriminative benchmark that naturally extends beyond the original three properties.
+Table 7. Extended MolOpt results on TPSA and HOMO-LUMO Gap (HLG). SR denotes Success Rate and WSR denotes Weighted Success Rate.
+Model TPSA SR TPSA WSR HLG SR HLG WSR
+MolT5-large 0.3568 0.0347 0.3460 0.0319
+ChemDFM-v1.5-8B 0.0634 0.0440 0.0920 0.0578
+Llama3.1-8B-Instruct 0.3654 0.1745 0.1760 0.0881
+5. Conclusion
+In this study, we introduced S2-Bench and OpenMolIns, the first benchmark and instruction tuning dataset for evaluating the capabilities of LLMs in open-domain natural language-driven molecule generation.
+By moving beyond traditional one-to-one text-to-molecule mappings, our benchmark accommodates the open-ended nature of molecular discovery.
+S2-Bench focuses on the realistic molecular reasoning and design, rather than simple pattern matching and memorization.
+Our comprehensive benchmarking of 31 LLMs not only highlights the challenges faced by current models and the limitations of existing targeted molecule generation datasets, but also demonstrates the substantial potential of LLMs for natural language–driven molecule discovery.
+Acknowledgments
+This research was supported by the National Natural Science Foundation of China (Grant No.: T2541073, 62372314), the Hong Kong Research Grants Council under the Theme-based Research Scheme (project no.
+T41-517/25-N), and the NSFC/RGC Joint Research Scheme (N_PolyU5179/25).
+The experimental part of this work was supported by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.
+Appendix A Testing Process
 In this section, we present the testing process used to calculate the success rate of each subtask.
@@
 0: Pass or Not (Bool) if $t$ is LogP then if LogPLogP($m^{g}$) >> LogPLogP($m^{o})$ and R is $\uparrow$ then return true else if LogPLogP($m^{g}$) << LogPLogP($m^{o})$ and R is $\downarrow$ then return true else return false else if $t$ is MR then if MRM\!R($m^{g}$) >> MRM\!R($m^{o})$ and R is $\uparrow$ then return true else if MRM\!R($m^{g}$) << MRM\!R($m^{o})$ and R is $\downarrow$ then return true else return false else if $t$ is QED then if QEDQ\!E\!D($m^{g}$) >> QEDQ\!E\!D($m^{o})$ and R is $\uparrow$ then return true else if QEDQ\!E\!D($m^{g}$) << QEDQ\!E\!D($m^{o})$ and R is $\downarrow$ then return true else return false
-The testing process for MolEdit task is demonstrated in Algorithm 1.
-However, in this process, we can not guarantee that the modified molecule differs the original molecule only in the specified functional group.
-The reasons are to follow:
-  • The complexity of molecular structures makes it difficult to use regular matching to monitor the exact number of functional groups present.
-This is because molecules can have a wide variety of bonding patterns and spatial arrangements that are not always amenable to simple, rule-based matching techniques.
-  • Modifying functional groups in specific positions can indeed lead to unintended changes in the molecule’s overall structure.
-This is due to the interconnected nature of atoms within a molecule, where altering one part can have a ripple effect on the stability and geometry of the entire structure.
-  • The tools and algorithms currently available may not have the necessary sophistication to predict and control all the possible outcomes of a functional group modification, particularly when dealing with complex molecules or reactions that are not fully understood.
-Therefore, the similarity score needs to be considered to ensure the accuracy and quality of the generation.
-For MolOpt, we adopt Algorithm 2 for the testing process. Similarly, merely passing the test does not necessarily mean that the generated molecule follows the principle of minimal change.
 Algorithm 3 MolCustom Testing Process
 0: Generated Molecule $m^{g}$, Subtask $t$, Atom List AA, Bond List $B$, Functional Group List $G$, Requirements $R$
-0: Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for atomatom in AA do         if |atom||atom|($m^{g}$) $\neq$ RR[atomatomthen            flag = false else if $t$ is BondNum then    for bondbond in $B$ do         if |bond||bond|($m^{g}$) $\neq$ RR[bondbondthen            flag = false else if $t$ is FunctionalGroup then    for groupgroup in $G$ do         if |group||group|($m^{g}$) $\neq$ RR[groupgroupthen            flag = false return flag
-For MolCustom, the testing process is shown in Algorithm 3, which test whether the specified number of the atom, bond, or functional group satisfies the requirements.
-Appendix F Detailed Results
-In this section, we first show the leaderboard of S2-Bench in Table 12, where Claude-3.5 achieves the second place with an average weighted success rate of 35.92%.
-Notably, via instruction tuning on our OpenMolIns dataset, Llama3.1-8B achieves the first place, which outperforms all the other LLMs.
-Rank Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
-1 Llama3.1-8B (OpenMolIns-xlarge) 8 58.79 39.33
-2 Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
-3 Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
-4 GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
-5 GPT-4o (OpenAI, 2023) - 49.08 32.29
-6 Claude-3 (Anthropic, 2024a) - 46.14 30.47
-7 Llama3.1-8B (OpenMolIns-large) 8 43.1 27.22
-8 Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
-9 Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
-10 Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
-11 Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
-12 GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
-13 Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
-14 Gemma3-12B (Team et al., 2025) 12 26.28 15.00
-15 Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 7 25.07 14.61
-16 Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
-17 Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
-18 chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
-19 Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
-20 Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
-21 yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
-22 Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
-23 BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
-24 MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
-25 Llama3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
-26 MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
-27 MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
-28 Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
-Table 12: Leaderboard of S2-Bench.
-Figure 5: Performance Comparison across several representative LLMs on S2-Bench.
-We also compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5.
-Among the proprietary models, it can be concluded that Claude-3.5 achieves a powerful performance in these subtasks, except for the DelComponent task and the BondNum task.
-In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5.
-More importantly, we can also observe that Llama3.1-8B (OpenMolIns-xlarge) outperforms most of the other LLMs across these nine subtasks, achieving the SOTA performance.
-Meanwhile, these radar maps also imply the difficulties of the subtasks.
-For example, for majority of the LLMs, they are struggling with the BondNum task.
-Even the most advanced proprietary LLMs like Claude-3.5 failed to achieve an ideal performance.
-Finally, we also include a reasoning LLM, Deepseek-R1-distill-Qwen-7B (Guo et al., 2025), as shown in Table 12, which achieves the best performance among 7 and 8B level open-source general LLMs, showing the potential of LLM reasoning in open-domain natural language-driven molecule generation tasks.
-Finally, we present the detailed experimental results of all the subtasks in Table 13, 14, and 15.
-Note that we do not provide results of vanilla galactica-125M, for it’s not instruction tuned and cannot respond to human instructions.
-All results of vanilla galactica-125M should be close to zeros.
+0: Pass or Not (Bool) flag = trueif $t$ is AtomNum then for atomatom in AA do if |atom||atom|($m^{g}$) $\neq$ RR[atomatom] then flag = false else if $t$ is BondNum then for bondbond in $B$ do if |bond||bond|($m^{g}$) $\neq$ RR[bondbond] then flag = false else if $t$ is FunctionalGroup then for groupgroup in $G$ do if |group||group|($m^{g}$) $\neq$ RR[groupgroup] then flag = false return flag
+Appendix B Detailed Results
+In this section, we show the leaderboard of S2-Bench in Table 8.
+The detailed experimental results of all the subtasks in Table 9, 10, and 11.
+Note that we do not provide results of vanilla galactica-125M, for it’s not instruction-tuned and cannot respond to human instructions.
+All results of vanilla galactica-125M should be close to zero.
+Table 8. Leaderboard of S2-Bench.
+Model # Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$ Rank
+Llama3.1-8B (OpenMolIns-xlarge) 8 58.79 39.33 1
+Claude-3.5 (Anthropic, 2024a) - 51.10 35.92 2
+Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80 3
+GPT-4-turbo (OpenAI, 2023) - 50.74 34.23 4
+GPT-4o (OpenAI, 2023) - 49.08 32.29 5
+Claude-3 (Anthropic, 2024b) - 46.14 30.47 6
+Llama3.1-8B (OpenMolIns-large) 8 43.1 27.22 7
+Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73 8
+Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93 9
+Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42 10
+Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89 11
+GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58 12
+Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18 13
+Gemma3-12B (Team et al., 2025) 12 26.28 15.00 14
+Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a) 7 25.07 14.61 15
+Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.12 13.78 16
+Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75 17
+chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7) 18
+Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6) 19
+ChemDFM-v1.5-8B (Zhao et al., 2025) 8 18.24 12.07 20
+ChemLLM-20B (Zhang et al., 2024) 20 16.23 9.76 21
+SmileyLlama (Cavanagh et al., 2026) 8 16.18 8.43 22
+Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10 23
+yi-1.5-9B (Young et al., 2024) 9 14.10 7.32 24
+Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81 25
+BioT5-base (Pei et al., 2023) 0.25 24.19 4.21 26
+MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89 27
+Llama3.2-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99 28
+MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0) 29
+MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9) 30
+Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15 31
+Table 9. Detailed results on MolEdit. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
 Models AddComponent DelComponent SubComponent
@@
 GPT-3.5-turbo (OpenAI, 2023) 0.5832 0.6545 0.3817 0.7980 0.3082 0.7797 0.2403 0.8468 0.2918 0.6333 0.1848 0.6822
-Claude-3.5 (Anthropic, 2024b) 0.6832 0.7017 0.4794 0.4414 0.5414 0.6678 0.3615 0.7960 0.8104 0.7310 0.5924 0.9588
-Claude-3 (Anthropic, 2024a) 0.6766 0.6840 0.4628 0.8180 0.5556 0.6408 0.3560 0.8984 0.6550 0.7159 0.4689 0.9184
+Claude-3.5 (Anthropic, 2024a) 0.6832 0.7017 0.4794 0.4414 0.5414 0.6678 0.3615 0.7960 0.8104 0.7310 0.5924 0.9588
+Claude-3 (Anthropic, 2024b) 0.6766 0.6840 0.4628 0.8180 0.5556 0.6408 0.3560 0.8984 0.6550 0.7159 0.4689 0.9184
 Gemini-1.5-pro (Deepmind, 2024) 0.7058 0.6792 0.4794 0.8254 0.7590 0.5949 0.4515 0.9158 0.7148 0.7139 0.5103 0.8684
@@
 Llama3-8B-Instruct (Dubey et al., 2024) 0.3914 0.6649 0.2602 0.5374 0.4348 0.5058 0.2199 0.5700 0.2602 0.6841 0.1780 0.4838
-Llama3.1-8B-Instruct (Dubey et al., 2024) 0.2992 0.6088 0.1822 0.4954 0.4336 0.5257 0.2279 0.5910 0.3401 0.6424 0.2185 0.5076
+Llama3.1-8B-Instruct (Dubey et al., 2024) 0.2992 0.6088 0.1822 0.4954 0.4416 0.5199 0.2296 0.5832 0.3401 0.6424 0.2185 0.5076
 Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.1868 0.6251 0.1168 0.3760 0.2018 0.3774 0.0762 0.3590 0.0602 0.6227 0.0375 0.3550
@@
 Gemma3-12B (Team et al., 2025) 0.4300 0.7081 0.3045 0.6324 0.4948 0.5158 0.2552 0.7168 0.4208 0.7154 0.3010 0.7616
-Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.2667 0.6904 0.1841 0.0090 0.6367 0.5921 0.3770 0.0736 0.2156 0.7380 0.1591 0.0356
+Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a) 0.2667 0.6904 0.1841 0.0090 0.6367 0.5921 0.3770 0.0736 0.2156 0.7380 0.1591 0.0356
+SmileyLlama (Cavanagh et al., 2026) 0.1990 0.5460 0.1088 0.7700 0.0740 0.5990 0.0445 0.7940 0.0170 0.5870 0.0102 0.7740
+ChemDFM-v1.5-8B (Zhao et al., 2025) 0.2942 0.6454 0.1899 0.9112 0.2056 0.6937 0.1426 0.9286 0.2198 0.6432 0.1414 0.8652
+ChemLLM-20B (Zhang et al., 2024) 0.1436 0.5978 0.0858 0.2108 0.2092 0.5734 0.1200 0.3098 0.3088 0.6805 0.2101 0.4936
 MolT5-small (Edwards et al., 2022) 0.1220 0.1027 0.0125 0.4490 0.1598 0.1125 0.0180 0.4504 0.0708 0.1029 0.0073 0.4876
@@
 Galactica-125M (OpenMolIns-xlarge) 0.5842 0.5859 0.3423 0.8438 0.6526 0.5084 0.3318 0.8286 0.1872 0.6024 0.1128 0.8538
-Table 13: Detailed results on MolEdit. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
+Table 10. Detailed results on MolOpt. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
 Models LogP MR QED
@@
 GPT-3.5-turbo (OpenAI, 2023) 0.4048 0.6327 0.2561 0.8540 0.4120 0.6263 0.2580 0.8486 0.3316 0.5635 0.1869 0.8354
-Claude-3.5 (Anthropic, 2024b) 0.7970 0.7124 0.5678 0.9422 0.6962 0.7112 0.4951 0.9110 0.5361 0.7042 0.3775 0.8604
-Claude-3 (Anthropic, 2024a) 0.7984 0.6067 0.4844 0.9096 0.6094 0.6398 0.3899 0.9062 0.4678 0.5855 0.2739 0.9044
+Claude-3.5 (Anthropic, 2024a) 0.7970 0.7124 0.5678 0.9422 0.6962 0.7112 0.4951 0.9110 0.5361 0.7042 0.3775 0.8604
+Claude-3 (Anthropic, 2024b) 0.7984 0.6067 0.4844 0.9096 0.6094 0.6398 0.3899 0.9062 0.4678 0.5855 0.2739 0.9044
 Gemini-1.5-pro (Deepmind, 2024) 0.7712 0.7022 0.5415 0.9274 0.7876 0.6744 0.5312 0.8926 0.4704 0.6077 0.2859 0.9484
@@
 Llama3-8B-Instruct (Dubey et al., 2024) 0.4642 0.3658 0.1698 0.6086 0.4332 0.4793 0.2076 0.5704 0.2568 0.4547 0.1168 0.6112
-Llama3.1-8B-Instruct (Dubey et al., 2024) 0.3990 0.4235 0.1690 0.5122 0.4336 0.5257 0.2279 0.5910 0.2655 0.4499 0.1194 0.6158
+Llama3.1-8B-Instruct (Dubey et al., 2024) 0.3990 0.4235 0.1690 0.5122 0.4134 0.4794 0.1982 0.5162 0.2655 0.4499 0.1194 0.6158
 Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.2220 0.4501 0.0999 0.2802 0.1908 0.2578 0.0492 0.3795 0.1210 0.3244 0.0393 0.2532
@@
 Gemma3-12B (Team et al., 2025) 0.4528 0.694 0.3142 0.5912 0.3256 0.3548 0.1155 0.4478 0.1666 0.0869 0.0145 0.3322
-Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.4048 0.4416 0.1788 0.5622 0.2896 0.4535 0.1313 0.2034 0.0835 0.4556 0.0380 0.6074
+Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a) 0.4048 0.4416 0.1788 0.5622 0.2896 0.4535 0.1313 0.2034 0.0835 0.4556 0.0380 0.6074
+SmileyLlama (Cavanagh et al., 2026) 0.4520 0.4960 0.2241 0.7860 0.4070 0.5180 0.2111 0.7730 0.2900 0.5170 0.1499 0.7800
+ChemDFM-v1.5-8B (Zhao et al., 2025) 0.2998 0.6613 0.1983 0.7582 0.2510 0.6831 0.1715 0.8170 0.3112 0.6600 0.2054 0.8072
+ChemLLM-20B (Zhang et al., 2024) 0.2992 0.5634 0.1686 0.3788 0.1316 0.5452 0.0717 0.293 0.2594 0.5739 0.1489 0.3856
 MolT5-small (Edwards et al., 2022) 0.2158 0.1052 0.0227 0.4302 0.2316 0.1011 0.0234 0.4420 0.2214 0.1031 0.0228 0.4326
@@
 Galactica-125M (OpenMolIns-xlarge) 0.7362 0.5744 0.4229 0.8902 0.7124 0.5697 0.4059 0.8612 0.5786 0.5677 0.3285 0.8626
-Table 14: Detailed results on MolOpt. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
+Table 11. Detailed results for subtasks on MolCustom. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
 Models AtomNum BondNum FunctionalGroup
@@
 GPT-3.5-turbo (OpenAI, 2023) 0.1070 0.5054 0.0541 0.6947 0.0518 0.6871 0.0356 0.5522 0.1136 0.6585 0.0748 0.8686
-Claude-3.5 (Anthropic, 2024b) 0.1928 0.6926 0.1335 0.6548 0.1058 0.6584 0.0697 0.8860 0.2364 0.6582 0.1556 0.8892
-Claude-3 (Anthropic, 2024a) 0.1044 0.6833 0.0713 0.5910 0.1042 0.6598 0.0688 0.8696 0.1816 0.9158 0.1663 0.6644
+Claude-3.5 (Anthropic, 2024a) 0.1928 0.6926 0.1335 0.6548 0.1058 0.6584 0.0697 0.8860 0.2364 0.6582 0.1556 0.8892
+Claude-3 (Anthropic, 2024b) 0.1044 0.6833 0.0713 0.5910 0.1042 0.6598 0.0688 0.8696 0.1816 0.9158 0.1663 0.6644
 Gemini-1.5-pro (Deepmind, 2024) 0.1742 0.6902 0.1202 0.6774 0.0708 0.6522 0.0462 0.8688 0.2486 0.6673 0.1659 0.9240
@@
 Gemma3-12B (Team et al., 2025) 0.0460 0.6017 0.0277 0.3414 0.0280 0.5973 0.0167 0.3318 0.0006 0.5940 0.0004 0.3224
-Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.1574 0.6611 0.1041 0.4815 0.1631 0.7264 0.1185 0.5758 0.0385 0.6248 0.0241 0.5634
+Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a) 0.1574 0.6611 0.1041 0.4815 0.1631 0.7264 0.1185 0.5758 0.0385 0.6248 0.0241 0.5634
+SmileyLlama (Cavanagh et al., 2026) 0.0002 0.6560 0.0001 0.6420 0.0116 0.5970 0.0069 0.6990 0.0052 0.5970 0.0031 0.7430
+ChemDFM-v1.5-8B (Zhao et al., 2025) 0.0112 0.6809 0.0076 0.6700 0.0172 0.6146 0.0106 0.8968 0.0312 0.6011 0.0188 0.8094
+ChemLLM-20B (Zhang et al., 2024) 0.0288 0.7055 0.0203 0.3530 0.0204 0.6786 0.0138 0.3678 0.0600 0.6452 0.0387 0.5442
 MolT5-small (Edwards et al., 2022) 0.0006 0.6586 0.0004 0.6610 0.0064 0.5980 0.0038 0.6202 0.0114 0.5287 0.0060 0.8354
@@
 Galactica-125M (OpenMolIns-xlarge) 0.1862 0.6899 0.1285 0.9308 0.1656 0.6887 0.1140 0.7952 0.2006 0.6445 0.1293 0.9162
-Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
-Appendix G Limitations
+Appendix C Limitations
 Although S2-Bench is carefully designed and well-validated through our experiments, we still observe several limitations:
-Data Distribution.
-In our data construction process, we allocate distributions to atoms, bonds, and functional groups with the aim of making our benchmark more reflective of real-world distributions.
-Nevertheless, the distribution we use is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios accurately.
-The Selection of Subtasks.
-For the selection of subtasks, there are various choices in MolOpt, we compeletly understand that there are multiple viable options, especially when there are enormous metrics for evaluating molecular properties.
-However, we need to 1) consider the consistency between the three major tasks in S2-Bench and balance their weights (which means we can not select too many metrics for MolOpt, otherwise it would make the benchmark solely an optimization benchmark), and 2) ensure that the metrics can be efficiently and robustly calculated.
-Therefore, we only select the three common metrics, LogP, MR, and QRD, for MolOpt.
-Appendix H Case Study
-Finally, we include the case study to better illustrate the performance gap among different models.
-Figure 6: Case study of AddComponent.
-In this example, we require LLMs to add a hydroxyl to the given molecule.
-Here, GPT-4o and Claude-3.5 barely modified the molecule and failed to add a hydroxyl to the given molecule, while other four models successfully passed the testing process.
-However, it can be clearly noted that although BioT5-base passed the testing process, the generated molecule of BioT5-base could be quite different from the original molecule.
-Figure 7: Case study of DelComponent.
-In this example, we require LLMs to remove an amine from the given molecule.
-Here, GPT-4o, Claude-3.5, and BioT5-base all removed two amines, while the remaining models successfully passed the testing process.
-Figure 8: Case study of SubComponent.
-In this example, we require LLMs to substitute the hydroxyl from the given molecule by a aldehyde.
-Here, although BioT5-base added a aldehyde, it failed to remove the hydroxyl.
-In contrast, the remaining models successfully passed the testing process.
-Figure 9: Case study of LogP.
-In this example, we require LLMs to modfify the molecule to have a higher LogP value.
-Here, after the calculation of LogP values, we found all the generated molecules satisfy the requirement.
-Figure 10: Case study of MR.
-In this example, we require LLMs to modify the molecule to have a higher MR value.
-Here, Llama3.1-8B-Instruct and Llama3.1-8B-Instruct (OpenMolIns-large) surprisingly failed the testing process.
-From the two errors, we could observe that the two models were trying to introduce extra rings to increase the MR value, while the ring numbers were misassigned.
-Figure 11: Case study of QED.
-In this example, we require LLMs to lower the QED value of the given molecule.
-Compared to LogP and MR, QED is more difficult to predict.
-We could observe that GPT-4o, BioT5-base, Llama3.1-8B-Instruct, and Llama3.1-8B-Instruct (OpenMolIns-large) all increased the QED value of the molecule, while Claude-3.5 and Galactica-125M (OpenMolIns-xlarge) successfully captured the requirement and lowered the QED value.
-Figure 12: Case study of AtomNum.
-In this example, we require LLMs to generate a molecule with 15 carbon atoms and 2 bromine atoms.
-Here, GPT-4o, Claude-3.5, BioT5-base, and Llama3.1-8B-Instruct all generated incorrect numbers of the specified atoms.
-Figure 13: Case study of BondNum.
-In this example, we require LLMs to generate a molecule with 6 aromatic bonds, which normally exist in benzene rings.
-Here, GPT-4o, Claude-3.5, Llama3.1-8B-Instruct (OpenMolIns-large), and Galactica-125M (OpenMolIns-xlarge) captured the concept and generated molecules with a benzene ring in their structures.
-In contrast, BioT5-base failed to identify the requirement and Llama3.1-8B-Instruct generated too may aromatic rings.
-Figure 14: Case study of FunctionalGroup. In this example, we require LLMs to generate a molecule with 2 benzene rings. Here, only Claude-3.5 failed to generate the correct number of benzene rings.
+Data Distribution. Our allocation of atoms, bonds, and functional groups is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios.
+Application to Complex Real-world Constraints.
+Incorporating constraints like stereochemistry, 3D geometry, ring/bridge topology, and basic synthesizability would represent more important aspects of real-world molecular design.
+Realism of Instructions. Template-generated prompts may not fully capture the richness of real-world chemist queries.
+Appendix D Broader Impacts
+S2-Bench firstly proposes the text-based open molecule generation task for LLMs, aiming at adopting LLMs as chemist assistants and molecule operators for molecule discovery.
+Hopefully, our benchmark could inspire more research in this promising area, thereby accelerating the pace of drug discovery and material design.
+Ultimately, the development in molecule discovery will benefit human welfare in various ways.
+For example, the effective medicines could help save people’s lives and expand their lifespan.
+Appendix E GenAI Disclosure
+During the preparation of this work, the author(s) used LLMs to improve the language and readability.
+After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.